### v1

Abstract
We formulate the Knapsack Contracts problem—a strategic version of the classic Stochastic Knapsack problem, which builds upon the inherent randomness shared by stochastic optimization and contract design. In this problem, the principal incentivizes agents to perform jobs with stochastic processing times, the realization of which depends on the agents’ efforts.
Algorithmically, we show that Knapsack Contracts can be viewed as Stochastic Knapsack with costs and multi-choice, features that introduce significant new challenges. We identify a crucial and economically meaningful parameter—the Return on Investment (ROI) value. We show that the Inverse of ROI (or IOR for short) precisely characterizes the extent to which the approximation guarantees for Stochastic Knapsack extend to its strategic counterpart.
For IOR of $\alpha$, we develop an algorithm that finds an $O(\alpha)$-approximation policy that does not rely on adaptivity. We establish matching $\Omega(\alpha)$ lower bounds, both on the adaptivity gap, and on what can be achieved without full distributional knowledge of the processing times. Taken together, our results show that IOR is fundamental to understanding the complexity and approximability of Knapsack Contracts, and bounding it is both necessary and sufficient for achieving non-trivial approximation guarantees. Our results highlight the computational challenges arising when stochasticity in optimization problems is controlled by strategic effort.
1 Introduction
The study of how incentives reshape classic algorithmic problems has been at the heart of algorithmic game theory research since its inception. In their seminal work, Nisan and Ronen [24] famously conjectured that adding incentives for truthfulness to the optimization problem of scheduling would make the problem extremely hard to approximate. Two decades later, intensive research has culminated in a proof of this conjecture [9]. Many additional works explore the intersection of algorithm design with truthfulness constraints [27, 25, 10, 6, 8, 22, 23].
Over recent years, the algorithmic game theory community has begun to explore incentive constraints beyond truthfulness. A recent focus is on an agent’s earnestness in exerting effort to complete a task, where the task is delegated to the agent by a principal. The main economic tool for incentivizing earnest effort is performance-based payment schemes, a.k.a. contracts.
Contracts are designed to create incentives that counteract moral hazard: When an agent exerts effort to carry out a job on behalf of a principal, this leads to a stochastic outcome that only partially reflects the agent’s effort. This stochasticity opens the possibility for the agent to claim they worked hard despite exerting a low level of effort, by blaming poor outcomes on “bad luck”. This phenomenon is known as moral hazard. A contract uses outcome-based payments and distributional knowledge to incentivize earnest effort, thus mitigating moral hazard. The goal of this work is to initiate the study of the intersection of algorithm design with contractual incentives for earnest effort.
1.1 Problem Formulation
To achieve our goal, we seek an algorithmic optimization problem with a natural connection to contractual incentives—similar to the connection between scheduling and truthfulness [24]. Our starting point is the observation that the field of contract design has a natural affinity to stochastic optimization, due to the nondeterminism inherent to both. Building upon this connection, we formulate a model of Knapsack Contracts, in which jobs that need to fit into a given time frame are delegated to agents. In our model, the agents have costs for carrying out the jobs, and the completion times (a.k.a. job lengths or sizes) are a stochastic outcome of the level of effort. The latter makes our model a generalization of Stochastic Knapsack [11]. We show that the agents’ costs lead to incentive constraints that make the problem significantly harder than classic Stochastic Knapsack. Since Stochastic Knapsack has a central place in stochastic optimization [3, 19, 20, 18, 28, 5, 2, 21], we view its study through the prism of contract design as a potential gateway to deeper exploration of the intersection between the two areas.
Example 1.
A PI (principal) wishes to run a set of experiments on the cloud before the deadline of a top AI conference. Each experimental task has a fixed value for the paper, which does not depend on how fast the cloud providers (agents) finish running the experiments. The PI does not benefit from experiments that finish running after the submission deadline. Despite this, she still needs to pay the providers for running them. The PI must spend substantial funds from her research grant to incentivize the agents to use powerful but costly resources when processing the experiments, resulting in shorter running times. This leads to an interesting tension between the PI’s benefit from an experiment, and how much they are willing to pay to incentivize an agent to run it quickly. The problem of optimizing the experiments thus has a mixed-sign objective.
Generalizing Example 1, we consider a single principal and multiple agents with jobs that must be executed sequentially. Each agent $i$ is associated with a fixed known value $v_{i}$ to the principal for completing their job before the deadline expires. Moreover, each agent can choose one of $m$ costly actions, and the chosen action $j\in[m]$ determines a distribution $F_{i}^{j}$ over the possible completion times of the job. The principal cannot directly observe the action taken by an agent, and can only inspect the completion times. To incentivize agent $i$, the principal designs a contract $t_{i}$ that maps the completion time $s_{i}$ to a monetary transfer. Each agent $i$ wants to maximize their own utility: the transfer received minus the cost $c_{i}^{j}$ exerted by them when taking action $j$. The principal wants to maximize the profit, which is the sum of values over all jobs completed before the deadline, minus the sum of transfers made to the agents (i.e., a mixed-sign objective). See Section 2 for a formal problem description.
Relation to vanilla contract design.
Usually in contract settings, the value created for the principal by the agent’s effort is stochastic. Our model captures settings in which the principal’s value for the job delegated to the agent is fixed, but the time it takes the agent to complete the job is a random variable, and if the job doesn’t finish on time then the resulting value is zero. This seemingly simple twist on the standard principal-agent model leads to new challenges.
Relation to Stochastic Knapsack.
In the standard Stochastic Knapsack problem, there are $n$ jobs to execute, each with a fixed value but stochastic size. A policy-maker executes jobs sequentially, gaining the value of any job that terminates before the deadline. This process ends once a job terminates after the deadline. The optimal policy for this problem can be highly complex, adaptive, and intractable. Yet, Dean et al. [11] show that there exists a simple non-adaptive policy that approximates, up to a constant factor, the reward of the optimal adaptive policy.
In contrast, in our model, even if a job ends after the deadline, the principal must pay the agent according to the contract, for processing it. As we show in Section 2.2, our model corresponds to a version of Stochastic Knapsack with job costs and multi-choice, and we show this makes a substantial difference relative to [11]. A first indication of the difference arises from the linear programming approach of [11], which they use to establish their main result. In their setting, the integrality gap of the LP is bounded; however, in Appendix A we demonstrate that the LP approach can grossly overshoot the true optimum in our setting. In fact, as we shall see, in our model there is no constant-approximation non-adaptive policy; any constant-approximation algorithm must adaptively choose the next job to execute. Intuitively, adaptivity is necessary to avoid the danger of incurring a very high cost, so high it overshadows any value the algorithm may obtain.
1.2 Our Results and Techniques
Return on Investment.
The example in Appendix A is based on a “hard instance” of Stochastic Knapsack with job costs. We show this instance (and in fact any other hard instance) has the following property: For some of the jobs, their cost almost “covers” their value. Thus, in order to incentivize such jobs, the principal has to pay the agent their cost, and therefore, they are left with a small return on investment (ROI). I.e., the ratio between the obtained profit and the transfer made is small. Such jobs are intuitively problematic—if their cost must be reimbursed by the principal, but their value is not gained since they finish after the deadline, the loss can be substantial. We essentially show that these are the only problematic jobs.
In order to parameterize this, we inspect the Inverse of the ROI (or IOR) of a job (as the ROI becomes small, the IOR grows). We introduce an economically-meaningful parameter $\alpha$, which represents the worst IOR among all jobs—see Definition 2.6. Our main finding is that IOR is precisely the parameter that determines the loss in approximation due to moving from the classic Stochastic Knapsack problem, to a contract design problem subject to incentive constraints:
  • We show that the principal can efficiently and non-adaptively compute a set of contracts, which give an $O(\alpha)$-approximation to the profit of the (intractable and highly adaptive) optimal policy of the principal. See Theorem 3.1.
  • We show the existence of an instance such that any non-adaptive policy of the principal is bounded away by a factor of $\Omega(\alpha)$ in terms of profit from the optimal adaptive policy. This establishes a tight $\Theta(\alpha)$ adaptivity gap for the Knapsack Contracts problem. See Theorem 4.1.
  • We establish an information gap: Consider any policy (adaptive or not) that has limited information about the jobs’ size distributions in the form of finitely many moments, as well as access to the probability of knapsack overflow at every step. No such policy can guarantee an approximation factor better than $O(\alpha)$ with respect to the profit of the optimal policy with full access to the size distributions. See Theorem 5.2.
We now describe our results in more detail as well as the techniques that enable them.
Positive results.
In Section 3, we provide an efficient algorithm that finds a non-adaptive set of contracts (see Section 2 for adaptivity notions), which give an $O(\alpha)$-approximation to the optimal profit the principal can obtain, where $\alpha$ is the IOR parameter (Theorem 3.1). Our algorithm needs to choose in advance the order in which agents are to execute their jobs and which effort level to incentivize each agent. Each effort level induces a different expected transfer needed in order to incentivize it. As a first step, in Lemma 2.1, we reduce the Knapsack Contract problem to a variant of Stochastic Knapsack that we term the Multi-Choice Stochastic Knapsack Problem with Costs (MSKC). In this purely algorithmic problem, each job is assigned one of multiple combinations of cost and size distribution. When a policy attempts to execute a job, it has to pay the cost up front.
Our algorithm formulates an LP (see Eq. (2)) for the MSKC problem, where the weight of each job $i$ and effort level $j$ corresponds to the profit of executing job $i$ at effort level $j$, and the size is the expected size of the job at effort level $j$. In Dean et al. [11], an LP formulation of the problem gives an upper bound on the reward of the optimal adaptive policy, which is approximated by the reward of the best non-adaptive policy. In our case, the LP is still an upper bound on the optimal profit of an adaptive policy, but this upper bound can be a factor $\alpha$ away (see Appendix A for an example). We show that there exists a non-adaptive set of contracts that yields an $O(\alpha)$-approximation to the LP’s value.
The way in which we construct the set of contracts is the following. We first efficiently find a well-structured solution to the LP where each job is taken integrally at up to one effort level, except for at most a single job that can be taken as a convex combination of two effort levels. This is done by taking an arbitrary solution of the LP, and then shifting weights along the Pareto curves of jobs in order to iteratively move jobs to be taken integrally, while keeping the same solution value. Once we obtain the well-structured solution, we either use a standard bang-per-buck ordering of the jobs selected, or take a single job if its expected profit is larger than a predefined threshold. This threshold is set in a way that will (i) prevent a large loss for the bang-per-buck order in case the IOR parameter $\alpha$ has a large value; and (ii) ensure a large profit when there’s a single job that gives such profit, but might have a low probability of being selected when jobs are executed in the bang-per-buck order.
Negative results.
In Section 4 and Section 5 we give further evidence for the role that IOR has in Knapsack Contracts. In Section 4 we inspect the adaptivity gap, considering three types of policies:
  1. 1.
    Fully adaptive policies that can choose the next job to execute (if any) and at which effort level based on the entire history.
  2. 2.
    Stopping-time adaptive policies that chooses a fixed set of contracts (job/effort level pairs) up front at a fixed order, but stops once any future sequence of jobs has a negative expected profit. These types of policies avoid a large loss that the principal can incur due to the transfer component of the contract problem and are used in the algorithm of Section 6.
  3. 3.
    Non-adaptive policies that chooses a fixed set of contracts at a fixed order, and the process terminates either when the last job in the order has been executed, or when there’s a job that is finished after the deadline.
Dean et al. [11] show that in the setting without costs, the reward of the optimal non-adaptive policy is within a factor 4 of the optimal adaptive policy. We show that when introducing costs, the ratio between these can be unbounded. Namely, for every IOR parameter $\alpha>1$, we show:
  • An instance where the ratio between the profit of stopping-time adaptive policies and non-adaptive policies is $\Omega(\alpha)$ (Section 4.1).
  • An instance where the ratio between the profit of fully adaptive policies and stopping-time adaptive policies is $\Omega(\alpha^{1/3})$ (Section 4.2).
Our instances are ones where each job has only a single effort level, demonstrating that the cost aspect by itself leads to a significant change relative to the setting without cost.
Information barrier.
In Section 5 we inspect how much information is needed to get an $O(\alpha)$ approximation in the purely algorithmic Stochastic Knapsack problem where there is an inherent cost for executing jobs (this corresponds to having a single effort level for each agent). In the purely algorithmic setting, our algorithm from Section 3 only uses first moment information (the expectation) of jobs’ sizes, along with the probability each job’s size will exceed the deadline, and obtains an $O(\alpha)$ approximation. We show that this information does not suffice if one wants to obtain a better approximation. Namely, for every finite $k$, there exists an instance where any (possibly adaptive) policy that can only access the $k$ moments of the size distribution, along with the probability that each job will exceed the deadline, is bounded away by a factor of $\Omega(\alpha)$ from the optimal policy (Theorem 5.2).
Bounded-size jobs.
In Section 6, we consider the case where the IOR parameter can be arbitrarily large, but each job’s probability to take more than $1-\delta$ of the length of the time budget is small for $\delta\in(0,1)$. We give an efficient algorithm that finds a stopping-time adaptive set of contracts, which guarantees an $O(1/\delta)$ approximation to the optimal profit of a fully-adaptive set of contracts.
In order to obtain this result, we first consider a variant of the problem where, when a job finishes after the deadline, the principal still obtains its reward, and then the process stops. For this variant, the algorithm from Section 3 gives a constant approximation to the LP value. We then consider running this algorithm while imagining the deadline is $\delta$-fraction of the actual deadline. We show that in this case, having bounded-size jobs is equivalent to obtaining the reward of the first job that finishes after the deadline. Moreover, halving the deadline only causes a factor $\delta$ loss in the objective. Note that (moderate) adaptivity is required for this variant as well, as non-adaptive policies lose a factor that grows with $\alpha$ (see Appendix B).
Ordered-Adaptive Contracts.
Motivated by real-world examples where jobs must be executed in a given order, in Appendix D we study the model where the order is given, but for each job in the order, the principal can decide whether to execute it, and if so, at which effort level. In Theorem D.1 we give a resource augmentation-type result: given a time budget larger by a factor of $1+\epsilon$ from the original time budget, one can compute an adaptive policy efficiently (in polynomial time in the number of jobs, effort levels, and $1/\epsilon$), which obtains a reward larger than the optimal adaptive policy for the original time budget. The policy, similar to [11, 5], defines a dynamic programming solution on carefully discretized rounded-down processing times. This result hints that our problem boils down to ordering the jobs; i.e., one may hope to find a nearly-optimal adaptive policy for our non-ordered model if one can efficiently find a near-optimal ordering of the jobs.
1.3 Related Work
The study of sequential contracts in [14] investigates the problem of incentivizing a single agent to perform multiple jobs in a sequential manner. In contrast, our work considers a sequence of contracts in which the jobs are executed sequentially by different agents, with no interdependence between the jobs. The works of [26, 15, 1, 16] consider contract design subject to a limited monetary budget constraining the transfers from principal to agent. In our work, we study a multi-agent setting under a time budget constraint on the agents’ actions, where time is limited and the principal must incentivize the agents to complete their jobs on time.
Classic contract theory.
Classic contract theory has a long-standing tradition in economics, focusing on incentive design under asymmetric information. A result by Carroll [7] shows that linear contracts are often robust across a range of environments, providing a foundation for the study of simpler contract forms. These classical models emphasize moral hazard, adverse selection, and trade-offs between risk and incentives.
Algorithmic contract theory.
In recent years, a series of studies has emerged that study contracts through the lens of the theory of computation, initiated by [12, 4]. The authors of [12] raise concerns regarding the complexity and unintuitiveness of optimal contracts, while exploring the approximation ratios of simpler contract forms, such as linear contracts. The multi-agent setting was initially explored by Babaioff et al. [4]. A recent survey paper [13] provides an overview of developments in algorithmic contract theory.
2 Model
In Section 2.1 we present our contract setting; we reduce this setting to a purely combinatorial setting in Section 2.2.
2.1 Contract Design Setting
An instance $I$ of the knapsack contracts problem consists of a principal, $n$ agents, and a knapsack of size $S$. Without loss of generality, let $S=1$, where $S$ represents a budget of some limited resource. As our running example, we assume $S$ is the time remaining until a set deadline. Each agent $i$ has a job which they can carry out to the benefit of the principal. Agent $i$ is represented by the following:
  • $v_{i}$: The reward which the principal obtains if agent $i$ completes their job before the deadline.
  • $\{(F_{i}^{j},c_{i}^{j})\}_{j\in[m]}$: The possible effort levels of agent ii. [^1] All agents have the same number of effort levels without loss of generality, since we can add “dummy” effort levels that would never be incentivized by the principal. Each effort level $j\in[m]$ determines a distribution $F_{i}^{j}$ over the time it takes the agent to complete the job, as well as a cost $c_{i}^{j}\geq 0$. If the agent chooses to exert effort level $j$, they incur cost $c_{i}^{j}$, and complete their job within time $s_{i}$ drawn (independently) from distribution $F_{i}^{j}$. We assume non-decreasing costs $c_{i}^{1}\leq\dots\leq c_{i}^{m}$. Let $\mathbf{F}_{i}=\left(F_{i}^{1},\ldots,F_{i}^{m}\right)$. Let $\mathrm{supp}(F_{i}^{j})$ denote the support of $F_{i}^{j}$, and $\mathrm{supp}(\mathbf{F}_{i})=\cup_{j\in[m]}\mathrm{supp}(F_{i}^{j}).$ We assume the support of each distribution is finite and is explicitly given.
The principal knows $I=(v_{i},\{F_{i}^{j},c_{i}^{j}\}_{j\in[m]})_{i\in[n]}$ in advance. After an agent completes a job, the principal observes the time it took the agent to complete it, but it does not observe the effort level chosen by the agent. The principal offers each agent $i$ a contract $t_{i}:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$, mapping $i$’s observed processing time to a transfer paid by the principal to the agent. Given contract $t_{i}$, agent $i$ chooses an effort level
$j\in\arg\max_{\ell}\operatorname{\mathbb{E}}_{s_{i}\sim F_{i}^{\ell}}[t_{i}(s_{i})]-c_{i}^{\ell}.$
Definition 2.1 (Implementable effort level).
For agent $i$, an effort level $j$ is implementable if there exists a contract $t_{i}$ for which this action maximizes the agent’s expected utility among all actions given the contract. In this case, we say contract $t_{i}$ incentivizes action $j$.
Any implementable effort level $j$ for agent $i$ has an optimal contract that incentivizes it, which can be obtained by solving the following LP:
$\min\quad$ $\sum_{s\in\mathrm{supp}(F_{i}^{j})}t_{i}(s)\cdot F_{i}^{j}(s)$
s.t. $\sum_{s\in\mathrm{supp}(F_{i}^{j})}t_{i}(s)\cdot F_{i}^{j}(s)-c_{i}^{j}\geq\sum_{s\in\mathrm{supp}(F_{i}^{j^{\prime}})}t_{i}(s)\cdot F_{i}^{j^{\prime}}(s)-c_{i}^{j^{\prime}}$ $\forall j^{\prime}\in[m]$
$t_{i}(s)\geq t_{i}(s^{\prime})$ $\forall s<s^{\prime}\in\mathrm{supp}(\mathbf{F}_{i})$ (1)
$t_{i}(s)\geq 0$ $\forall s\in\mathrm{supp}(\mathbf{F}_{i})$
Constraint (1) ensures monotonicity of payments. This is required as without it, an agent might have an incentive to hover around after completing the job and get a higher payment.
The principal’s problem can then be described as: when hiring agent $i$, the principal chooses an effort level $j$ to incentivize, and offers the optimal contract that incentivizes that effort level. The principal’s collected reward is the sum of the rewards of the jobs that the agents get to complete with a total processing time less than the budget $S$. The principal aims to maximize their profit: the expected total reward minus the expected sum of transfers made to the agents.
Adaptive versus non-adaptive policies.
Definition 2.2.
(Fully-Adaptive Policies). A fully-adaptive policy, $\varphi:2^{I}\times[0,1]\rightarrow[n]\times[m]\cup\{\perp\}$, takes as input the unassigned jobs $J\subseteq I$ and the remaining time budget $S^{\prime}\leq 1$, and chooses which agent to hire next and at which implementable effort level. This procedure is repeated until:
  • an agent finishes their job after the deadline; or
  • the policy decides not to execute another job (outputs $\perp$ as the next job).
Let $\mathrm{profit}(\varphi)$ be a random variable representing the profit obtained by the principal when using policy $\varphi$ on instance $I$. We denote by $\mathrm{ADAPT}(I)=\max_{\varphi}\operatorname{\mathbb{E}}[\mathrm{profit}(\varphi)]$ the optimum expected profit obtained by an adaptive policy for instance $I$. When referring to an algorithm $\mathcal{A}$ that is given an instance $I$ and returns a policy $\varphi_{I}$, we denote by $\mathcal{A}(I)=\mathrm{profit}(\varphi_{I})$ the random variable representing the profit obtained by the principal when using the policy outputted by algorithm $\mathcal{A}$’s policy on instance $I$.
Definition 2.3.
(Non-Adaptive Policies) A non-adaptive policy is an ordering of agents $\pi:[n]\rightarrow[n]\cup\{\perp\}\times[m]$. For some $i\in[n]$, $\pi(i)$ outputs the $i$th selected agent and the effort level the policy incentivizes the agent to exert.
Let $\Pi$ be the set of non-adaptive policies. We denote by $\mathrm{NON\_ADAPT}(I)=\max_{\pi\in\Pi}\operatorname{\mathbb{E}}[\mathrm{profit}(\pi)]$ the optimum expected profit obtained by a nonadaptive policy for instance $I$.
Definition 2.4.
(Stopping-Time Adaptive Policies) Similarly to non-adaptive policies, a stopping-time adaptive policy chooses a non-adaptive ordering of the agents and their effort levels $\pi$. However, it may adaptively choose to stop at any point. Without loss of generality, we assume the policy stops if every remaining sequence of agents yields a negative expected profit.
Let $\Pi^{\prime}$ be the set of stopping-time adaptive policies. $\mathrm{STOP\_ADAPT}(I)=\max_{\pi\in\Pi^{\prime}}\operatorname{\mathbb{E}}[\mathrm{profit}(\pi)]$ is the optimum expected profit obtained by a stopping-time policy for instance $I$.
2.2 Multi-Choice Stochastic Knapsack Problem With Costs
We define a generalization of Stochastic Knapsack, where items have multiple choices of assignment, and each is associated with a size distribution, and a cost. We show that our Knapsack Contracts problem reduces to this problem. This implies that our algorithmic results can be obtained in this setting, and the results will carry through to the contracts setting.
Definition 2.5 (Multi-Choice Stochastic Knapsack Problem With Costs).
An instance of the multi-choice stochastic knapsack problem with costs (or MSKC for short) consists of $n$ items, where each item has $m$ choices, each choice has a different cost and size distribution. That is $I=\{v_{i},\{p_{i}^{j},F_{i}^{j}\}_{j\in[m]}\}_{i\in[n]}.$ An item $i$ can be selected with a single choice $j$. In this case, the cost $p_{i}^{j}$ must be incurred up front, and the item’s size is realized from distribution $F_{i}^{j}$. There is a budget of size $S$ ($S=1$ without loss of generality). A policy selects item-choice pairs sequentially, and terminates either when it decides not to select any item, or when the total size of selected items exceeds the budget. The profit of a policy is the total value of items selected that do not exceed the budget minus the total cost of the selected items, including the cost of an item that exceeds the budget, if one exists.
Adaptive, Stopping-time Adaptive and Non-adaptive policies for the MSKC problem are defined analogously to the way they are defined for the Knapsack Contracts problem. The following lemma (Proof in Appendix F) shows that in order to find an approximately optimal policy for the Knapsack Contracts problem, it suffices to find an approximately optimal policy for the MSKC problem.
Lemma 2.1 (Reduction to Multi-Choice Stochastic Knapsack with Costs).
Given an efficient algorithm that finds a policy that $c$-approximates the optimal policy MSKC problem, one can efficiently find a policy that $c$-approximates the optimal policy for the knapsack contracts problem.
2.3 Inverse of Return on Investment Parameter
Consider an instance of the Knapsack Contracts problem. For a job and implementable effort level pair, $(i,j)$, let
$\tilde{t}_{i}^{j}=\operatorname{\mathbb{E}}_{s\sim F_{i}^{j}}[t_{i}(s)],$
where $t_{i}$ is the contract that minimizes the expected payment while incentivizing effort level $j$, as given by the LP in Definition 2.1. That is, $\tilde{t}_{i}^{j}$ is the minimal expected transfer required in order to incentivize agent $i$ to take action $j$.
The hardness results presented in the paper critically exploit problem instances characterized by pairs (job, action) with high values $v_{i}$ and expected transfers $\tilde{t}_{i}^{j}$, but a comparatively low net profit
$v_{i}\cdot\Pr_{s_{i}^{j}\sim F_{i}^{j}}[s_{i}^{j}\leq 1]-\tilde{t}_{i}^{j}.$
In such cases, a job may yield a small positive profit of $v_{i}-\tilde{t}_{i}^{j}$ when the knapsack is empty, yet incur a large negative penalty of up to $\tilde{t}_{i}^{j}$ when capacity is low. Moreover, this fluctuation can be abrupt: even an $\epsilon$ increase in load may decrease the expected profit to $\tilde{t}_{i}^{j}$. We show that this drastic fluctuation property is an inherent obstacle in our model. To do so, we define the Inverse of the Return on Investment parameter:
Definition 2.6 (Inverse of ROI parameter).
We define
$\mathrm{ROI}(i,j)\ =\ \frac{v_{i}\cdot\Pr_{s_{i}^{j}\sim F_{i}^{j}}[\,s_{i}^{j}\leq 1\,]\;-\;\tilde{t}_{i}^{j}}{\tilde{t}_{i}^{j}},$
and
$\alpha\;=\;\max_{i,j\in\mathsf{I}_{i}}\frac{1}{\mathrm{ROI}(i,j)}\,$,
where $\mathsf{I}_{i}$ is the set of implementable actions for job $i$. We sometimes refer to $\alpha$ as the Inverse of ROI (or IOR) parameter.
Intuitively, a large value of $\alpha$ signifies the presence of at least one job where for some effort level, the transfer required to incentivize it is much larger than the net profit from successfully executing the job before the deadline. We show a $\Theta(\alpha)$ adaptivity gap and an $\Omega(\alpha)$ information gap (Section 5), suggesting that $\alpha$ captures precisely the “hardness” aspect for our problem.
Analogously, we define the IOR parameter for the MSKC problem as in Definition 2.6, but using $p_{i}^{j}$ instead of $\tilde{t}_{i}^{j}$. Given a Knapsack Contracts instance with an $\alpha$ IOR parameter, applying the reduction in Lemma 2.1 yields an MSKC instance with the same $\alpha$ as the analogous IOR parameter. This implies that any approximation as a function of $\alpha$ a policy in the MSKC instance would obtain will imply the same approximation as a function of $\alpha$ in the Stochastic Knapsack instance.
3 $O(\alpha)$-Approximate Non-Adaptive Contracts
In this section, we show how to find a non-adaptive set of contracts $\pi:[n]\to[n]\times[m]\cup\{\perp\}$ that obtains an $O(\alpha)$ approximation to the optimal profit the principal can obtain by any (possibly highly adaptive) policy.
By Lemma 2.1, we focus on finding an $O(\alpha)$-approximate non-adaptive policy for the MSKC problem, which yields an $O(\alpha)$-approximate non-adaptive set of contracts for the Knapsack Contracts problem. Our main theorem is the following:
Theorem 3.1.
There exists a polynomial time algorithm for MSKC that computes for every input $I$ a non-adaptive policy $\pi_{I}:[n]\to[n]\times[m]\cup\{\perp\}$ such that
$\frac{\mathrm{ADAPT}(I)}{\operatorname{\mathbb{E}}[\mathrm{profit}(\pi_{I})]}\leq O(\alpha).$
We first define an LP that both guides our solution and gives an upper bound on the performance of every adaptive policy. For every $i\in[n],j\in[m]$ and for a knapsack of size $t$, denote by $w_{ij}=v_{i}\cdot\Pr_{{s_{i}^{j}\sim F_{i}^{j}}}[s_{i}^{j}\leq t]-p_{i}^{j}$ and $\mu_{ij}=\operatorname{\mathbb{E}}_{s_{i}^{j}\sim F_{i}^{j}}[\min\{s_{i}^{j},t\}]$. The first parameter can be viewed as a proxy to the profit from job $i$ when executed at effort level $j$ and the second parameter is the expected truncated size.
Consider the following LP, parameterized by the size of the knapsack $t$.
$\Phi(t)=\max\left\{\sum_{ij}w_{ij}\,x_{ij}\;:\;\sum_{ij}\mu_{ij}\,x_{ij}\,\leq\,t,\quad\forall\,i,\;\sum_{j}x_{ij}\,=\,1,\quad x_{ij}\,\geq\,0\right\}.$ (2)
We assume, without loss of generality, that every item $i$ includes a dummy “null” choice $j_{\bot}$ with zero size and zero value; selecting this choice represents the case where none of the actual choices for the item are chosen. In addition, we will use the solution to this LP to guide our algorithm. First, we characterize the structure of optimal solutions:
Lemma 3.2.
There exists an optimal solution to $\Phi(t)$ satisfying the following properties:
  1. 1.
    At most one item $i^{*}$ has exactly two choices $j,j^{\prime}$ such that $x_{i^{*}j},x_{i^{*}j^{\prime}}>0$.
  2. 2.
    Every other item $i\neq i^{*}$ has at most one choice $j$ with $x_{ij}=1$.
Moreover, such a solution can be computed in polynomial time.
Note that every solution returned by the Simplex algorithm satisfies the above properties; the basic feasible solution of the LP has at most $n+1$ non-negative variables, and the constraints ensure that for each item $i$, there is at least one choice $j$ with $x_{ij}>0$, hence at most one item has two non-zero choices. However, the Simplex algorithm may run at exponential time in the worst case. Thus, we describe a polynomial time algorithm to compute such a solution.
First, for each item $i$, we characterize the choices $j$ for which there exists an optimal solution with $x_{ij}=0$. We then show that there always exists an optimal solution in which at most two choices have positive values.
Definition 3.1 (Single Dominance).
For an item $i$, choice $j^{\prime}$ single dominates choice $j$ if:
$(1)$ $\quad w_{ij^{\prime}}\geq w_{ij}$
$(2)$ $\quad\mu_{ij^{\prime}}\leq\mu_{ij}$
Definition 3.2 (Double Dominance).
For an item $i$, choices $j^{\prime}$ and $j^{\prime\prime}$ double dominate choice $j$ if:
$(1)$ $\quad w_{ij^{\prime}}<w_{ij}<w_{ij^{\prime\prime}}$
$(2)$ $\quad\mu_{ij^{\prime}}\leq\mu_{ij}\leq\mu_{ij^{\prime\prime}}$
$(3)$ $\quad(\mu_{ij}-\mu_{ij^{\prime}})/(w_{ij}-w_{ij^{\prime}})\geq(\mu_{ij^{\prime\prime}}-\mu_{ij})/(w_{ij^{\prime\prime}}-w_{ij})$
Finally, we say that a choice $j$ is dominated if it is either single-dominated or double-dominated.
Lemma 3.3.
There exists an optimal solution $x$ to $\Phi(t)$ in which every dominated choice is assigned zero weight; namely, for any item $i$ and any choice $j$ that is either single‑dominated or double‑dominated, we have $x_{ij}=0$.
Proof.
Suppose that for some feasible solution $x$, we have $x_{ij}\neq 0$.
First, if $j^{\prime}$ single-dominates $j$, define a new solution by setting
$x^{*}_{ij^{\prime}}=x_{ij^{\prime}}+x_{ij},\quad x^{*}_{ij}=0~.$
This yields a feasible solution whose objective value is at least as large as that of $x$. Hence, there exists an optimal solution in which $x_{ij}=0$.
Second, if $j^{\prime}$ and $j^{\prime\prime}$ double-dominate $j$, define a new solution by setting
$x^{*}_{ij^{\prime}}=x_{ij^{\prime}}\;+\;x_{ij}\cdot\frac{w_{ij^{\prime\prime}}-w_{ij}}{w_{ij^{\prime\prime}}-w_{ij^{\prime}}},\quad x^{*}_{ij^{\prime\prime}}=x_{ij^{\prime\prime}}\;+\;x_{ij}\cdot\frac{w_{ij}-w_{ij^{\prime}}}{w_{ij^{\prime\prime}}-w_{ij^{\prime}}},\quad x^{*}_{ij}=0~.$
This again yields a feasible solution whose objective value is at least as large as that of $x$ (see Fig. 1 for illustration, and Appendix E for the full proof). Therefore, in either case, there exists an optimal solution in which $x_{ij}=0$.
[…]
Figure 1: Visualization of the removal of double-dominated points. The hollow circles ($j^{\prime},j^{\prime\prime}$) represent two choices for an item. The red point ($j$) is ”double-dominated” because it lies above the line segment representing the convex combination of $j^{\prime}$ and $j^{\prime\prime}$. In the LP relaxation, this dominated point can be replaced by its vertical projection (the green point) onto the segment, which yields a solution that is no worse in objective value and potentially better in terms of feasibility ($\mu$).
Lemma 3.4.
There exists an optimal solution $x$ to $\Phi(t)$ in which, for every item $i$, at most two choices satisfy $x_{ij}>0$.
Proof.
Let $x$ be an optimal solution in which no choice is dominated; such a solution exists by Lemma 3.3. Without loss of generality we may further assume that no additional choices remain and that
$w_{i,j}\;<\;w_{i,j+1}\qquad\text{for every item }i\text{ and every }j\in\{1,\dots,m-1\}.$
Lemma 3.3 then also yields
$\mu_{i,j}\;<\;\mu_{i,j+1}\qquad\text{for all }i\text{ and }j\in\{1,\dots,m-1\}.$
Consider the pairs $\{\bigl{(}w_{ij},\mu_{ij}\bigr{)}\}_{j\in[m]}$ for a fixed index $i$. Connecting consecutive points by straight-line segments (the standard piecewise-linear interpolation) yields a polyline with non-decreasing slopes. Consequently, the boundary is convex.
Suppose for some item $i$ there are strictly more than two choices that satisfy $x_{ij}>0$. Define $\tilde{\mu}=\sum_{j}x_{ij}\mu_{ij}$. Let $j^{\prime}=\arg\max_{j}\bigl{\{}\mu_{ij}\leq\tilde{\mu}\bigr{\}}$ and $j^{\prime\prime}=\arg\min_{j}\bigl{\{}\mu_{ij}\geq\tilde{\mu}\bigr{\}}$. By the definition of $\tilde{\mu}$ and by our assumption that at least two choices satisfy $x_{ij}>0$, such indices $j^{\prime}$ and $j^{\prime\prime}$ always exist. Set $\lambda=\frac{\mu_{ij^{\prime\prime}}-\tilde{\mu}}{\mu_{ij^{\prime\prime}}-\mu_{ij^{\prime}}}$. Consider the solution defined by $x^{*}_{ij^{\prime}}=\lambda$, $x^{*}_{ij^{\prime\prime}}=1-\lambda$, and $x^{*}_{ij}=0$ for all $j\notin\{j^{\prime}j^{\prime\prime}\}$. Note that this construction satisfies $\sum_{j}x^{*}_{ij}\,\mu_{ij}=\tilde{\mu}$. Therefore, it yields a feasible solution. Since the resulting point lies on the boundary of the convex set, its objective value is at least as large as that of the original solution $x$, (see Fig. 2 for illustration).
[…]
Figure 2: Visualization of the set of undominated points, $\Gamma_{i}$. The blue line represents the efficient (Pareto) frontier formed by undominated choices and their convex combinations. An arbitrary combination of choices that lies above this frontier (the red point) is inefficient. It can be strictly improved by projecting it horizontally onto the frontier (the green point), which achieves the same $\mu$ value for a greater $w$ value.
We are now ready to prove Lemma 3.2.
Proof of Lemma 3.2.
By Lemmas 3.3, 3.4, it follows that there exists an optimal solution $x$ where at most two choices have positive values.
Consider a solution where there exist indices $i_{1},i_{2}$ and $j^{\prime}_{1},j^{\prime\prime}_{1},j^{\prime}_{2},j^{\prime\prime}_{2}$ such that $x_{i_{k},j^{\prime}_{k}}>0$ and $x_{i_{k},j^{\prime\prime}_{k}}>0$ for each $k\in\{1,2\}$. Let
$d_{k}=\frac{w_{i_{k},j^{\prime}_{k}}-w_{i_{k},j^{\prime\prime}_{k}}}{\mu_{i_{k},j^{\prime}_{k}}-\mu_{i_{k},j^{\prime\prime}_{k}}}$
be the current slope of the $k$-th item, and assume without loss of generality that $d_{1}\geq d_{2}$. Define
$\tau=\frac{\mu_{i_{1},j^{\prime}_{1}}-\mu_{i_{1},j^{\prime\prime}_{1}}}{\mu_{i_{2},j^{\prime}_{2}}-\mu_{i_{2},j^{\prime\prime}_{2}}}.$
For $\epsilon>0$, set
$x^{*}_{i_{1},j^{\prime}_{1}}=x_{i_{1},j^{\prime}_{1}}+\epsilon,\quad x^{*}_{i_{1},j^{\prime\prime}_{1}}=x_{i_{1},j^{\prime\prime}_{1}}-\epsilon,$
$x^{*}_{i_{2},j^{\prime}_{2}}=x_{i_{2},j^{\prime}_{2}}-\tau\cdot\epsilon,\quad x^{*}_{i_{2},j^{\prime\prime}_{2}}=x_{i_{2},j^{\prime\prime}_{2}}+\tau\cdot\epsilon.$
By construction, for any $\epsilon>0$ we have
$\sum_{j}x_{i_{1},j}\,\mu_{i_{1},j}+\sum_{j}x_{i_{2},j}\,\mu_{i_{2},j}=\sum_{j}x^{*}_{i_{1},j}\,\mu_{i_{1},j}+\sum_{j}x^{*}_{i_{2},j}\,\mu_{i_{2},j}.$
By choosing the smallest $\epsilon$ for which $x^{*}_{i_{1},j^{\prime}_{1}}=1$, or $x^{*}_{i_{2},j^{\prime}_{2}}=0$, or $x^{*}_{i_{1},j^{\prime\prime}_{1}}=0$, or $x^{*}_{i_{2},j^{\prime\prime}_{2}}=1$, we obtain a feasible solution whose objective value is at least as large as that of the original solution $x$, and whose number of items with two positive choices is strictly smaller. By iteratively applying this procedure, we eventually obtain a solution in which at most a single item has two positive choices.
Finally, we observe that all these transformations are applicable in polynomial time. Therefore, given any solution to $\Phi(t)$, it is possible to maintain a solution satisfying the above properties in polynomial time.
[…]
Figure 3: For an optimal solution $x$, items $i_{1}$ and $i_{2}$ each have two choices with nonzero values. The solution for each item is a convex combination of the two choices, corresponding to the red dots in the graphs $\Gamma_{i_{1}}$ and $\Gamma_{i_{2}}$, which lie on the segments connecting the two choices. As the slope of the first item is larger, the modified solution indicated by the dashed arrows illustrates one possible update that maintain feasibility and increases the objective value.
For a set of item choices $\mathcal{S}\subset[n]\times[m]$, denote $size(\mathcal{S})=\sum_{(i,j)\in\mathcal{S}}s_{i}^{j}$ and $\mu(S)=\sum_{(i,j)\in\mathcal{S}}\mu_{ij}$. The following Lemmas are direct extensions of key lemmas from Dean et al. [11], originally established for the classic Stochastic Knapsack problem with a budget of 1. We adapt these lemmas to our knapsack variant with multi-choice and costs, and for a general budget $t$. In this section, we only need the case where $t=1$; however, in Section 6 we use the generalized versions. The proofs are presented in Appendix E for completeness.
Lemma 3.5.
For any set of item choices $\mathcal{S}$:
$\Pr[size(\mathcal{S})<t]\geq 1-\frac{1}{t}\mu(\mathcal{S}).$
Lemma 3.6.
For any adaptive policy for the MSKC problem with budget $t$, let $\mathcal{S}$ be the random set of item choices that the policy inserts, then $\operatorname{\mathbb{E}}[\mu(\mathcal{S})]\leq 2t$.
Lemma 3.7.
For any constant $c\geq 2$ and budget $t$,
$\mathrm{ADAPT}\leq\Phi(ct)\leq c\cdot\Phi(t)~,$
where $\mathrm{ADAPT}$ is the optimal adaptive policy for budget $t$.
We are now ready to present our algorithm, SKC (Stochastic Knapsack with Costs):
Input: Instance II with items $i\in[n]$, choices $j\in[m]$, values $v_{i}$, costs $p_{i}^{j}$, distributions $F_{i}^{j}$
Output: Non-adaptive policy $\pi$
// Step 1: LP Solution
Solve linear program $\Phi(1)$ to obtain optimal solution $x$ with the structure from Lemma 3.2;
Let $\mathcal{C}=\{(i,j)\mid x_{ij}>0\}$ be the candidate item-choice pairs;
// Step 2: Virtual Item Construction
if $\exists$ item $i^{*}$ with two choices $j,j^{\prime}$ such that $x_{i^{*},j},x_{i^{*},j^{\prime}}>0$ then
   Remove $(i^{*},j)$ and $(i^{*},j^{\prime})$ from $\mathcal{C}$;
   Create virtual item $(i^{*},j^{*})$ with:
 $w_{i^{*},j^{*}}\leftarrow w_{i^{*},j}x_{i^{*},j}+w_{i^{*},j^{\prime}}x_{i^{*},j^{\prime}}$
 $\mu_{i^{*},j^{*}}\leftarrow\mu_{i^{*},j}x_{i^{*},j}+\mu_{i^{*},j^{\prime}}x_{i^{*},j^{\prime}}$
   Add $(i^{*},j^{*})$ to $\mathcal{C}$;
 // Randomly select between original choices
   When attempting to insert $(i^{*},j^{*})$, choose $j$ with probability $x_{i^{*},j}$ and $j^{\prime}$ with probability $x_{i^{*},j^{\prime}}$;
end if
// Step 3: Item Insertion with Threshold Check
$w_{\max}\leftarrow\max_{(i,j)\in\mathcal{C}}w_{ij}$
$\mathcal{T}\leftarrow\frac{\Phi(1)}{4(1+\alpha)}$
if $w_{\max}>\mathcal{T}$ then
   Insert item $(i_{\max},j_{\max})=\arg\max_{(i,j)\in\mathcal{C}}w_{ij}$;
 return policy that executes only this item;
end if
else
   Sort items in $\mathcal{C}$ by decreasing density $\frac{w_{ij}}{\mu_{ij}}$;
 return policy that executes items in decreasing density order;
end if
Algorithm 1 SKC (Stochastic Knapsack with Costs)
An overview of the algorithm.
We begin by solving $\Phi(1)$ to obtain an optimal solution $x$ with the properties established in Lemma 3.2: each item $i$ has one choice $j$ such that $x_{ij}=1$, except for possibly one item $i^{*}$ with two choices $j,j^{\prime}$ such that $x_{i^{*}j}+x_{i^{*}j^{\prime}}=1$. Let $\mathcal{C}$ denote the set of item-choice pairs selected by the LP, and let $r=|\mathcal{C}|$. If there exists an item with two positive choices, we remove these two choices and replace them with a single virtual choice; whenever the algorithm attempts to insert this virtual item, it randomly selects between the two original choices with probabilities proportional to their values in $x$. Since each item now has a single (possibly virtual) choice, we refer to each item-choice pair $(i,j)$ by the item $i$. Note that from the properties of our solution $x$, it holds that $\sum_{k=1}^{r}{w_{k}}=\Phi(1)$ and $\sum_{k=1}^{r}{\mu_{k}}\leq 1$. We then define a threshold $\mathcal{T}(\alpha)$ according to which we decide whether to insert the item with the maximum profit $w_{\text{max}}$. If the threshold is not met, we insert the items in decreasing order of their density $\frac{w_{i}}{\mu_{i}}$, until the knapsack is full or no more items remain in the candidate set $\mathcal{C}$. Our threshold determines whether or not following the greedy ordering is worth the risk of overflow compared to the alternative of inserting the best item. The bigger $w_{\text{max}},\alpha$ are, the more likely we are to insert the best item.
Note that this algorithm can be easily de-randomized by choosing the choice among the two choices that maximizes the expected gain.
Claim 3.8.
$\mathbb{E}[\textbf{SKC}]\geq\mathcal{T}=\frac{\Phi(1)}{4(1+\alpha)}$
Proof.
First, if $w_{\max}>\mathcal{T}$, the algorithm attempts only item $(i_{\max},j_{\max})$, gaining an expected net reward of $w_{\max}$. Since $w_{\max}>\mathcal{T}$, we have $\mathbb{E}[\textbf{SKC}]=w_{\max}>\mathcal{T}$, as required.
Now, consider the case $w_{\max}\leq\mathcal{T}$. Define $\textbf{SKC}_{k}$ as the contribution of item $k$ to the algorithm’s profit. It holds that
$\operatorname{\mathbb{E}}[\textbf{SKC}]=\sum_{k=1}^{r}{\operatorname{\mathbb{E}}[\textbf{SKC}_{k}]}~.$
We want to lower bound $\operatorname{\mathbb{E}}[\textbf{SKC}_{k}]$. If item $k$ is inserted and does not overflow, we obtain a reward of $v_{k}-p_{k}\geq w_{k}$. On the other hand, if the item overflows, we pay a cost of $p_{k}$. By the definition of the IOR parameter (Definition 2.6), $\alpha\geq\frac{p_{k}}{v_{k}\Pr[s_{k}\leq 1]-p_{k}}=p_{k}/w_{k}$, which implies that an overflowing job incurs a cost of $p_{k}\leq\alpha w_{k}\leq\alpha w_{\max}$.
We get that
$\operatorname{\mathbb{E}}[\textbf{SKC}_{k}]$ $=(v_{k}-p_{k})\cdot\Pr[k\mbox{ successfully inserted}]-p_{k}\cdot\Pr[k\mbox{ overflows}]$
$\geq(v_{k}-p_{k})\left(1-\sum_{j=1}^{k}{\mu_{j}}\right)-p_{k}\cdot\Pr[k\mbox{ overflows}]$
$\geq w_{k}\left(1-\sum_{j=1}^{k}{\mu_{j}}\right)-\alpha\cdot w_{\max}\cdot\Pr[k\mbox{ overflows}],$ (3)
where the first inequality follows from Lemma 3.5 (with $t=1$), the second inequality follows from the definition of $w_{k}\leq v_{k}-p_{k}$ and the argument above that $p_{k}\leq\alpha\,w_{\max}$.
In the above, this bound is also true for the virtual item $i^{*}$: the item’s expected truncated size is $\mu_{i^{*}}$, and its expected profit conditioned on the event that it was successfully inserted is at least $w_{i^{*}}$.
Summing over all $k\in[r]$, we have that
$\operatorname{\mathbb{E}}[\textbf{SKC}]$ $\geq\sum_{k=1}^{r}{w_{k}\left(1-\sum_{j=1}^{k}{\mu_{j}}\right)}-\alpha w_{\max}\sum_{k=1}^{r}{\Pr[k\mbox{ overflows}]}$
$\geq\Phi(1)-\sum_{k=1}^{r}{w_{k}}\sum_{j=1}^{k}{\mu_{j}}-\alpha w_{\max}$ $(*)$
$=\Phi(1)-\sum_{1\leq j\leq k\leq r}{w_{k}\mu_{j}}-\alpha w_{\max}$
$=\Phi(1)-\sum_{1\leq j<k\leq r}{w_{k}\mu_{j}}-\sum_{k=1}^{r}{w_{k}\mu_{k}}-\alpha w_{\max}$
$\geq\Phi(1)-\frac{1}{2}\sum_{k,j=1}^{r}{w_{k}\mu_{j}}-\sum_{k=1}^{r}{w_{k}\mu_{k}}-\alpha w_{\max}$ $(**)$
$\geq\frac{\Phi(1)}{2}-\sum_{k=1}^{r}{w_{k}\mu_{k}}-\alpha w_{\max}$ $\sum_{k,j=1}^{r}{w_{k}\mu_{j}}\leq\Phi(1)$
$\geq\frac{\Phi(1)}{2}-w_{\max}\cdot\left(\sum_{k=1}^{r}{\mu_{k}}+\alpha\right)$
$\geq\frac{\Phi(1)}{2}-w_{\max}(1+\alpha)$ $\sum_{k=1}^{r}{\mu_{k}}\leq 1$
$\geq\frac{\Phi(1)}{4}\geq\frac{\Phi(1)}{4(1+\alpha)}=\mathcal{T}~.$ $w_{\max}\leq\mathcal{T}=\frac{\Phi(1)}{4(1+\alpha)}.$ (4)
In ()(*) we use the fact that $\sum_{k\in[r]}\Pr[k\mbox{ overflows}]\leq 1$ as at most one item overflows, and $\sum_{k=1}^{r}{w_{k}}=\Phi(1)$. In ()(**) we use the decreasing order of density: for every $j<k$ it holds that $w_{k}\mu_{j}\leq w_{j}\mu_{k}$, and thus $w_{k}\mu_{j}\leq\frac{1}{2}\left(w_{k}\mu_{j}+w_{j}\mu_{k}\right)$.
We are ready to prove the main theorem of this section:
Proof of Theorem 3.1.
From 3.8 and Lemma 3.7, we conclude:
$\frac{\mathrm{ADAPT}}{\operatorname{\mathbb{E}}[\textbf{SKC}]}\leq\frac{2\Phi(1)}{\frac{\Phi(1)}{4(1+\alpha)}}=8(1+\alpha)=O(\alpha)~.$
In our proof, we show that our algorithm finds a policy that approximates the LP value up to a factor $O(\alpha)$. In [11], a similar LP helped guide the solution as its value was a proxy (up to constant factors) of the optimal policy for the stochastic knapsack problem. In our case, this LP provides an upper bound to the optimal policy (see Lemma 3.7), but its value can be larger by a factor $\Omega(\alpha)$ than the profit of the optimal policy (see Appendix A). This shows that we cannot hope for a tighter approximation to the LP’s value.
In Section 4, we show that every non-adaptive policy cannot have a better approximation than $O(\alpha)$ to the optimal adaptive policy (Theorem 4.1).
4 Adaptivity Gaps
In this section we study the power of different policies. In Section 4.1, we start by comparing Non-Adaptive policies (Definition 2.3) to Stopping-time Adaptive policies (Definition 2.4). Theorem 4.1 shows that the profit of a principal that uses the latter can be higher by a factor of $\Omega(\alpha)$ (which is tight according to Theorem 3.1). In Section 4.2 we study the power of Fully-Adaptive policies (Definition 2.2), and show they are more powerful than policies that fix the order of jobs and only decide when to stop processing new jobs. Namely, the gap between Fully-Adaptive policies and Stopping-time Adaptive policies can be as high as $\Omega(\alpha^{1/3})$ (Theorem 4.2). We note that in the instances that give the lower bound, all jobs have a single choice, and therefore, in order to incentivize jobs, we need to pay their cost exactly. Specifically, when determining the IOR, we can substitute the expected payment in Definition 2.6 with the jobs’ costs.
4.1 Stopping-time Adaptive vs. Non-Adaptive $\Omega(\alpha)$ Gap
We show that there exists a $\Omega(\alpha)$ gap between the optimal Non-Adaptive policy and the optimal Stopping-time Adaptive policy. Recall that both models require the algorithm to choose an item ordering at the beginning; the stopping-time adaptive policy can then adaptively choose when to halt, while the non-adaptive policy must attempt all items in the chosen order.
Theorem 4.1.
For every $\alpha$ there exists an instance $I$ of the knapsack-contracts problem such that
$\frac{\mathrm{STOP\_ADAPT}(I)}{\mathrm{NON\_ADAPT}(I)}=\Omega(\alpha).$
Proof.
For a constant $\alpha$, the claim is trivially true. We prove the claim for every $\alpha>3$. We build an instance $I$ as follows. Consider some $\epsilon<\frac{1}{2}$ and any $\gamma\in(0,\tfrac{1}{2})$, where $\epsilon$ will be used to determine $\alpha$. Our instance has $n$ jobs for $n=\lceil\frac{1}{\epsilon}\rceil$. Each job has value $v=2$ and cost $c=2-\epsilon$.
The size distribution of job $i\in[n]$ is
$s_{i}\;=\;\begin{cases}a_{i},&\text{with probability }1-\epsilon,\\[4.0pt] b_{i},&\text{with probability }\epsilon,\end{cases}$
where
$a_{i}=\gamma^{\,n-i+1},\qquad b_{i}=1-\displaystyle\sum_{j=1}^{i-1}a_{j}$
By Definition 2.6, we have
$\alpha=\frac{2-\epsilon}{\epsilon}\;=\;\Theta\!\left(\frac{1}{\epsilon}\right).$
Note that for every $\alpha>3$, we can choose an $\epsilon<1/2$ in order to set the IOR of our instance to be $\alpha$.
Since for every $i$, $\sum_{j=1}^{i-1}a_{j}+b_{i}=1,$ a stopping-time adaptive policy can safely schedule the jobs sequentially until it encounters the first job $i$ whose realized size equals $b_{i}$, and then halt.
We show that since $b_{i}+a_{i^{\prime}}>1$ for every pair $i<i^{\prime}$, attempting to schedule job $i^{\prime}$ without considering the realization of earlier jobs’ size yields negative expected profit. Consequently, the optimal $\mathrm{NON\_ADAPT}$ policy executes at most a single job. Our proof combines these two facts to yield the desired bound.
Formally, consider a stopping‑time adaptive policy that executes jobs in increasing index order ($i=1,2,\ldots$) until it encounters the first job whose size realizes as $s_{i}=b_{i}$, and then stops. Let $X$ denote the number of jobs the policy executes. We have
$\operatorname{\mathbb{E}}[X]\;=\;\sum_{i=1}^{n}\Pr[X\geq i]\;=\;\sum_{i=1}^{n}(1-\epsilon)^{i-1}\;=\;\frac{1-(1-\epsilon)^{n}}{\epsilon}.$
Since $n\geq 1/\epsilon$, we get that $\operatorname{\mathbb{E}}[X]\geq\frac{1-e^{-1}}{\epsilon}$. Therefore the expected profit is at least
$\mathrm{STOP\_ADAPT}(I)\;=\;\operatorname{\mathbb{E}}[X]\,(v-c)\;\geq\;\frac{1-e^{-1}}{\epsilon}\,\cdot\,\epsilon\;=\;1-e^{-1}.$
Consider now a non‑adaptive policy. This policy fixes the number of jobs to execute, $k$, and an ordering $\pi=i_{1},i_{2},\ldots,i_{k}$. For each $\ell\in[k]$, let $\mathrm{profit}(\ell)$ denote the profit from the $\ell$‑th job in the sequence. Clearly, $\operatorname{\mathbb{E}}[\mathrm{profit}(1)]=\epsilon$ for every non‑empty sequence of jobs. We will show that for every sequence with $k>1$ we have $\operatorname{\mathbb{E}}[\mathrm{profit}(k)]<0$, which implies that it is always better to execute at most a single job, since we can iteratively remove jobs and only improve the expected profit of the policy.
Given a sequence $\pi=i_{1},i_{2},\ldots,i_{k}$ of jobs, consider first the case in which $i_{\ell}>i_{k}$ for some earlier job in the sequence ($\ell<k$). If $i_{k}$ is reached and its size realises as $s_{i_{k}}=b_{i_{k}}$, the job terminates after the deadline, because $s_{i_{\ell}}+s_{i_{k}}\geq a_{i_{\ell}}+b_{i_{k}}>1$. Hence
[…]
$\geq\Pr[i_{k}\text{ is reached}]\Bigl{(}\Pr[s_{i_{k}}=b_{i_{k}}](-c)+\Pr[s_{i_{k}}=a_{i_{k}}](v-c)\Bigr{)}$
$=\Pr[i_{k}\text{ is reached}]\bigl{(}\epsilon(-c)+(1-\epsilon)(v-c)\bigr{)}$
$=\Pr[i_{k}\text{ is reached}]\bigl{(}-2\epsilon+\epsilon^{2}+\epsilon-\epsilon^{2}\bigr{)}$
$=-\,\epsilon\,\Pr[i_{k}\text{ is reached}]\;<\;0.$
Now consider the case where $i_{\ell}<i_{k}$ for all $\ell<k$. Let $j=\arg\max_{\ell<k}i_{\ell}$, the largest‑index job among the first k1k-1 scheduled. Note that, $i_{j}<i_{k}$ by our assumption. We have that
$\operatorname{\mathbb{E}}[\mathrm{profit}(k)]$ $=$ $\operatorname{\mathbb{E}}[\mathrm{profit}(k)|i_{k}\mbox{ is reached}]\cdot\Pr[i_{k}\mbox{ is reached}]$
$=$ $\operatorname{\mathbb{E}}[\mathrm{profit}(k)|s_{i_{\ell}}=a_{i_{\ell}}\quad\forall\ell<k]\cdot\Pr[s_{i_{\ell}}=a_{i_{\ell}}\quad\forall\ell<k]$
$+\operatorname{\mathbb{E}}[\mathrm{profit}(k)|i_{k}\mbox{ is reached}\ \wedge\ \exists\ell<k:\ s_{i_{\ell}}=b_{i_{\ell}}]\cdot\Pr[i_{k}\mbox{ is reached}\ \wedge\ \exists\ell<k:\ s_{i_{\ell}}=b_{i_{\ell}}]$
$=$ $(v-c)\cdot(1-\epsilon)^{k-1}+(-c)\cdot\Pr[i_{k}\mbox{ is reached}\ \wedge\ \exists\ell<k:\ s_{i_{\ell}}=b_{i_{\ell}}]$
$=$ $\epsilon\cdot(1-\epsilon)^{k-1}+(-c)\cdot\Pr[i_{k}\mbox{ is reached}\ \wedge\ \exists\ell<k:\ s_{i_{\ell}}=b_{i_{\ell}}]$
$\leq$ $\epsilon\cdot(1-\epsilon)^{k-1}+(-c)\cdot\Pr[s_{i_{j}}=b_{i_{j}}\ \wedge\ (s_{i_{\ell}}=a_{i_{\ell}}\quad\forall\ell<k,\ell\neq j)]$
$=$ $\epsilon\cdot(1-\epsilon)^{k-1}+(-2+\epsilon)\cdot(1-\epsilon)^{k-2}\cdot\epsilon$
$=$ $\epsilon\cdot(1-\epsilon)^{k-2}(1-\epsilon-2+\epsilon)$
$=$ $-\epsilon\cdot(1-\epsilon)^{k-2}<0.$
In the above, the third equality follows since $i_{k}$ always overflows in case some $s_{i_{\ell}}=b_{i_{\ell}}$ for some $i_{\ell}<i_{k}$, as $b_{i_{\ell}}+a_{i_{k}}>1$. The first inequality follows from the fact that c-c is negative, and $\Pr[i_{k}\mbox{ is reached}\ \wedge\ \exists\ell<k:\ s_{i_{\ell}}=b_{i_{\ell}}]\geq\Pr[s_{i_{j}}=b_{i_{j}}\ \wedge\ (s_{i_{\ell}}=a_{i_{\ell}}\quad\forall\ell<k,\ell\neq j)]$ since for the event $s_{i_{j}}=b_{i_{j}}$ and for all other $\ell<k$, $s_{i_{\ell}}=a_{i_{\ell}},$ we have that $i_{k}$ is reached. Since
$\sum_{\ell<k,\ell\neq j}a_{i_{\ell}}\leq\sum_{\ell=1}^{i_{j}-1}a_{\ell}=1-b_{i_{j}},$
and there will be no overflowing job before $i_{k}$ is executed.
We conclude that the optimal non-adaptive policy only executes a single job, and obtains a profit of $\epsilon$, while the optimal stopping-time adaptive policy obtains a profit of 1, as argued above. Therefore,
$\frac{\mathrm{STOP\_ADAPT}(I)}{\mathrm{NON\_ADAPT}(I)}\geq\frac{1-e^{-1}}{\epsilon}=\Omega(\alpha).$
4.2 Fully-adaptive vs. Stopping-time Adaptive $\Omega(\alpha^{1/3})$ Gap
We further show that there is an $\Omega(\alpha^{1/3})$ gap between the optimal stopping–time adaptive policy and the optimal fully adaptive policy. Recall that the stopping-time adaptive model requires the algorithm to fix an ordering of the items at the outset, and it halts if all future sequences of jobs yield a negative expected profit.
Theorem 4.2.
For every $\alpha$ there exists an instance $I$ of the knapsack-contracts problem such that
$\frac{\mathrm{ADAPT}(I)}{\mathrm{STOP\_ADAPT}(I)}=\Omega\bigl{(}{\alpha}^{1/3}\bigr{)}.$
In order to prove the above theorem, we define an instance $I$ consisting of an infinite collection of jobs, each with with value $v=15/\epsilon$ and cost $c=15/\epsilon-\epsilon^{2}$. As in Section 4.1, $\epsilon$ is used to determine $\alpha$. Again, as the claim is trivial for bounded $\alpha$, we prove the claim for large values of $\alpha$.
The job set is partitioned into $1/\epsilon$ types of jobs $L_{1},\ldots,L_{1/\epsilon}$, each containing an infinite number of jobs. A job of type $L_{i}$ has the size distribution
Probability Size
$1-\epsilon-\epsilon^{2}$ 0
$\epsilon$ $a_{i}$
$\epsilon^{2}$ $b_{i}$
where
$a_{i}=\gamma^{\,n-i+1},\qquad b_{i}=1-\displaystyle\sum_{j=1}^{i-1}a_{j}$
for $\gamma<\tfrac{1}{2}$.
By the definition of the IOR parameter (Definition 2.6),
$\alpha=\frac{15/\epsilon-\epsilon^{2}}{\epsilon^{2}}=15/\epsilon^{3}-1=O\left(\frac{1}{\epsilon^{3}}\right)$
For every large enough $\alpha$, we can choose an appropriate $\epsilon$ that sets the IOR to be $\alpha$.
We first analyze the expected reward from a fully adaptive policy.
Lemma 4.3.
For the instance given above,
$\mathrm{ADAPT}(I)\;\geq\;\frac{1}{2e}.$
Proof.
Consider an adaptive policy that selects jobs by type $L_{1}$, then type $L_{2}$, and so on until type $L_{1/\epsilon}$, according to the following rule:
  • When selecting jobs of type $L_{i}$, the policy continues to select such jobs as long as $s_{i}=0$.
  • Once $s_{i}=a_{i}$, the policy moves on to select jobs of type $L_{i+1}$ if $i<1/\epsilon$.
  • If $s_{i}=b_{i}$, the policy terminates.
Notice that the execution order ensures each executed job finishes within the time budget, so the policy obtains a reward of exactly $v-c=\epsilon^{2}$ for every executed job. Let $X$ be a random variable for the number of jobs the adaptive policy completes. Since the profit from each job is $\epsilon^{2}$, the policy gets a reward of $\operatorname{\mathbb{E}}[X]\cdot\epsilon^{2}$. Let $X_{i}$ be the number of jobs of type $L_{i}$ completed by the policy. Notice that
[…]
By the description of the policy,
$\operatorname{\mathbb{E}}\bigl{[}X_{i}\mid\varphi\text{ executes jobs of type }L_{i}\bigr{]}=\frac{1}{\epsilon+\epsilon^{2}}.$
We also have
[…]
The second inequality follows because in order for the policy to reach jobs of type $L_{1/\epsilon}$, each time a job of type $L_{i}$ (for $i<1/\epsilon$) has its size realized at a value larger than 0, the policy moves on to jobs of type $L_{i+1}$ only if the job’s size is realized at $a_{i}$, which happens with probability $\frac{\epsilon}{\epsilon+\epsilon^{2}}$. Since this must occur at most $1/\epsilon$ times, the probability is at least
$\left(\frac{\epsilon}{\epsilon+\epsilon^{2}}\right)^{1/\epsilon}.$
Therefore,
$\operatorname{\mathbb{E}}[X_{i}]\;\geq\;\frac{1}{e(\epsilon+\epsilon^{2})}\;>\;\frac{1}{2e\epsilon},$
which implies
$\operatorname{\mathbb{E}}[X]=\sum_{i=1}^{1/\epsilon}\operatorname{\mathbb{E}}[X_{i}]\;>\;\frac{1}{2e\epsilon}\cdot\frac{1}{\epsilon}=\frac{1}{2e\epsilon^{2}}.$
This yields
$\mathrm{ADAPT}(I)\;>\;\epsilon^{2}\cdot\frac{1}{2e\epsilon^{2}}\;=\;\frac{1}{2e},$
as desired. ∎
We next bound the expected reward from a stopping-time adaptive policy.
Lemma 4.4.
For the instance given above,
$\mathrm{STOP\_ADAPT}(I)\;\leq\;4\epsilon.$
Proof.
Consider some sequence $\pi$ of jobs. Let $Y$ denote the following event: the policy is to process a job of type $L_{i}$, when some prior job of type $L_{i}$ was processed with a size realized to be larger than 0. We claim that once event $Y$ occurs, then for every sequence of future jobs, the expected profit is negative. Consider first the expected profit of only processing the next job, of type $L_{i}$. This is at most
$(1-\epsilon^{2})(v-c)+\epsilon^{2}(-c)=(1-\epsilon^{2})\cdot\epsilon^{2}-\epsilon^{2}(15/\epsilon-\epsilon^{2})=-15\epsilon+\epsilon^{2}<0,$
since if the next job’s size is realized to be $b_{i}$, then the job finishes after the deadline for sure.
As the profit of any job is at most $v-c=\epsilon^{2},$ this implies that for any sequence to have a positive profit, it has to be of length larger than $(15\epsilon-\epsilon^{2})/\epsilon^{2}$. Consider a partition of the sequence into blocks of size
$k=\frac{1}{\epsilon+\epsilon^{2}}+\frac{1}{\epsilon},$
If the sequence ends at a block of size smaller than $k$, we merge the last two blocks together to form a single block, meaning that the size of each block is between $k$ and $2k$. Denote by $r$ the number of blocks in the sequence. Consider some block $j\in[r]$, and let $n_{j}(i)$ denote the number of jobs of type $L_{i}$ it contains. From our observation, if $n_{j}(i)>1$, then the expected profit of processing the second job of type $L_{i}$ in the block is at most $15\epsilon+\epsilon^{2} \ZoharcommentThis$ is only true given $Y$, right? I think the proof might be more clear if it is partitioned (e.g. having a couple of clearly stated claims in the proof), in which case the expected profit of the entire block is at most
$15\epsilon+\epsilon^{2}+\epsilon^{2}(2k-1)<-15\epsilon+2\epsilon^{2}+4\epsilon=-11\epsilon+2\epsilon^{2}<0.$
Therefore, to have a positive expected profit, for every $i$, if $b_{j}(i)>1$, then the first $b_{j}(i)-1$ jobs of type $L_{i}$ must have size 0, and in total, at least
$\sum_{i=1}^{1/\epsilon}{\max\{n_{j}(i)-1,0\}}\geq k-\frac{1}{\epsilon}=\frac{1}{\epsilon+\epsilon^{2}}$
jobs must have size 0. This occurs with probability at most
$(1-(\epsilon+\epsilon^{2}))^{\frac{1}{\epsilon+\epsilon^{2}}}<\frac{1}{2}$~
in which case the expected profit of the block is at most $2k\epsilon^{2}\leq 4\epsilon$.
Else, with probability at least $1/2$, the expected profit of the block is at most $11\epsilon+2\epsilon^{2}$. Thus, the expected profit of a block bounded given by
$\frac{1}{2}(-11\epsilon+2\epsilon^{2})+4\epsilon=-\frac{3}{2}\epsilon+\epsilon^{2}<0.$
Thus, we get that once event $Y$ occurs, since every sequence of jobs has a negative profit for the policy, it will halt. Let $X$ be a random variable of the number of jobs processed before event $Y$ occurs. Since every job has profit at most $\epsilon^{2}$, we have that $\mathrm{STOP\_ADAPT}(I)\leq\operatorname{\mathbb{E}}[X]\epsilon^{2}$. We again break $\pi$ into blocks of size $k$. As analyzed above, for every block, we have that event $Y$ occurs with probability at least $1/2$; therefore, the expected number of blocks processed is bounded by 2, and the expected number of jobs processed is bounded by $2k$. Therefore,
$\mathrm{STOP\_ADAPT}(I)\leq\operatorname{\mathbb{E}}[X]\epsilon^{2}\leq 2k\epsilon^{2}\leq 4\epsilon,$
as desired. ∎
Combining Lemmas 4.3 and 4.4 yields Theorem 4.2.
5 $\Omega(\alpha)$ Informational Gap
In this section we establish a lower bound that exposes a fundamental information gap in the problem of Stochastic Knapsack with Costs. Previous constant-factor approximation algorithms for Stochastic Knapsack (without costs) have used limited distributional information, specifically the expected (or truncated) sizes and the probability of fitting within the current capacity. In Proposition 5.1 we demonstrate that this information is insufficient. We show that any algorithm, even one that finds an adaptive policy, which relies solely on such limited data, cannot achieve a constant-factor approximation in the presence of costs. More generally, we prove in Theorem 5.2 that even providing the first $k$ moments of each distribution for any constant $k$ is insufficient. This showcases the inherent difficulty of the added costs, implying that future constant-approximation algorithms will need to utilize almost the full distributional structure in novel ways.
5.1 Warm-Up: First Moment
Proposition 5.1.
Let $\mathcal{A}$ be any algorithm whose knowledge of the item size distributions is restricted to the expected values $\operatorname{\mathbb{E}}[s_{i}]$, the truncated expected values $\mu_{i}=\operatorname{\mathbb{E}}[\min\{s_{i},1\}]$, and oracle access to the function $q_{i}(B)=\Pr[s_{i}\leq B]$ for each $i\in[n]$, where $B$ is the current remaining knapsack capacity. For every $\alpha$ there exists an input instance $I$ on which the algorithm’s performance is bounded by:
$\frac{\mathrm{ADAPT}(I)}{\operatorname{\mathbb{E}}[\mathcal{A}(I)]}\geq\Omega(\alpha).$
The proof relies on constructing two types of items, one “good” and one “bad”. The optimal strategy would be to insert only the “good” items, and any algorithm that starts by inserting a “bad” item will result in a solution with a value that is only a $\frac{1}{\alpha}$-fraction of OPT. We design the size distributions to be indistinguishable to algorithm $\mathcal{A}$, forcing $\mathcal{A}$ to arbitrarily pick an item to insert into the knapsack. By creating an instance where the number of “bad” items far exceeds the number of “good” items, algorithm $\mathcal{A}$ will almost surely pick a “bad” item first, resulting in our lower bound.
Proof of Proposition 5.1.
Fix $\epsilon\in(0,1)$. The two item types are defined as follows:
Property Type 1 Item (“good”) Type 2 Item (“bad”)
Reward $v_{i}$ 22 22
Cost $p_{i}$ $2-\epsilon$ $2-\epsilon$
Size $s_{i}$ dist. $s_{1}=\begin{cases}0,&\text{w.p. }1-\epsilon,\\ 1,&\text{w.p. }\epsilon;\end{cases}$ $s_{2}=\begin{cases}\epsilon/2,&\text{w.p. }1-\frac{\epsilon}{2-\epsilon},\\ 1,&\text{w.p. }\frac{\epsilon}{2-\epsilon}.\end{cases}$
Note that in this instance the expected values and the truncated expected values are identical: $\operatorname{\mathbb{E}}[s_{1}]=\operatorname{\mathbb{E}}[s_{2}]=\mu_{1}=\mu_{2}=\epsilon$ and $q_{1}(1)=q_{2}(1)=1$, hence, algorithm $\mathcal{A}$ cannot distinguish between the two items. Inserting one item of type 2 results in a net profit of $\epsilon$, and reduces the remaining capacity $B$ by at least $\epsilon/2$. Inserting any further items (of either types) results in a negative expected profit:
$\operatorname{\mathbb{E}}[\text{inserting item of type $i$}\;|\;B\leq 1-\epsilon/2]$ $\leq\Pr[s_{i}\neq 1]\cdot v_{i}-p_{i}$
$\leq(1-\frac{\epsilon}{2-\epsilon})\cdot 2-(2-\epsilon)$
$=\epsilon(1-\frac{2}{2-\epsilon})\;<\;0.$
Now, consider an input instance $I$ with $n>\frac{1}{\epsilon}$ items of type 1 and $n^{2}-n$ items of type 2. The optimal policy will attempt to insert items of type 1, and stop if an item’s size instantiates $s_{i}=1$, or if no more items of type 1 remain. The expected number of items successfully inserted by OPT is $\frac{1}{\epsilon}$, each item contributing $\epsilon$ net profit, resulting in
$\mathrm{ADAPT}(I)\geq\frac{1}{\epsilon}\cdot\epsilon=1.$
Since $\mathcal{A}$’s only strategy is to randomly select an item to insert, with probability $1-\frac{1}{n}$ it begins by inserting a “bad” item. By taking $n\to\infty$, the expected profit of $\mathcal{A}$ approaches $\epsilon$:
$\lim_{n\to\infty}\operatorname{\mathbb{E}}[\mathcal{A}(I)]=\epsilon.$
Thus for a large enough $n$, the performance ratio of the algorithm can be bounded:
$\frac{\mathrm{ADAPT}(I)}{\operatorname{\mathbb{E}}[\mathcal{A}(I)]}\geq\frac{1}{2\epsilon}\geq\frac{1}{4}\alpha.$
5.2 $k$ Moments
We generalize the previously shown distributional informational gap to show that, even an algorithm with knowledge of the first $k$ (or truncated) moments of the distribution, has an $\Omega(\alpha)$ gap compared to an algorithm with knowledge of the exact distribution.
Theorem 5.2.
Fix $k\geq 1$. Let $\mathcal{A}$ be any algorithm whose knowledge of the item size distributions is restricted to the moment $\operatorname{\mathbb{E}}[s_{i}^{j}]$, and truncated moment $\mu_{i}^{j}=\operatorname{\mathbb{E}}[(\min\{s_{i},1\})^{j}]$ for every item $i\in[n]$ and moment $j\in[k]$, and oracle access to the function $q_{i}(B)=\Pr[s_{i}\leq B]$, where $B$ is the current remaining knapsack capacity. For every $\alpha$ there exists an input instance $I$ on which the algorithm’s performance is bounded by:
$\frac{\mathrm{ADAPT}(I)}{\operatorname{\mathbb{E}}[\mathcal{A}(I)]}\geq\Omega(\alpha).$
Similarly to the previous proof, we construct two types of items, one “good” and one “bad”, which are indistinguishable to $\mathcal{A}$ yet are critically different for the optimal algorithm.
Proof of Theorem 5.2.
Fix $\epsilon\in(0,1)$. Fix distinct sizes
$t_{j}:=1-2^{-j},\quad j=1,2,\dots,k,$
$t_{k+1}:=1.$
The two item types are defined as follows. For sufficiently small $\delta>0$ and for $p_{j}\in[0,1]$ to be determined later:
Property Type 1 Item (“good”) Type 2 Item (“bad”)
Reward $v_{i}$ 22 22
Cost $c_{i}$ $2-\epsilon$ $2-\epsilon$
Size $s_{i}$ dist. $s_{1}=\begin{cases}0,&\text{w.p. }1-2\epsilon\\ t_{j},&\text{w.p. }\tfrac{\epsilon}{k+1},\;j=1,\dots,k+1\\ 1,&\text{w.p. }\epsilon\end{cases}$ $s_{2}=\begin{cases}\delta,&\text{w.p. }1-2\epsilon\\ t_{j},&\text{w.p. }p_{j},\;j=1,\dots,k+1\\ 1,&\text{w.p. }\epsilon\end{cases}$
Note that since $\Pr[s_{i}\leq 1]=1$, the truncated moments are identical to the original moments. We now focus on finding the correct parameters such that the two distributions have the same moments; We seek $p_{j}$ with $\sum_{j}p_{j}=\epsilon$, satisfying for each $r=1,\dots,k$:
$\operatorname{\mathbb{E}}[s_{1}^{r}]=\operatorname{\mathbb{E}}[s_{2}^{r}].$
That is, for $r=1,\dots,k$:
$\sum_{j=1}^{k+1}\frac{\epsilon}{k+1}\,t_{j}^{r}+\epsilon$ $=(1-2\epsilon)\,\delta^{r}+\sum_{j=1}^{k+1}p_{j}\,t_{j}^{r}+\epsilon$
$\iff\sum_{j=1}^{k+1}t_{j}^{r}\,\left(\frac{\epsilon}{k+1}-p_{j}\right)$ $=(1-2\epsilon)\,\delta^{r}.$
Denote $x_{j}:=\frac{\epsilon}{k+1}-p_{j}$. Then $\sum_{j}x_{j}=0$ (since $\sum p_{j}=\epsilon$ and $\sum\frac{\epsilon}{k+1}=\epsilon$). We get the following system of equations:
$\sum_{j=1}^{k+1}x_{j}$ $=0;$ (5)
$\sum_{j=1}^{k+1}t_{j}^{r}\,x_{j}$ $=(1-2\epsilon)\,\delta^{r},\quad\text{for }r=1,\dots,k.$ (6)
We now rewrite (5) and (6) in matrix form $Ax=b$, where
$A=\begin{pmatrix}1&1&\cdots&1\\ t_{1}&t_{2}&\cdots&t_{k+1}\\ t_{1}^{2}&t_{2}^{2}&\cdots&t_{k+1}^{2}\\ \vdots&\vdots&&\vdots\\ t_{1}^{k}&t_{2}^{k}&\cdots&t_{k+1}^{k}\end{pmatrix},x=\begin{pmatrix}x_{1}\\ \vdots\\ x_{k+1}\end{pmatrix},b=\begin{pmatrix}0\\ (1-2\epsilon)\delta\\ \vdots\\ (1-2\epsilon)\delta^{k}\end{pmatrix}$.
Note that $A$ is a $(k+1)\times(k+1)$ matrix where the first row corresponds to the zero-sum constraint (5), and the subsequent $k$ rows correspond to the moment-matching conditions (6). This is a transposed Vandermonde matrix [17]. Since the $t_{j}$ are distinct, it yields a unique solution $x=A^{-1}b$. Hence
$p_{j}=\frac{\epsilon}{k+1}-x_{j}.$
It is left to prove that $p_{j}\geq 0$ to conclude that $s_{2}$ is indeed a valid distribution. Note that since $A$ and $A^{-1}$ are functions of $k$ (and the fixed $t_{j}$ values), then after fixing $k$, $x_{j}$ is a function of $\epsilon$ and $\delta$:
$x_{j}(\epsilon,\delta)=\sum_{r=1}^{k}(A^{-1})_{j,r+1}[(1-2\epsilon)\delta^{r}].$
For a fixed $\epsilon$, we have that $|x_{j}(\delta)|=O(\delta)$. Thus, for $p_{j}=\frac{\epsilon}{k+1}-x_{j}$ to be non-negative, we can choose $\delta$ sufficiently small such that $\frac{\epsilon}{k+1}$ dominates the $x_{j}$ terms, ensuring $p_{j}\geq 0$. Hence, the two distributions are indistinguishable to $\mathcal{A}$.
Inserting a single item of type 2 results in a net profit of $\epsilon$, and reduces the remaining capacity $B$ by at least $\delta$. Inserting any further items (of either types) results in a negative expected profit:
$\operatorname{\mathbb{E}}[\text{inserting item of type $i$}\;|\;B\leq 1-\delta]$ $\leq\Pr[s_{i}\neq 1]\cdot v_{i}-c_{i}$
$=(1-\epsilon)\cdot 2-(2-\epsilon)$
$=-\epsilon<0.$
We conclude the proof by considering an input instance $I$ with $n>\frac{1}{\epsilon}$ items of type 1 and $n^{2}-n$ items of type 2. The optimal policy does as well as the following policy: it will attempt to insert items of type 1, and stop if an item’s size instantiates $s_{i}\neq 0$, or if no more items of type 1 remain. The expected number of items successfully inserted by OPT is $\frac{1}{2\epsilon}$, each item contributing $\epsilon$ net profit, resulting in
$\mathrm{ADAPT}(I)\geq\frac{1}{2\epsilon}\cdot\epsilon=\frac{1}{2}.$
Since $\mathcal{A}$’s only strategy is to randomly select an item to insert, with probability $1-\frac{1}{n}$ it begins by inserting a “bad” item. By taking $n\to\infty$, the expected profit of $\mathcal{A}$ approaches $\epsilon$:
$\lim_{n\to\infty}\operatorname{\mathbb{E}}[\mathcal{A}(I)]=\epsilon.$
Thus for a large enough $n$, the performance ratio of the algorithm can be bounded:
$\frac{\mathrm{ADAPT}(I)}{\operatorname{\mathbb{E}}[\mathcal{A}(I)]}\geq\frac{\frac{1}{2}}{2\epsilon}\geq\frac{1}{8}\alpha.$
6 $O(1/\delta)$-approximation for Bounded-Size Jobs
In this section, we assume that all jobs are size-bounded with high probability; specifically, we assume that for some $\delta\in(0,1)$, for all $i,j$,
$\Pr_{s_{i}^{j}\sim F_{i}^{j}}[s_{i}^{j}>1-\delta]\leq\frac{1}{2n\alpha}.$
By Lemma 2.1, we focus on providing an approximately optimal policy for the MSKC problem.
As a first step, we analyze a modified problem, where there is no bound on the job sizes, and if a job exceeds the deadline, the principal collects the job’s reward, and then the process terminates. For this case, we give a policy that approximates the profit of the optimal policy that does not receive the reward of the job that overflows. We then show how to transform this policy into a policy that handles bounded-size jobs and does not collect the reward of overflowing jobs.
Our algorithm for the case where jobs can overflow, SKC-OF, runs SKC with budget $\delta$ and a different threshold $\mathcal{T}=\Phi(\delta)/4$. Note that, unlike the SKC algorithm, this threshold is independent of $\alpha$. We denote by $\textbf{SKC-OF}(t)$, the algorithm that runs SKC with a budget $t$, and a threshold $\mathcal{T}=\Phi(t)/4$. Specifically, $\textbf{SKC-OF}=\textbf{SKC-OF}(\delta)$. We have the following.
Lemma 6.1.
For any constant $c\geq 1$,
$\operatorname{\mathbb{E}}[\textbf{SKC-OF}(t)]\geq\frac{1}{c}\Phi(ct)/4.$
Proof.
In case $w_{\max}\geq\Phi(t)/4$, we only take $w_{\max}$ and obtain a profit of
$v_{i_{\max}}-p_{i_{\max}}^{j_{\max}}\geq w_{\max}\geq\Phi(t)/4\geq\frac{1}{c}\Phi(ct)/4,$
where the last inequality follows Lemma 3.7. Therefore, assume $w_{\max}<\Phi(t)/4$.
Recall that $\textbf{SKC-OF}_{k}$ is the profit of item $k$ in the SKC-OF algorithm, as defined in Section 3. In our case, the profit from item $k$ will be as large as in the MSKC instance, and will not incur a cost in case of an overflow. Therefore, by adapting Eq. (3), our bound on $\textbf{SKC-OF}_{k}$ will be
$\operatorname{\mathbb{E}}[\textbf{SKC-OF}_{k}]\geq w_{k}\left(1-\frac{1}{t}\sum_{j=1}^{k}{\mu_{j}}\right).$
We adapt the bound on SKC-OF from Eq. (4) accordingly to be
$\operatorname{\mathbb{E}}[\textbf{SKC-OF}]\geq\Phi(t)/2-w_{\max}\geq\Phi(t)/4\geq\frac{1}{c}\Phi(ct)/8,$ (7)
where the second inequality follows our assumption that $w_{\max}\leq\Phi(t)/4$ and the last inequality follows Lemma 3.7. ∎
We now turn to the bounded job size case. Our algorithm for this case, SKC-BOUND, will run SKC-OF, and pretend that the budget is $\delta$, i.e. it will terminate if the total size of jobs exceeds $\delta$.
Theorem 6.2.
$\frac{\mathrm{ADAPT}}{\operatorname{\mathbb{E}}[\textbf{SKC-BOUND}]}\leq O\left(\frac{1}{\delta}\right)$
Proof.
If all items are instantiated with sizes $s_{i}^{j}<1-\delta$, then since the policy terminates when we exceed the budget of $\delta$, the policy never overflows, achieving the expected value of SKC-OF. However, if $s_{i}^{j}>1-\delta$ for some $i,j$, the policy terminates and an overflow may occur, resulting in a loss of at most $\alpha w_{\max}$. We have that
$\operatorname{\mathbb{E}}[\textbf{SKC-BOUND}]$ $\geq\operatorname{\mathbb{E}}[\textbf{SKC-OF}]-\alpha w_{\max}\cdot\Pr[\exists\text{ item $i$ with }s_{i}^{j}>1-\delta]$
$\geq\Phi(\delta)/4-\alpha w_{\max}\cdot\sum_{i=1}^{n}\Pr[s_{i}^{j}>1-\delta]$
$\geq\Phi(\delta)/4-\alpha w_{\max}\cdot n\cdot\frac{1}{2n\alpha}$
$\geq\Phi(\delta)/4-\Phi(\delta)/8$
$=\Phi(\delta)/8.$
The theorem follows from Lemma 3.7,
$\frac{\mathrm{ADAPT}}{\operatorname{\mathbb{E}}[\textbf{SKC-BOUND}]}\leq\frac{\Phi(2)}{\Phi(\delta/8)}\leq\frac{\frac{16}{\delta}\Phi(\delta/8)}{\Phi(\delta/8)}=O\left(\frac{1}{\delta}\right)~.$
We note that SKC-BOUND determines a non-adaptive order of jobs-effort levels (as it applies SKC), but it applies an adaptive stopping rule, as it stops after the total realized sizes first exceed $\delta$. In Appendix B, we show that even for bounded-sized jobs, any non-adaptive policy is bounded away by a factor $\Omega(\sqrt{\alpha})$ from the optimal policy (Theorem B.1).
In Appendix C, we show that if indeed the principal gets the reward of any job that is initiated before the deadline (even if it finishes after the deadline), then we have a constant adaptivity gap Lemma C.1.
7 Conclusion and Future Directions
In this work, we initiate the study of contract design where rewards are fixed, but the principal incentivizes the agents to invest effort in finishing their jobs on time. We identify an economically meaningful Inverse of the Return-on-Investment parameter, $\alpha$, which captures many aspects of the problem’s complexity. Dean et al. [11] show a constant adaptivity gap, while we show that in our case, the adaptivity gap is $\Theta(\alpha)$. As the optimal adaptive contract is intractable, a remaining open problem is whether one can efficiently compute an adaptive contract whose profit is a good approximation to the optimal adaptive contract.
In Section 4.2, we consider a policy with limited adaptivity: the order is non-adaptive, but the policy stops processing jobs once it expects a negative profit. We show that for such policies, their profit can be a factor $\Omega(\alpha^{1/3})$ away from the optimal profit. As our algorithm in Section 3 guarantees an $O(\alpha)$-approximation, this leaves a gap between the lower and upper bound for such policies. It is also interesting to consider other limited-adaptivity policies, e.g., a policy that predetermines the agents’ order, but not their effort levels. Another direction is quantifying the profit loss from using a family of restricted contracts, such as contracts that do not pay agents who finish after the deadline.
Finally, many additional stochastic optimization problems, such as stochastic routing and others, have natural strategic variants were the uncertainty is controlled by an agent’s actions. E.g., in routing, service times are likely to result from a strategic choice of an economic entity. Marrying the classic algorithms developed for such problems with contractual incentives is an intriguing avenue for future research.
Appendix A LP Overconfidence: An $\Omega(\alpha)$ Overestimation of the True Profit
In this section, we provide an instance of the MSKC problem where the value of the LP introduced in Eq. (2) is larger by a factor $\Omega(\alpha)$ from the obtainable value of any policy. This is in stark contrast to the case without costs, where Dean et al. [11] show a non-adaptive policy who’s reward is a constant factor away from the value of the analogous LP. As our non-adaptive policy in Section 3 gives $O(\alpha)$-approximation to the value of the LP in Eq. (2), this approximation is the best possible with respect to the value of the LP.
In our instance, we have an infinite number of identical jobs with value $1$ and cost $1-\epsilon$. The size-distribution of the jobs is
$s=\begin{cases}\epsilon^{n},&\text{w.p. }1-\epsilon,\\ 1,&\text{w.p. }\epsilon;\end{cases}.$
By Definition 2.6, the value of the IOR parameter is $\alpha=\frac{1-\epsilon}{\epsilon}=O(1/\epsilon)$. For every job $i$, we have $\mu_{i}=(1-\epsilon)\cdot\epsilon^{n}+\epsilon\cdot 1=\epsilon^{n}+\epsilon-\epsilon^{n+1}\leq 2\epsilon$ and $w_{i}=\epsilon$. Setting $x_{i}=$ for $1/2\epsilon$ is a feasible assignment to the LP, which implies that $\Phi(1)\geq\frac{1}{2\epsilon}\cdot\epsilon=1/2.$ We next bound the reward of any policy for this instance.
For any policy, after executing the first job, we have that the any job fits with probability at most $1-\epsilon$; thus, we have that the benefit from trying to execute any job after the first one is at most $(1-\epsilon)\epsilon-\epsilon\cdot(1-\epsilon)=0.$ Therefore, the policy that only executes one job is optimal, and it obtains a reward of $\epsilon\leq\frac{\epsilon}{2}\Phi(1)=\Phi(1)/\Omega(\alpha).$
Appendix B The Need for Adaptivity for Bounded-Size Jobs
Theorem B.1.
There exists an instance $I$ of the Knapsack Contracts problem such that
$\frac{\mathrm{STOP\_ADAPT}(I)}{\mathrm{NON\_ADAPT}(I)}=\Omega\bigl{(}\sqrt{\alpha}\bigr{)}.$
Proof.
Let item:
Property Item
Reward $v$ 22
Cost $c$ $2-\epsilon^{2}$
Size $s$ dist. $s_{1}=\begin{cases}\epsilon^{n},&\text{w.p. }1-\epsilon,\\ \frac{1}{2},&\text{w.p. }\epsilon;\end{cases}$
We consider the case where $S^{\prime}<\frac{1}{2}$, and compute the expected profit from each job.
$\operatorname{\mathbb{E}}[w|S^{\prime}<\frac{1}{2}]=v\cdot\Pr[s\leq S^{\prime}]-c=2\cdot(1-\epsilon)-(2-\epsilon^{2})=2-2\epsilon-2+\epsilon^{2}=-2\epsilon+\epsilon^{2}<0$
In expectation, we can include approximately $1/\epsilon$ jobs before reaching a state where $S^{\prime}<\frac{1}{2}$ so, the expected number of jobs executed by the optimal stopping-time adaptive policy is $1/\epsilon$, and therefore, it obtains an expected profit of $\frac{1}{\epsilon}\cdot\epsilon^{2}=\epsilon.$
Consider now a non-adaptive policy. This policy decides on the number of jobs to execute, and ends whether it overflows or completes the number of jobs it sets to execute. Consider first a policy that only executes one job. Its reward is exactly $v-c=\epsilon^{2}$. Let $\mathrm{profit}(i)$ be the expected profit from the $i$th job in the sequence. Since the first job always finishes before the deadline, it is clear that $\mathrm{profit}(1,2)=c-v=\epsilon^{2}$. Consider now $\mathrm{profit}(i)$ for $i>2$. We have that
$\mathrm{profit}(i)$ $=$ $\operatorname{\mathbb{E}}[\mathrm{profit}(i)|s_{j}=\epsilon^{n}\ \forall j<i]\cdot\Pr[s_{j}=\epsilon^{n}\quad\forall j<i]$
$+\operatorname{\mathbb{E}}[\mathrm{profit}(i)|\exists!j<i:\ s_{j}>\epsilon^{n}]\cdot\Pr[\exists!j<i:\ s_{j}>\epsilon^{n}]$
$=$ $(v-c)\cdot(1-\epsilon)^{i-1}+((v-c)\cdot(1-\epsilon)-c\cdot\epsilon)\cdot\Pr[\exists!j<i:\ s_{j}>0]$
$=$ $\epsilon^{2}\cdot(1-\epsilon)^{i-1}+((1-\epsilon)v-c)\cdot\Pr[\exists!j<i:\ s_{j}>0]$
$=$ $\epsilon^{2}\cdot(1-\epsilon)^{i-1}+(\epsilon^{2}-2\epsilon)\cdot\Pr[\exists!j<i:\ s_{j}>0]$
$\leq$ $\epsilon^{2}\cdot(1-\epsilon)^{i-1}+(\epsilon^{2}-2\epsilon)\cdot\Pr[s_{i-1}>0\ \wedge\ s_{j}=0\quad\forall j<i-1]$
$=$ $\epsilon^{2}\cdot(1-\epsilon)^{i-1}+(\epsilon^{2}-2\epsilon)\cdot(1-\epsilon)^{i-2}\cdot\epsilon$
$=$ $\epsilon\cdot(1-\epsilon)^{i-2}(\epsilon\cdot(1-\epsilon)+\epsilon^{2}-2\epsilon)$
$\leq$ $\epsilon\cdot(1-\epsilon)^{i-2}(\epsilon^{2}-\epsilon)$
$<$ $0.$
In the above, the first inequality follows from the fact that $\epsilon^{2}-2\epsilon$ is negative, and that the probability that exactly one of the first i1i-1 jobs has size greater than $\epsilon^{n}$ is larger than the probability that only job i1i-1 has size greater than $\epsilon^{n}$. By Definition 2.6, the value of the IOR parameter is $\alpha=\frac{2-\epsilon^{2}}{\epsilon^{2}}=O(1/\epsilon^{2})$.
Therefore, we have that a non-adaptive policy only loses from having a sequence larger than 2, and the optimal non-adaptive policy has a profit of $2\epsilon^{2}$. We conclude that for this instance, we have $\frac{\mathrm{STOP\_ADAPT}(I)}{\mathrm{NON\_ADAPT}(I)}=\frac{1}{2\epsilon}=O(\sqrt{\alpha}).$
Appendix C Constant Approximation for MSKC with Overflows
We first provide an upper-bound for expected profit of the optimal adaptive policy that can overflow as well. Recall the definition of $\Phi$ given in Eq (2). Let $\mathrm{ADAPT}_{OF}$ be the optimal adaptive policy that can overflow.
Lemma C.1.
For every instance of the knapsack contract problem $I$, $\mathrm{ADAPT}_{OF}(1)\leq 2\Phi(1)+w_{max}$.
Proof.
let $\mathcal{S}$ be the random set of item choices that the policy inserts, and $w_{ij}=v_{i}\cdot\Pr_{{s_{i}^{j}\sim F_{i}^{j}}}[s_{i}^{j}\leq 1]-p_{i}^{j}$ (like in Section 3), $w_{\max}=\max_{(i,j)\in\mathcal{C}}w_{ij}$ (Algorithm 1). $\Phi(t)$ is defined in Eq. (2), .
$\mathrm{ADAPT}_{OF}(1)$ $=$ $\sum_{ij}w_{ij}\cdot\Pr[(i,j)\in\mathcal{S}]$ (8)
$=$ $\sum_{ij}w_{ij}\cdot\Pr[(i,j)\in\mathcal{S}\wedge s_{i}\leq 1]+\sum_{ij}w_{ij}\cdot\Pr[(i,j)\in\mathcal{S}\wedge s_{i}>1]$
$\leq$ $\sum_{ij}w_{ij}\cdot\Pr[s_{i}\leq 1]\cdot\Pr[(i,j)\in\mathcal{S}]+w_{max}\sum_{ij}\Pr[(i,j)\in\mathcal{S}\wedge s_{i}>1]$
$\leq$ $\sum_{ij\in\mathcal{S}}w_{ij}\cdot x_{ij}+w_{max}$
$\leq$ $\Phi(2)+w_{max}$
$\leq$ $2\Phi(1)+w_{max}.$
Here the first inequity follows since the probability $s_{i}>1$ is independent from the fact $(i,j)\in\mathcal{S}$. The last inequality follows from the fact that, at each run of the adaptive policy, at most one element has a size larger than 1, and therefore the sum of probabilities of the right summand is at most 1. By Lemma 3.6, we know that $\operatorname{\mathbb{E}}[\mu(\mathcal{S})]\leq 2$, therefore, we get $\sum_{ij\in A}w_{ij}\cdot x_{ij}\leq\Phi(2)$. ∎
Theorem C.2.
$\textbf{SKC}_{OF}$ is a 9-approximation algorithm to $\mathrm{ADAPT}_{OF}$.
Proof.
By Equations (7),(8):
$\frac{ADAPT_{OF}}{\textbf{SKC-OF}}$ $\leq$ $\frac{2\Phi(1)+w_{max}}{\Phi(1)/4}.$
If $w_{max}>\Phi(1)/4$: so $\textbf{SKC-OF}=w_{max}$ and:
$\frac{ADAPT_{OF}}{\textbf{SKC-OF}}$ $\leq$ $\frac{8w_{max}+w_{max}}{w_{max}}=9.$
Else:
$\frac{ADAPT_{OF}}{\textbf{SKC-OF}}$ $\leq$ $\frac{2\Phi(1)+\Phi(1)/4}{\Phi(1)/4}=9.$
Appendix D Ordered Adaptive Contracts
In many real-world scenarios, the principal lacks the flexibility to reorder agents at will. For example, this can be due to procedural requirements, like a fixed sequence of suppliers in a supply chain, or pre-determined project plan with sequential dependencies. In these settings, the principal’s adaptive power is limited to deciding, after observing the outcome of each task, whether or not to continue with the next agent, and if so – which effort level to incentivize. We define the Ordered Adaptive Contracts model:
Definition D.1 (Ordered Adaptive Contracts Model).
In the Ordered Adaptive Contracts model, the principal is given a fixed ordering $\sigma=(1,2,\ldots,n)$ of the $n$ agents. For each agent $i$ in order, the principal offers a contract $t_{i}$, and the agent chooses an effort level $j$ to maximize their utility. The agents execute their jobs sequentially in the given order. The principal observes the completion time of each job and may adaptively decide whether to proceed to the next agent or stop, but cannot change the order. The objective is to maximize the expected net profit (total value from completed jobs before the budget is exhausted minus total payments to agents), given the fixed order $\sigma$.
Since the order is fixed, we are interested in comparing our policy’s performance to the performance of the optimal policy operating on the same order of agents. By using the same reduction to the multi-agent contract problem as shown in Lemma 2.1, we focus on solving the matching stochastic knapsack with costs problem. In a similar manner to the algorithm that finds an ordered adaptive policy presented in [11] and to the PTAS presented in [5] for the regular Stochastic Knapsack variant (single-choice, no-cost), we use a resource augmentation approach: the policy is allowed a knapsack of size $(1+\epsilon)$, while the optimal adaptive policy has a knapsack of size $1$. Our algorithm runs in time $O(\frac{n^{3}m}{\epsilon^{2}})$, offering a tradeoff between the augmented size of the knapsack and the running time of the algorithm.
We use a dynamic programming algorithm with discretized sizes. We denote by $\mathcal{D}(i,k)$ the optimal expected value that can be obtained from inserting items i,,ni,...,n, with $\delta k$ capacity left in the knapsack, for some small $\delta>0$ to be determined later. We are interested in computing $\mathcal{D}(1,\lfloor\frac{1}{\delta}\rfloor)$. If all of the real sizes are indeed multiples of $\delta k$, then one can obtain an optimal policy by computing:
$\mathcal{D}(i,k)=\max\left\{\mathcal{D}(i+1,k),\max_{j\in[m]}\left\{v_{i}\cdot\Pr[s_{i}^{j}\leq\delta k]-p_{i}^{j}+\sum_{t=0}^{k}{\mathcal{D}(i+1,k-t)\cdot\Pr[s_{i}^{j}=t\delta]}\right\}\right\}.$
With the base cases of $\mathcal{D}(n+1,k)=0$ for all $k$, and $\mathcal{D}(i,k)=0$ for $k<0$.
However, when the distributions are arbitrary, a slight deviation from the discretization can result in an arbitrarily bad solution. To overcome this, we must utilize the fact our policy is allowed a knapsack of size $1+\epsilon$. We do so by rounding down the true distribution at each step to the nearest multiple of $\delta$. By picking a small enough $\delta$, we ensure that the overall rounding error is not too large.
Theorem D.1.
There exists an $O(\frac{n^{3}m}{\epsilon^{2}})$ time algorithm that finds a policy for the ordered adaptive knapsack with costs problem, where the policy has a knapsack of size $(1+\epsilon)$, and the optimal adaptive policy has a knapsack of size $1$. The policy achieves an expected profit of at least the profit of the optimal policy with a knapsack of size $1$.
Proof.
Let OAC be our algorithm and $\mathrm{ADAPT}$ be the optimal adaptive policy. We set the discretization parameter $\delta=\epsilon/n$.
First, consider a modified problem where all item sizes $s_{i}^{j}$ are rounded down to the nearest multiple of $\delta$: $\hat{s}_{i}^{j}=\delta\lfloor s_{i}^{j}/\delta\rfloor$. Let $\widehat{\mathrm{ADAPT}}$ be the optimal expected profit for this problem with rounded-down sizes. Since for any policy, the probability of successfully fitting a set of items can only increase when item sizes are reduced, the optimal value for the rounded-down problem is an upper bound on the optimal value for the original problem. Thus, $\widehat{\mathrm{ADAPT}}\geq\mathrm{ADAPT}.$
Our algorithm solves the dynamic program $\mathcal{D}$ described above for the rounded-down sizes. This DP computes the exact optimal policy for the rounded-down problem. Denote by $\varphi_{\textbf{OAC}}$ the policy derived from this DP. When given a knapsack of capacity $1$, this policy achieves the optimal value for the rounded-down problem. Let $\widehat{\mathrm{profit}}(\varphi_{\textbf{OAC}},\text{capacity}=1)$ be a random variable of the profit of the policy with the original capacity and rounded-down job sizes. We have that $\operatorname{\mathbb{E}}[\widehat{\mathrm{profit}}(\varphi_{\textbf{OAC}},\text{capacity}=1)]=\widehat{\mathrm{ADAPT}}$. Combining these, we have $\operatorname{\mathbb{E}}[\widehat{\mathrm{profit}}(\varphi_{\textbf{OAC}},\text{capacity}=1)]\geq\mathrm{ADAPT}.$
Now, we analyze the performance of our algorithm, which executes the policy $\varphi_{\textbf{OAC}}$ on the original item sizes $s_{i}^{j}$ with an augmented capacity of $1+\epsilon$. Consider any execution path of $\varphi_{\textbf{OAC}}$ where a sequence of items and choices $\mathcal{S}\subset[n]\times[m]$ is attempted and successfully completed. According to the policy $\varphi_{\textbf{OAC}}$ (which was computed for capacity 1 and sizes $\hat{s}$), the sum of the rounded-down sizes must be no more than 1:
$\sum_{(i,j)\in\mathcal{S}}\hat{s}_{i}^{j}\leq 1.$
The true total size of these items is $\sum_{(i,j)\in\mathcal{S}}s_{i}^{j}$. We know that for each item, $s_{i}^{j}=\hat{s}_{i}^{j}+\eta_{i}^{j}$, where $0\leq\eta_{i}^{j}<\delta$ is the rounding error. The total size in the original problem is:
$\sum_{(i,j)\in\mathcal{S}}s_{i}^{j}=\sum_{(i,j)\in\mathcal{S}}\hat{s}_{i}^{j}+\sum_{(i,j)\in\mathcal{S}}\eta_{i}^{j}.$
Since at most $n$ items can be placed in the knapsack, the total accumulated rounding error is bounded:
$\sum_{(i,j)\in\mathcal{S}}\eta_{i}^{j}\leq n\cdot\delta=n\cdot(\epsilon/n)=\epsilon.$
Therefore, the true total size of the items is:
$\sum_{(i,j)\in\mathcal{S}}s_{i}^{j}\leq\sum_{(i,j)\in\mathcal{S}}\hat{s}_{i}^{j}+\epsilon\leq 1+\epsilon.$
This shows that any sequence of items that fits within capacity 1 in the rounded-down instance is guaranteed to fit within the augmented capacity of $1+\epsilon$ in the original instance. Since the policy $\varphi_{\textbf{OAC}}$ makes decisions based on the sequence of outcomes (which are independent of the rounding), every successful path under capacity 1 with sizes $\hat{s}$ is also a successful path under capacity $1+\epsilon$ with sizes $s$. Moreover, it crucially follows that an item can cause an overflow under capacity $1+\epsilon$ with sizes $s$ only if it causes an overflow under capacity 1 with sizes $\hat{s}$. Therefore, the expected profit can only be higher:
$\operatorname{\mathbb{E}}[\mathrm{profit}(\varphi_{\textbf{OAC}},\text{capacity}=1+\epsilon)]\geq\operatorname{\mathbb{E}}[\widehat{\mathrm{profit}}(\varphi_{\textbf{OAC}},\text{capacity}=1)].$
Chaining the inequalities, we get the final result:
$\textbf{OAC}=\operatorname{\mathbb{E}}[\mathrm{profit}(\varphi_{\textbf{OAC}},\text{capacity}=1+\epsilon)]\geq\operatorname{\mathbb{E}}[\widehat{\mathrm{profit}}(\varphi_{\textbf{OAC}},\text{capacity}=1)]\geq\mathrm{ADAPT}.$
The dynamic program $\mathcal{D}$ has $O(n)\cdot O(n/\epsilon)$ states, and computing each state takes up to $O(m\cdot n/\epsilon)$ time. This results in a running time of $O(\frac{n^{3}m}{\epsilon^{2}})$.
Appendix E Omitted Proofs from Section 3
Full proof of Lemma 3.3.
For two solutions $x,x^{*}$, denote by $\Delta_{w}$ and $\Delta_{\mu}$ the difference in the total values and the total expected truncated sizes, respectively, between $x^{*}$ and $x$. We want to show that our constructed solution $x^{*}$ has at least as much value ($\Delta_{w}\geq 0$) and at most as much expected truncated size ($\Delta_{\mu}\leq 0$) as the original solution $x$, to conclude that $x^{*}$ is a feasible optimal solution.
For the single dominance case, we define
$x^{*}_{ij^{\prime}}=x_{ij^{\prime}}+x_{ij},\quad x^{*}_{ij}=0~.$
We have that
$\Delta_{w}$ $=(x_{ij^{\prime}}^{*}-x_{ij^{\prime}})w_{ij^{\prime}}+(x_{ij}^{*}-x_{ij})w_{ij}$
$=x_{ij}w_{ij^{\prime}}-x_{ij}w_{ij}$
$=x_{ij}(w_{ij^{\prime}}-w_{ij})$
$\geq 0$ $\text{(since $w_{ij^{\prime}}\geq w_{ij}$)}~,$
$\Delta_{\mu}$ $=(x_{ij^{\prime}}^{*}-x_{ij^{\prime}})\mu_{ij^{\prime}}+(x_{ij}^{*}-x_{ij})\mu_{ij}$
$=x_{ij}\mu_{ij^{\prime}}-x_{ij}\mu_{ij}$
$=x_{ij}(\mu_{ij^{\prime}}-\mu_{ij})$
$\leq 0$ $\text{(since $\mu_{ij^{\prime}}\leq\mu_{ij}$)}~.$
For the double dominance case, we define
$x^{*}_{ij^{\prime}}=x_{ij^{\prime}}\;+\;x_{ij}\cdot\frac{w_{ij^{\prime\prime}}-w_{ij}}{w_{ij^{\prime\prime}}-w_{ij^{\prime}}},\quad x^{*}_{ij^{\prime\prime}}=x_{ij^{\prime\prime}}\;+\;x_{ij}\cdot\frac{w_{ij}-w_{ij^{\prime}}}{w_{ij^{\prime\prime}}-w_{ij^{\prime}}},\quad x^{*}_{ij}=0~.$
We have that
$\Delta_{w}$ $=(x_{ij^{\prime}}^{*}-x_{ij^{\prime}})w_{ij^{\prime}}+(x_{ij^{\prime\prime}}^{*}-x_{ij^{\prime\prime}})w_{ij^{\prime\prime}}+(x_{ij}^{*}-x_{ij})w_{ij}$
$=x_{ij}w_{ij^{\prime}}\cdot\frac{w_{ij^{\prime\prime}}-w_{ij}}{w_{ij^{\prime\prime}}-w_{ij^{\prime}}}+x_{ij}w_{ij^{\prime\prime}}\cdot\frac{w_{ij}-w_{ij^{\prime}}}{w_{ij^{\prime\prime}}-w_{ij^{\prime}}}-x_{ij}w_{ij}$
$=x_{ij}\left(\frac{w_{ij^{\prime}}(w_{ij^{\prime\prime}}-w_{ij})+w_{ij^{\prime\prime}}(w_{ij}-w_{ij^{\prime}})-w_{ij}(w_{ij^{\prime\prime}}-w_{ij^{\prime}})}{w_{ij^{\prime\prime}}-w_{ij^{\prime}}}\right)$
$=0~,$
$\Delta_{\mu}$ $=(x_{ij^{\prime}}^{*}-x_{ij^{\prime}})\mu_{ij^{\prime}}+(x_{ij^{\prime\prime}}^{*}-x_{ij^{\prime\prime}})\mu_{ij^{\prime\prime}}+(x_{ij}^{*}-x_{ij})\mu_{ij}$
$=x_{ij}\mu_{ij^{\prime}}\cdot\frac{w_{ij^{\prime\prime}}-w_{ij}}{w_{ij^{\prime\prime}}-w_{ij^{\prime}}}+x_{ij}\mu_{ij^{\prime\prime}}\cdot\frac{w_{ij}-w_{ij^{\prime}}}{w_{ij^{\prime\prime}}-w_{ij^{\prime}}}-x_{ij}\mu_{ij}$
$=x_{ij}\left(\frac{\mu_{ij^{\prime}}(w_{ij^{\prime\prime}}-w_{ij})+\mu_{ij^{\prime\prime}}(w_{ij}-w_{ij^{\prime}})-\mu_{ij}(w_{ij^{\prime\prime}}-w_{ij^{\prime}})}{w_{ij^{\prime\prime}}-w_{ij^{\prime}}}\right)~.$
From double dominance we have that
$\frac{\mu_{ij}-\mu_{ij^{\prime}}}{w_{ij}-w_{ij^{\prime}}}\geq\frac{\mu_{ij^{\prime\prime}}-\mu_{ij}}{w_{ij^{\prime\prime}}-w_{ij}}$
Since $w_{ij}-w_{ij^{\prime}}>0$ and $w_{ij^{\prime\prime}}-w_{ij}>0$, we can cross-multiply:
$(\mu_{ij}-\mu_{ij^{\prime}})(w_{ij^{\prime\prime}}-w_{ij})\geq(\mu_{ij^{\prime\prime}}-\mu_{ij})(w_{ij}-w_{ij^{\prime}})$
Expanding both sides gives:
$\mu_{ij}w_{ij^{\prime\prime}}-\mu_{ij}w_{ij}-\mu_{ij^{\prime}}w_{ij^{\prime\prime}}+\mu_{ij^{\prime}}w_{ij}\geq\mu_{ij^{\prime\prime}}w_{ij}-\mu_{ij^{\prime\prime}}w_{ij^{\prime}}-\mu_{ij}w_{ij}+\mu_{ij}w_{ij^{\prime}}$
We can cancel the $\mu_{ij}w_{ij}$ term from both sides. Rearranging the terms to match the numerator of our expression for $\Delta_{\mu}$ yields:
$\mu_{ij^{\prime}}(w_{ij^{\prime\prime}}-w_{ij})+\mu_{ij^{\prime\prime}}(w_{ij}-w_{ij^{\prime}})-\mu_{ij}(w_{ij^{\prime\prime}}-w_{ij^{\prime}})\leq 0~.$
This confirms the numerator is non-positive. Since $w_{ij^{\prime\prime}}>w_{ij^{\prime}}$ we conclude that $\Delta_{\mu}\leq 0$.
Proof of Lemma 3.5.
$\Pr[size(\mathcal{S})\geq t]$ $=\Pr[\min\{size(\mathcal{S}),t\}\geq t]$
$\leq\frac{1}{t}\operatorname{\mathbb{E}}[\min\{size(S),t\}]$
$=\frac{1}{t}\operatorname{\mathbb{E}}[{\min\{\sum_{(i,j)\in\mathcal{S}}{s_{i}^{j}},t\}}]$
$\leq\frac{1}{t}\operatorname{\mathbb{E}}[\sum_{(i,j)\in\mathcal{S}}{\min\{s_{i}^{j},t\}}]$
$=\frac{1}{t}\sum_{(i,j)\in\mathcal{S}}{\operatorname{\mathbb{E}}[\min\{s_{i}^{j},t\}]}$
$=\frac{1}{t}\sum_{(i,j)\in\mathcal{S}}{\mu_{ij}}=\frac{1}{t}\mu(S)$
Proof of Lemma 3.6.
Let $\mathcal{S^{-}}$ be the (random) set of item choices that the policy inserts except the final item choice. By the stopping condition
$size(\mathcal{S}^{-})<t~.$
It follows that for every $(i,j)\in\mathcal{S}^{-}$, $s_{i}^{j}<t$ and therefore $\mu_{ij}=\operatorname{\mathbb{E}}[\min\{s_{i}^{j},t\}]=\operatorname{\mathbb{E}}[s_{i}^{j}]$. Denote by $(i^{*},j^{*})$ the final item choice made by the policy. Then,
$\operatorname{\mathbb{E}}[\mu(\mathcal{S})]$ $=\operatorname{\mathbb{E}}[\sum_{(i,j)\in\mathcal{S}}{\mu_{ij}}]=\operatorname{\mathbb{E}}[\sum_{(i,j)\in\mathcal{S}^{-}}{\mu_{ij}}]+\operatorname{\mathbb{E}}[\mu_{i^{*}j^{*}}]$
$=\operatorname{\mathbb{E}}[\sum_{(i,j)\in\mathcal{S}^{-}}{s_{i}^{j}}]+\operatorname{\mathbb{E}}[s_{i^{*}}^{j^{*}}]=\operatorname{\mathbb{E}}[size(\mathcal{S}^{-})]+\operatorname{\mathbb{E}}[s_{i^{*}}^{j^{*}}]$
$\leq t+t=2t~.$
Proof of Lemma 3.7.
For any adaptive policy, denote by $\mathcal{S}$ the random set of pairs $(i,j)$ that the policy inserts. Set $x_{ij}=\Pr[(i,j)\in\mathcal{S}]$. From the previous lemma, the expected mass that $\mathrm{ADAPT}$ attempts to insert is $\operatorname{\mathbb{E}}(\mu(\mathcal{S}))=\sum_{ij}\mu_{ij}x_{ij}\leq 2t$, therefore, it is a feasible solution to $\Phi(2t)$ with $\sum_{ij}w_{ij}x_{ij}\leq\Phi(2t)$. Since $w_{ij}x_{ij}$ is an upper bound on the expected profit of inserting the item choice $(i,j)$, we have that
$\mathrm{ADAPT}\leq\sum_{ij}{w_{ij}x_{ij}}\leq\Phi(2t)\leq\Phi(ct)~,$
for any $c\geq 2$. Now, given an optimal solution $x$ to $\Phi(ct)$, we can construct a feasible solution $x^{\prime}$ to $\Phi(t)$ by setting $x^{\prime}_{ij}=\frac{x_{ij}}{c}$ for all $i,j$, and then increasing the value of $x^{\prime}_{ij_{\bot}}$ to ensure that $\sum_{j}x^{\prime}_{i,j}=1$ for all $i$. This results in a feasible solution to $\Phi(t)$ with $\sum_{ij}w_{ij}x^{\prime}_{ij}=\frac{1}{c}\sum_{ij}w_{ij}x_{ij}$. We conclude that
$\mathrm{ADAPT}\leq\Phi(ct)\leq c\cdot\Phi(t)$
Appendix F Proof of Lemma 2.1
Proof.
Given a knapsack contracts instance, for each agent $i\in[n]$, we create a corresponding item $i$ in the MSKC instance. For each implementable effort level $j$ of agent $i$, we create a corresponding choice $j$ for item $i$. The parameters are mapped as follows:
  • The value of item $i$ is set to the principal’s reward $v_{i}$.
  • The size of choice $(i,j)$ is a random variable with the distribution $F_{i}^{j}$.
  • The cost of choice $(i,j)$ is set to $p_{i}^{j}=\operatorname{\mathbb{E}}_{s\sim F_{i}^{j}}[t_{i}(s)]$, where $t_{i}$ is the contract that minimizes the expected payment, as given by the LP in Definition 2.1.
Now, consider any policy $\varphi$ for the contracts problem, and denote by $\varphi^{\prime}$ the corresponding policy for the knapsack problem. $\varphi$ specifies the choice of (agent, effort level) pairs for the contracts problem, and $\varphi^{\prime}$ specifies the corresponding (item, choice) pairs for the knapsack problem.
It is left to show that the expected profits are identical. Denote by $S$ the set of successfully completed jobs in the contracts and knapsack problems, and by $A$ the set of agents/items attempted in the contracts and knapsack problems. Notice that $S$ and $A$ have exactly the same distribution under $\varphi$ and $\varphi^{\prime}$, as both policies make the same choices under the same randomization.
The profit for policy $\varphi$ is a random variable that depends on the (possibly random) choices of the policy and the random sizes of the attempted jobs. Its expectation is given by
$\operatorname{\mathbb{E}}[\mathrm{profit}(\varphi)]=\operatorname{\mathbb{E}}\left[\sum_{(i,j)\in S}{v_{i}}-\sum_{(i,j)\in A}{t_{i}(s_{i})}\right]=\operatorname{\mathbb{E}}\left[\sum_{(i,j)\in S}{v_{i}}\right]-\operatorname{\mathbb{E}}\left[\sum_{(i,j)\in A}{t_{i}(s_{i})}\right].$ (9)
Expanding the transfer term, we get:
$\operatorname{\mathbb{E}}\left[\sum_{(i,j)\in A}{t_{i}(s_{i})}\right]$ $=\operatorname{\mathbb{E}}\left[\sum_{(i,j)\in A}{t_{i}(s_{i})}\right]=\operatorname{\mathbb{E}}\left[\sum_{(i,j)}{\mathbb{1}_{(i,j)\in A}\cdot t_{i}(s_{i})}\right]=\sum_{(i,j)}\operatorname{\mathbb{E}}\left[{\mathbb{1}_{(i,j)\in A}\cdot t_{i}(s_{i})}\right]$
$\stackrel{{\scriptstyle(\star)}}{{=}}\sum_{(i,j)}\operatorname{\mathbb{E}}\left[\mathbb{1}_{(i,j)\in A}\right]\cdot\operatorname{\mathbb{E}}_{s_{i}\sim F_{i}^{j}}\left[t_{i}(s_{i})\right]\stackrel{{\scriptstyle(\star\star)}}{{=}}\sum_{(i,j)}\operatorname{\mathbb{E}}\left[\mathbb{1}_{(i,j)\in A}\right]\cdot p_{i}^{j}$
$=\operatorname{\mathbb{E}}\left[\sum_{(i,j)}\mathbb{1}_{(i,j)\in A}\cdot p_{i}^{j}\right]=\operatorname{\mathbb{E}}\left[\sum_{(i,j)\in A}p_{i}^{j}\right],$
where equality $(\star)$ is due to the independence between $\mathbb{1}_{(i,j)\in A}$ and $s_{i}^{j}$, and equality $(\star\star)$ is due to the definition of $p_{i}^{j}$. Plugging this back in Eq. 9 yields:
$\operatorname{\mathbb{E}}[\mathrm{profit}(\varphi)]=\operatorname{\mathbb{E}}\left[\sum_{(i,j)\in S}{v_{i}}-\sum_{(i,j)\in A}p_{i}^{j}\right]=\operatorname{\mathbb{E}}[\mathrm{profit}(\varphi^{\prime})]~,$
which implies the desired.
Thus, given a $\varphi^{\prime}$ that $c$-approximate the optimal policy in the constructed instance of the MSKC problem, the corresponding $\varphi$ $c$-approximates the optimal policy in the original knapsack contracts instance. ∎

### v2

@@
 Abstract
-We formulate the Knapsack Contracts problem—a strategic version of the classic Stochastic Knapsack problem, which builds upon the inherent randomness shared by stochastic optimization and contract design.
-In this problem, the principal incentivizes agents to perform jobs with stochastic processing times, the realization of which depends on the agents’ efforts.
-Algorithmically, we show that Knapsack Contracts can be viewed as Stochastic Knapsack with costs and multi-choice, features that introduce significant new challenges.
-We identify a crucial and economically meaningful parameter—the Return on Investment (ROI) value.
-We show that the Inverse of ROI (or IOR for short) precisely characterizes the extent to which the approximation guarantees for Stochastic Knapsack extend to its strategic counterpart.
-For IOR of $\alpha$, we develop an algorithm that finds an $O(\alpha)$-approximation policy that does not rely on adaptivity.
-We establish matching $\Omega(\alpha)$ lower bounds, both on the adaptivity gap, and on what can be achieved without full distributional knowledge of the processing times.
-Taken together, our results show that IOR is fundamental to understanding the complexity and approximability of Knapsack Contracts, and bounding it is both necessary and sufficient for achieving non-trivial approximation guarantees.
-Our results highlight the computational challenges arising when stochasticity in optimization problems is controlled by strategic effort.
+We revisit the Stochastic Knapsack problem, where a policy-maker chooses an execution order for jobs with fixed values and stochastic running-times, aiming to maximize the value completed by a deadline.
+Dean et al.
+(FOCS’04) show that simple non-adaptive policies can approximate the (highly adaptive) optimum, initiating the study of adaptivity gaps.
+We introduce an economically motivated generalization in which each job also carries an execution cost.
+This uncovers new applications, most notably a new and natural variant of contract design: jobs are processed by agents who choose among effort levels that induce different processing-time distributions, and the principal must decide which jobs to run and what payments induce the intended effort.
+With costs, the objective becomes mixed-sign: value from completed jobs must be balanced against costs of execution, and running a job that misses the deadline can create negative utility.
+This changes the algorithmic picture, and the adaptivity gap is no longer constant.
+We give an economic explanation: the performance of non-adaptive policies is governed by jobs’ return on investment (ROI) — utility over cost — which can be arbitrarily small.
+We introduce a hierarchy of increasingly adaptive policies, trading off simplicity and adaptivity.
+We prove near-tight guarantees across the hierarchy, showing that with costs the adaptivity gap is $\Theta(\alpha)$, where $1/\alpha$ is the ROI.
+Higher in the hierarchy, we identify an efficiently computable policy with limited adaptivity that is approximately-optimal.
+Analogous to the centrality of ROI in economics, we believe our ROI-based, simple-vs-optimal approach to adaptivity may be useful for additional stochastic optimization and online problems with mixed-sign objectives.
 1 Introduction
-The study of how incentives reshape classic algorithmic problems has been at the heart of algorithmic game theory research since its inception.
-In their seminal work, Nisan and Ronen [24] famously conjectured that adding incentives for truthfulness to the optimization problem of scheduling would make the problem extremely hard to approximate.
-Two decades later, intensive research has culminated in a proof of this conjecture [9].
-Many additional works explore the intersection of algorithm design with truthfulness constraints [27, 25, 10, 6, 8, 22, 23].
-Over recent years, the algorithmic game theory community has begun to explore incentive constraints beyond truthfulness.
-A recent focus is on an agent’s earnestness in exerting effort to complete a task, where the task is delegated to the agent by a principal.
-The main economic tool for incentivizing earnest effort is performance-based payment schemes, a.k.a.
-contracts.
-Contracts are designed to create incentives that counteract moral hazard: When an agent exerts effort to carry out a job on behalf of a principal, this leads to a stochastic outcome that only partially reflects the agent’s effort.
-This stochasticity opens the possibility for the agent to claim they worked hard despite exerting a low level of effort, by blaming poor outcomes on “bad luck”.
-This phenomenon is known as moral hazard.
-A contract uses outcome-based payments and distributional knowledge to incentivize earnest effort, thus mitigating moral hazard.
-The goal of this work is to initiate the study of the intersection of algorithm design with contractual incentives for earnest effort.
-1.1 Problem Formulation
-To achieve our goal, we seek an algorithmic optimization problem with a natural connection to contractual incentives—similar to the connection between scheduling and truthfulness [24].
-Our starting point is the observation that the field of contract design has a natural affinity to stochastic optimization, due to the nondeterminism inherent to both.
-Building upon this connection, we formulate a model of Knapsack Contracts, in which jobs that need to fit into a given time frame are delegated to agents.
-In our model, the agents have costs for carrying out the jobs, and the completion times (a.k.a.
-job lengths or sizes) are a stochastic outcome of the level of effort.
-The latter makes our model a generalization of Stochastic Knapsack [11].
-We show that the agents’ costs lead to incentive constraints that make the problem significantly harder than classic Stochastic Knapsack.
-Since Stochastic Knapsack has a central place in stochastic optimization [3, 19, 20, 18, 28, 5, 2, 21], we view its study through the prism of contract design as a potential gateway to deeper exploration of the intersection between the two areas.
-Example 1.
-A PI (principal) wishes to run a set of experiments on the cloud before the deadline of a top AI conference.
-Each experimental task has a fixed value for the paper, which does not depend on how fast the cloud providers (agents) finish running the experiments.
-The PI does not benefit from experiments that finish running after the submission deadline.
-Despite this, she still needs to pay the providers for running them.
-The PI must spend substantial funds from her research grant to incentivize the agents to use powerful but costly resources when processing the experiments, resulting in shorter running times.
-This leads to an interesting tension between the PI’s benefit from an experiment, and how much they are willing to pay to incentivize an agent to run it quickly.
-The problem of optimizing the experiments thus has a mixed-sign objective.
-Generalizing Example 1, we consider a single principal and multiple agents with jobs that must be executed sequentially.
-Each agent $i$ is associated with a fixed known value $v_{i}$ to the principal for completing their job before the deadline expires.
-Moreover, each agent can choose one of $m$ costly actions, and the chosen action $j\in[m]$ determines a distribution $F_{i}^{j}$ over the possible completion times of the job.
-The principal cannot directly observe the action taken by an agent, and can only inspect the completion times.
-To incentivize agent $i$, the principal designs a contract $t_{i}$ that maps the completion time $s_{i}$ to a monetary transfer.
-Each agent $i$ wants to maximize their own utility: the transfer received minus the cost $c_{i}^{j}$ exerted by them when taking action $j$.
-The principal wants to maximize the profit, which is the sum of values over all jobs completed before the deadline, minus the sum of transfers made to the agents (i.e., a mixed-sign objective).
-See Section 2 for a formal problem description.
-Relation to vanilla contract design.
-Usually in contract settings, the value created for the principal by the agent’s effort is stochastic.
-Our model captures settings in which the principal’s value for the job delegated to the agent is fixed, but the time it takes the agent to complete the job is a random variable, and if the job doesn’t finish on time then the resulting value is zero.
-This seemingly simple twist on the standard principal-agent model leads to new challenges.
-Relation to Stochastic Knapsack.
-In the standard Stochastic Knapsack problem, there are $n$ jobs to execute, each with a fixed value but stochastic size.
-A policy-maker executes jobs sequentially, gaining the value of any job that terminates before the deadline.
-This process ends once a job terminates after the deadline.
-The optimal policy for this problem can be highly complex, adaptive, and intractable.
-Yet, Dean et al.
-[11] show that there exists a simple non-adaptive policy that approximates, up to a constant factor, the reward of the optimal adaptive policy.
-In contrast, in our model, even if a job ends after the deadline, the principal must pay the agent according to the contract, for processing it.
-As we show in Section 2.2, our model corresponds to a version of Stochastic Knapsack with job costs and multi-choice, and we show this makes a substantial difference relative to [11].
-A first indication of the difference arises from the linear programming approach of [11], which they use to establish their main result.
-In their setting, the integrality gap of the LP is bounded; however, in Appendix A we demonstrate that the LP approach can grossly overshoot the true optimum in our setting.
-In fact, as we shall see, in our model there is no constant-approximation non-adaptive policy; any constant-approximation algorithm must adaptively choose the next job to execute.
-Intuitively, adaptivity is necessary to avoid the danger of incurring a very high cost, so high it overshadows any value the algorithm may obtain.
-1.2 Our Results and Techniques
+In many economic decision-making settings, actions have stochastic outcomes and are associated not only with value but also with inherent costs.
+While most of the existing literature focuses on actions that generate value alone, in this paper we study how incorporating costs can significantly change the algorithmic design of effective solution methods.
+Understanding how costs affect decision-making scenarios is a central challenge in designing practical, near-optimal policies in many practical settings.
+Stochastic Knapsack and Adaptivity Gaps.
+In this work, we consider one of the classic stochastic decision problems studied in the seminal paper of Dean et al.
+[10]: the Stochastic Knapsack problem.
+In this problem, a policy-maker designs a policy that sequentially executes jobs, each associated with a known value and an uncertain processing time.
+The policy-maker faces a fixed deadline and obtains value only from jobs that complete before this deadline; any job finishing after the deadline yields zero value.
+While processing times are uncertain, the policy-maker is assumed to know their distributions, and seeks a policy that maximizes the expected total value obtained before the deadline.
+A key insight of Dean et al.
+[10] is the trade-off between policy simplicity and optimality, formalized through the notion of adaptivity gap.
+The optimal policy may adaptively choose the next job to process based on the realized processing times of previously executed jobs.
+Such adaptivity can be extremely complex: the optimal policy can be represented as a decision tree, whose size might be exponential in the number of jobs.
+Motivated by this potential complexity, Dean et al.
+[10] study simple non-adaptive policies, in which jobs are ordered in advance and executed sequentially until the deadline, collecting rewards from those that complete in time.
+Despite their simplicity, they show that such policies achieve a constant-factor approximation of the optimal adaptive policy, implying that the adaptivity gap is bounded by a constant.
+While the results of Dean et al.
+[10] demonstrate the surprising power of simple non-adaptive policies, they rely on the assumption that jobs can be executed without incurring costs.
+Costly Job Execution and Multiple Execution Modes.
+In many fundamental applications, assuming job execution is cost-free is unrealistic.
+We therefore study a setting with costly job execution, motivated by applications such as cloud computing or contracting with workers, where executing a job requires explicit compensation and thus incurs a cost.
+Moreover, each job may be executed at multiple levels of effort, where higher effort corresponds to higher cost, and different efforts induce different distributions over processing times.
+For example, the choice of the machine used to process a job, or of the specialist hired to perform a task, affects both the execution cost and the processing-time distribution.
+Although these extensions may appear modest, they fundamentally change the role of adaptivity.
+Notably, we show that when execution costs are present, the adaptivity gap of non-adaptive policies can be unbounded.
+Nevertheless, we identify a natural and economically meaningful parameter of the setting—capturing a notion of return on investment (ROI)—that governs the adaptivity gap in the presence of costs.
+We then show that a limited amount of carefully-applied adaptivity suffices to obtain policies whose performance is within a constant factor of the optimal adaptive policy.
 Return on Investment.
-The example in Appendix A is based on a “hard instance” of Stochastic Knapsack with job costs.
-We show this instance (and in fact any other hard instance) has the following property: For some of the jobs, their cost almost “covers” their value.
-Thus, in order to incentivize such jobs, the principal has to pay the agent their cost, and therefore, they are left with a small return on investment (ROI).
-I.e., the ratio between the obtained profit and the transfer made is small.
-Such jobs are intuitively problematic—if their cost must be reimbursed by the principal, but their value is not gained since they finish after the deadline, the loss can be substantial.
+We notice that “hard instances” of our problem have the following structural property: For some of the jobs, their cost almost eclipses their value.
+Thus, when executing such jobs, they provide a small return on investment (ROI).
+More formally, the ratio between the obtained utility (value minus cost) from running the job and the incurred cost is small.
+Such jobs are intuitively problematic in the knapsack context — their execution adds to the policy-maker’s cost, but their value might be lost if they finish after the deadline, in which case the policy-maker’s loss can be substantial.
 We essentially show that these are the only problematic jobs.
-In order to parameterize this, we inspect the Inverse of the ROI (or IOR) of a job (as the ROI becomes small, the IOR grows).
+Hierarchy of Adaptivity.
+Dean et al. [10] studied the Stochastic Knapsack problem by considering two extremes:
+  • Non-adaptive policies. These policies choose an ordered set of jobs. The set is then processed in order until either all jobs are completed or the deadline is met, whichever occurs first.
+  • Fully-adaptive policies. This is the most general policy class: there is no predetermined job order. The policy chooses which job to run next based on the observed runtime history.
+When costs are introduced, simple non-adaptive policies no longer guarantee a constant approximation to the value of the optimal fully-adaptive policies.
+We explore the landscape through the simplicity-optimality lens, defining and analyzing the following intermediate classes of policies:
+  • Stopping-adaptive policies.
+Policies in this class choose an ordered set of jobs, but stop executing once, given the time left until the deadline, any future sequence of jobs results in a negative expected utility.
+  • Skipping-adaptive policies. These policies also commit to an ordered list, but unlike stopping-adaptive policies, they may adaptively skip some of the listed jobs.
+See Section 2.2 for formal definitions and how the decision tree policy representation size decreases along the hierarchy.
+In the appendix we also consider non-stopping policies, which are non-adaptive policies that never halt and execute all jobs in the ordered set even after the deadline has passed.
+When there are multiple execution modes in addition to costs, the above classes of policies extend naturally by ordering pairs of (job, execution mode) rather than just ordering jobs.
+Motivating Example.
+To illustrate the differences among these policy types, consider the following motivating example, in which a contractor hires subcontractors to perform jobs under a time constraint.
+The contractor’s level of commitment and flexibility varies across policy classes.
+Under a fully-adaptive policy, the contractor never commits to an execution order in advance.
+After each observed runtime, the contractor may choose any remaining job and execution mode next, redesigning the plan on the fly as a function of the realized history and the remaining time.
+Such a high degree of adaptivity, however, can significantly complicate the system, both for the contractor, who must continuously re-optimize decisions, and for the subcontractors, who face uncertainty regarding their scheduling times and effort level.
+We therefore turn to simpler policy classes that impose an underlying fixed execution order on the jobs, specified by a priority list of job–execution-level pairs determined in advance.
+Under non-adaptive policies, once this priority list is fixed, the contractor is obliged to continue attempting jobs in order whenever any time remains.
+In this setting, a job that is highly unlikely to finish on time will still get executed,
+incurring its execution cost without obtaining its value. As we show, this behavior can be highly inefficient, particularly when the return on investment of the final job is low.
+To mitigate this issue, stopping-adaptive policies introduce a limited form of adaptivity by allowing the contractor to stop execution once future jobs cannot yield positive utility.
+Unfortunately, as we demonstrate, this additional adaptivity is insufficient to guarantee a constant-factor approximation.
+Finally, we consider a class of skipping-adaptive policies that augments a fixed priority order with job-specific time thresholds.
+A job is executed upon reaching its position in the order only if the remaining time before the deadline is at least its threshold; otherwise, it is skipped.
+We show this rather simple policy class is expressive enough to guarantee a constant-factor approximation.
+Application to Contract Design — Knapsack Contracts.
+Our initial motivation in studying the stochastic knapsack with costs problem was to study a variant of the principal-agent model, which we term Knapsack Contracts.
+In this model, the value of each executed task is fixed, but the duration of the task may depend on the effort level of the (privately taken) action of an agent.
+Moreover, the principal has a deadline by which they need to complete all tasks.
+Using contracts to incentivize timely execution of tasks is prevalent in many time-sensitive tasks such as food delivery and contractors outsourcing tasks.
+In Appendix B, we formally present the Knapsack Contracts model.
+We show that the Knapsack Contracts problem is reducible to the multi-choice stochastic knapsack with costs problem.
+Thus, providing approximation guarantees to the latter automatically translates to approximation guarantees for the former.
+Proposition 1.1.
+[Reduction to Multi-choice Stochastic Knapsack with Costs] Given an efficient algorithm that finds a policy that $c$-approximates the optimal policy for the MSKC problem, one can efficiently find a policy that $c$-approximates the optimal policy for the Knapsack Contracts problem.
+1.1 Our Contributions and Techniques
+In Section 2.1, we formally define the problem through which we explore the simplicity-optimality landscape when introducing costs, the Multi-choice Stochastic Knapsack with Costs problem (or MSKC).
+In this problem, each job is associated with multiple levels of execution, where each level has a different execution cost and induces a different size distribution.
+The decision-maker computes a policy to choose the (job, level) execution order, obtaining a utility of the value of jobs completed by the deadline minus the cost of all executed jobs.
+In Section 2.2 we formally define the different policies in our adaptiveness hierarchy.
+As mentioned above, hard instances are ones where for some jobs, their cost almost “covers” their value.
+Thus, when executing such jobs, the policy is left with a small return on investment (ROI).
+In order to parameterize this, we inspect the Inverse of the ROI (or IOR) of a job (as the ROI diminishes, the IOR grows).
 We introduce an economically-meaningful parameter $\alpha$, which represents the worst IOR among all jobs—see Definition 2.1.
-Our main finding is that IOR is precisely the parameter that determines the loss in approximation due to moving from the classic Stochastic Knapsack problem, to a contract design problem subject to incentive constraints:
-  • We show that the principal can efficiently and non-adaptively compute a set of contracts, which give an $O(\alpha)$-approximation to the profit of the (intractable and highly adaptive) optimal policy of the principal.
-See Theorem 3.1.
-  • We show the existence of an instance such that any non-adaptive policy of the principal is bounded away by a factor of $\Omega(\alpha)$ in terms of profit from the optimal adaptive policy.
-This establishes a tight $\Theta(\alpha)$ adaptivity gap for the Knapsack Contracts problem.
-See Theorem 4.1.
-  • We establish an information gap: Consider any policy (adaptive or not) that has limited information about the jobs’ size distributions in the form of finitely many moments, as well as access to the probability of knapsack overflow at every step.
-No such policy can guarantee an approximation factor better than $O(\alpha)$ with respect to the profit of the optimal policy with full access to the size distributions.
-See Theorem 5.2.
-We now describe our results in more detail as well as the techniques that enable them.
-Positive results.
-In Section 3, we provide an efficient algorithm that finds a non-adaptive set of contracts (see Section 2 for adaptivity notions), which give an $O(\alpha)$-approximation to the optimal profit the principal can obtain, where $\alpha$ is the IOR parameter (Theorem 3.1).
-Our algorithm needs to choose in advance the order in which agents are to execute their jobs and which effort level to incentivize each agent.
-Each effort level induces a different expected transfer needed in order to incentivize it.
-As a first step, in Lemma 2.1, we reduce the Knapsack Contract problem to a variant of Stochastic Knapsack that we term the Multi-Choice Stochastic Knapsack Problem with Costs (MSKC).
-In this purely algorithmic problem, each job is assigned one of multiple combinations of cost and size distribution.
-When a policy attempts to execute a job, it has to pay the cost up front.
-Our algorithm formulates an LP (see Eq.
-(2)) for the MSKC problem, where the weight of each job $i$ and effort level $j$ corresponds to the profit of executing job $i$ at effort level $j$, and the size is the expected size of the job at effort level $j$.
+We show this parameter is closely tied to the performance of non-adaptive policies in our problem.
+An overview of our main findings is given in Figure 1.
+[…]
+Figure 1: Hierarchy of policy adaptivity levels showing the approximation gaps between different classes.
+1.1.1 $\Theta(\alpha)$ Adaptivity Gap.
+In Section 3, we study the power of non-adaptive policies. Our results are the following.
+Upper Bound.
+We give an efficient algorithm for computing an ordered set of (job, execution level) pairs that give an $O(\alpha)$-approximation to the utility of the (intractable and highly adaptive) optimal policy (theorem 3.2).
+Our algorithm formulates an LP for the MSKC problem where each pair $(i,j)$ has weight equal to the utility from executing job $i$ at execution level $j$, and size equal to the expected size of job $i$ at execution level $j$.
 In Dean et al.
-[11], an LP formulation of the problem gives an upper bound on the reward of the optimal adaptive policy, which is approximated by the reward of the best non-adaptive policy.
-In our case, the LP is still an upper bound on the optimal profit of an adaptive policy, but this upper bound can be a factor $\alpha$ away (see Appendix A for an example).
-We show that there exists a non-adaptive set of contracts that yields an $O(\alpha)$-approximation to the LP’s value.
-The way in which we construct the set of contracts is the following.
-We first efficiently find a well-structured solution to the LP where each job is taken integrally at up to one effort level, except for at most a single job that can be taken as a convex combination of two effort levels.
-This is done by taking an arbitrary solution of the LP, and then shifting weights along the Pareto curves of jobs in order to iteratively move jobs to be taken integrally, while keeping the same solution value.
-Once we obtain the well-structured solution, we either use a standard bang-per-buck ordering of the jobs selected, or take a single job if its expected profit is larger than a predefined threshold.
-This threshold is set in a way that will (i) prevent a large loss for the bang-per-buck order in case the IOR parameter $\alpha$ has a large value; and (ii) ensure a large profit when there’s a single job that gives such profit, but might have a low probability of being selected when jobs are executed in the bang-per-buck order.
-Negative results.
-In Section 4 and Section 5 we give further evidence for the role that IOR has in Knapsack Contracts. In Section 4 we inspect the adaptivity gap, considering three types of policies:
-  1. 1.
-    Fully adaptive policies that can choose the next job to execute (if any) and at which effort level based on the entire history.
-  2. 2.
-    Stopping-time adaptive policies that chooses a fixed set of contracts (job/effort level pairs) up front at a fixed order, but stops once any future sequence of jobs has a negative expected profit.
-These types of policies avoid a large loss that the principal can incur due to the transfer component of the contract problem and are used in the algorithm of Section 6.
-  3. 3.
-    Non-adaptive policies that chooses a fixed set of contracts at a fixed order, and the process terminates either when the last job in the order has been executed, or when there’s a job that is finished after the deadline.
-Dean et al.
-[11] show that in the setting without costs, the reward of the optimal non-adaptive policy is within a factor 4 of the optimal adaptive policy.
-We show that when introducing costs, the ratio between these can be unbounded.
-Namely, for every IOR parameter $\alpha>1$, we show:
-  • An instance where the ratio between the profit of stopping-time adaptive policies and non-adaptive policies is $\Omega(\alpha)$ (Section 4.1).
-  • An instance where the ratio between the profit of fully adaptive policies and stopping-time adaptive policies is $\Omega(\alpha^{1/3})$ (Section 4.2).
-Our instances are ones where each job has only a single effort level, demonstrating that the cost aspect by itself leads to a significant change relative to the setting without cost.
-Information barrier.
-In Section 5 we inspect how much information is needed to get an $O(\alpha)$ approximation in the purely algorithmic Stochastic Knapsack problem where there is an inherent cost for executing jobs (this corresponds to having a single effort level for each agent).
-In the purely algorithmic setting, our algorithm from Section 3 only uses first moment information (the expectation) of jobs’ sizes, along with the probability each job’s size will exceed the deadline, and obtains an $O(\alpha)$ approximation.
-We show that this information does not suffice if one wants to obtain a better approximation.
-Namely, for every finite $k$, there exists an instance where any (possibly adaptive) policy that can only access the $k$ moments of the size distribution, along with the probability that each job will exceed the deadline, is bounded away by a factor of $\Omega(\alpha)$ from the optimal policy (Theorem 5.2).
-Bounded-size jobs.
-In Section 6, we consider the case where the IOR parameter can be arbitrarily large, but each job’s probability to take more than $1-\delta$ of the length of the time budget is small for $\delta\in(0,1)$.
-We give an efficient algorithm that finds a stopping-time adaptive set of contracts, which guarantees an $O(1/\delta)$ approximation to the optimal profit of a fully-adaptive set of contracts.
-In order to obtain this result, we first consider a variant of the problem where, when a job finishes after the deadline, the principal still obtains its reward, and then the process stops.
-For this variant, the algorithm from Section 3 gives a constant approximation to the LP value.
-We then consider running this algorithm while imagining the deadline is $\delta$-fraction of the actual deadline.
-We show that in this case, having bounded-size jobs is equivalent to obtaining the reward of the first job that finishes after the deadline.
-Moreover, halving the deadline only causes a factor $\delta$ loss in the objective.
-Note that (moderate) adaptivity is required for this variant as well, as non-adaptive policies lose a factor that grows with $\alpha$ (see Appendix B).
-Ordered-Adaptive Contracts.
-Motivated by real-world examples where jobs must be executed in a given order, in Appendix D we study the model where the order is given, but for each job in the order, the principal can decide whether to execute it, and if so, at which effort level.
-In Theorem D.1 we give a resource augmentation-type result: given a time budget larger by a factor of $1+\epsilon$ from the original time budget, one can compute an adaptive policy efficiently (in polynomial time in the number of jobs, effort levels, and $1/\epsilon$), which obtains a reward larger than the optimal adaptive policy for the original time budget.
-The policy, similar to [11, 5], defines a dynamic programming solution on carefully discretized rounded-down processing times.
-This result hints that our problem boils down to ordering the jobs; i.e., one may hope to find a nearly-optimal adaptive policy for our non-ordered model if one can efficiently find a near-optimal ordering of the jobs.
-1.3 Related Work
-The study of sequential contracts in [14] investigates the problem of incentivizing a single agent to perform multiple jobs in a sequential manner.
-In contrast, our work considers a sequence of contracts in which the jobs are executed sequentially by different agents, with no interdependence between the jobs.
-The works of [26, 15, 1, 16] consider contract design subject to a limited monetary budget constraining the transfers from principal to agent.
-In our work, we study a multi-agent setting under a time budget constraint on the agents’ actions, where time is limited and the principal must incentivize the agents to complete their jobs on time.
-Classic contract theory.
-Classic contract theory has a long-standing tradition in economics, focusing on incentive design under asymmetric information.
-A result by Carroll [7] shows that linear contracts are often robust across a range of environments, providing a foundation for the study of simpler contract forms.
-These classical models emphasize moral hazard, adverse selection, and trade-offs between risk and incentives.
-Algorithmic contract theory.
-In recent years, a series of studies has emerged that study contracts through the lens of the theory of computation, initiated by [12, 4].
-The authors of [12] raise concerns regarding the complexity and unintuitiveness of optimal contracts, while exploring the approximation ratios of simpler contract forms, such as linear contracts.
-The multi-agent setting was initially explored by Babaioff et al.
-[4].
-A recent survey paper [13] provides an overview of developments in algorithmic contract theory.
+[10], a similar LP upper-bounds the utility of the optimal adaptive policy and can be approximated via the best non-adaptive policy.
+In our setting, the LP remains an upper bound on the optimal adaptive utility, but it may be a factor $\Omega(\alpha)$ away (see Section D.3).
+We show that some non-adaptive ordered set of job-level pairs achieves an $O(\alpha)$-approximation to the LP value.
+We construct this set as follows.
+First, we efficiently obtain a well-structured LP solution in which every job is assigned integrally to at most one execution level, except possibly one job that may be split as a convex combination of two execution levels.
+We start from an arbitrary LP solution and shift weight along each job’s Pareto curve, iteratively making jobs integral while preserving the objective value.
+Given the resulting solution, we either apply the standard bang-per-buck ordering to the selected jobs, or take a single job if its expected utility exceeds a fixed threshold.
+The threshold is chosen to balance the potential loss when overflowing and the utility when selecting a high-utility job that is not part of the LP solution.
+Lower Bound.
+We show the existence of an instance such that any non-adaptive policy of the principal is bounded away by a factor of $\Omega(\alpha)$ in terms of utility from the optimal adaptive policy.
+This establishes a tight $\Theta(\alpha)$ adaptivity gap for the Knapsack Contracts problem (theorem 3.1).
+In fact, in the given instance, the gap exists between non-adaptive and stopping-adaptive policies.
+1.1.2 A Constant Approximation with Limited Adaptivity
+Motivated by the dependency of non-adaptive policies on ROI, in Section 4, we consider policies that go beyond non-adaptive policies, but still exhibit limited adaptivity.
+Our goal is to devise simple near-optimal policies, irrespective of the IOR parameter.
+In Section 4.1 we first inspect whether policies that take a fixed, non-adaptive execution order, but can adaptively halt execution when future jobs are non-profitable, approximate the optimal utility.
+This is arguably the simplest class of policies that goes beyond non-adaptive policies, and can avoid incurring a high cost if the remaining time is insufficient.
+We show that this does not suffice; namely, we provide an instance where the ratio between the optimal utility and the utility of stopping-adaptive policies is $\Omega(\alpha^{1/3})$.
+In this instance, a good order of execution is obvious, but an optimal policy can skip over jobs that yield negative utility.
+This motivates us to study the next class of policies.
+In Section 4.2 we give an efficient algorithm to compute a skipping-adaptive policy that gives a constant approximation to the utility of the optimal fully-adaptive policy.
+The policy is computed via a global time-indexed LP, an approach taken in [20, 24, 3], adapted to our setting.
+1.1.3 Non-Stopping Policies
+In Appendix F, we study an even more demanding setup, where the policy chooses an ordered set of items in advance, and incurs the cost of all of them, not only items that are inserted before the capacity is met.
+In the motivating example, this corresponds to a contractor who pre-commits to a fixed slate of subcontractors and must pay all agreed costs, even if earlier overruns make later jobs unprofitable or irrelevant.
+In the setting without costs (e.g., [10]), non-stopping and non-adaptive policies coincide.
+With costs, however, we show an $\Omega(\alpha)$ separation between them, and give an $O(\alpha)$-approximation non-stopping algorithm for the fully-adaptive optimum.
+1.1.4 Information Gap
+Our algorithm for computing a constant-approximation policy uses a substantial amount of distributional information, while our non-adaptive algorithm only uses the first moments of the size distributions and the overflow probability.
+In Appendix C, we establish an information gap: Consider any policy (adaptive or not) that has limited information about the jobs’ size distributions in the form of finitely many moments, as well as access to the probability of knapsack overflow for a single item.
+No such policy can guarantee an approximation factor better than $O(\alpha)$ with respect to the utility of the optimal policy with full access to the size distributions (theorem C.2).
+We note that the $O(\alpha)$ non-adaptive policy only uses the expectation (first moment information) and probability of overflow.
+1.1.5 Relaxations.
+In Appendix G, we study mild relaxations under which simpler policies achieve a constant-factor approximation.
+In Section G.2, we allow the IOR parameter to be arbitrarily large, but assume that for each job, the probability of consuming more than $1-\delta$ of the budget is small, for $\delta\in(0,1)$.
+We give an efficient algorithm that outputs a stopping-adaptive policy with an $O(1/\delta)$-approximation to the optimal fully-adaptive utility.
+We also show that non-adaptive policies fail to achieve an $\alpha$-independent approximation, even when job sizes are bounded by half the total time.
+To prove this, we first study a variant where if a job completes after the deadline, the policy still receives its value and the process stops.
+For this variant, the algorithm from Section 3.2 gives a constant approximation to the LP value.
+We then run the same algorithm while treating the deadline as a $\delta$-fraction of the true deadline.
+In this view, bounded-size jobs correspond exactly to collecting the value of the first job that finishes after the (imagined) deadline, and shrinking the deadline costs only a factor $\delta$ in the objective.
 2 Model
-In Section 2.1 we present our contract setting; we reduce this setting to a purely combinatorial setting in Section 2.2.
-2.1 Contract Design Setting
-An instance $I$ of the knapsack contracts problem consists of a principal, $n$ agents, and a knapsack of size $S$.
-Without loss of generality, let $S=1$, where $S$ represents a budget of some limited resource.
-As our running example, we assume $S$ is the time remaining until a set deadline.
-Each agent $i$ has a job which they can carry out to the benefit of the principal.
-Agent $i$ is represented by the following:
-  • $v_{i}$: The reward which the principal obtains if agent $i$ completes their job before the deadline.
-  • $\{(F_{i}^{j},c_{i}^{j})\}_{j\in[m]}$: The possible effort levels of agent ii.
-[^1] All agents have the same number of effort levels without loss of generality, since we can add “dummy” effort levels that would never be incentivized by the principal.
-Each effort level $j\in[m]$ determines a distribution $F_{i}^{j}$ over the time it takes the agent to complete the job, as well as a cost $c_{i}^{j}\geq 0$.
-If the agent chooses to exert effort level $j$, they incur cost $c_{i}^{j}$, and complete their job within time $s_{i}$ drawn (independently) from distribution $F_{i}^{j}$.
-We assume non-decreasing costs $c_{i}^{1}\leq\dots\leq c_{i}^{m}$.
-Let $\mathbf{F}_{i}=\left(F_{i}^{1},\ldots,F_{i}^{m}\right)$.
-Let $\mathrm{supp}(F_{i}^{j})$ denote the support of $F_{i}^{j}$, and $\mathrm{supp}(\mathbf{F}_{i})=\cup_{j\in[m]}\mathrm{supp}(F_{i}^{j}).$ We assume the support of each distribution is finite and is explicitly given.
-The principal knows $I=(v_{i},\{F_{i}^{j},c_{i}^{j}\}_{j\in[m]})_{i\in[n]}$ in advance.
-After an agent completes a job, the principal observes the time it took the agent to complete it, but it does not observe the effort level chosen by the agent.
-The principal offers each agent $i$ a contract $t_{i}:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$, mapping $i$’s observed processing time to a transfer paid by the principal to the agent.
-Given contract $t_{i}$, agent $i$ chooses an effort level
-$j\in\arg\max_{\ell}\operatorname{\mathbb{E}}_{s_{i}\sim F_{i}^{\ell}}[t_{i}(s_{i})]-c_{i}^{\ell}.$
-Definition 2.1 (Implementable effort level).
-For agent $i$, an effort level $j$ is implementable if there exists a contract $t_{i}$ for which this action maximizes the agent’s expected utility among all actions given the contract.
-In this case, we say contract $t_{i}$ incentivizes action $j$.
-Any implementable effort level $j$ for agent $i$ has an optimal contract that incentivizes it, which can be obtained by solving the following LP:
-$\min\quad$ $\sum_{s\in\mathrm{supp}(F_{i}^{j})}t_{i}(s)\cdot F_{i}^{j}(s)$
-s.t.
-$\sum_{s\in\mathrm{supp}(F_{i}^{j})}t_{i}(s)\cdot F_{i}^{j}(s)-c_{i}^{j}\geq\sum_{s\in\mathrm{supp}(F_{i}^{j^{\prime}})}t_{i}(s)\cdot F_{i}^{j^{\prime}}(s)-c_{i}^{j^{\prime}}$ $\forall j^{\prime}\in[m]$
-$t_{i}(s)\geq t_{i}(s^{\prime})$ $\forall s<s^{\prime}\in\mathrm{supp}(\mathbf{F}_{i})$ (1)
-$t_{i}(s)\geq 0$ $\forall s\in\mathrm{supp}(\mathbf{F}_{i})$
-Constraint (1) ensures monotonicity of payments. This is required as without it, an agent might have an incentive to hover around after completing the job and get a higher payment.
-The principal’s problem can then be described as: when hiring agent $i$, the principal chooses an effort level $j$ to incentivize, and offers the optimal contract that incentivizes that effort level.
-The principal’s collected reward is the sum of the rewards of the jobs that the agents get to complete with a total processing time less than the budget $S$.
-The principal aims to maximize their profit: the expected total reward minus the expected sum of transfers made to the agents.
-Adaptive versus non-adaptive policies.
+In this section we introduce our problem, the Inverse ROI parameter, and our policy hierarchy.
+2.1 The Multi-choice Stochastic Knapsack with Costs Problem (MSKC)
+We generalize Stochastic Knapsack along two axes: (i) items can have multiple tiers, called choices, and (ii) each choice is associated with a cost.
+An instance $I=\{v_{i},\{p_{i}^{j},F_{i}^{j}\}_{j\in[m]}\}_{i\in[n]}$ of the multi-choice stochastic knapsack problem with costs (or MSKC for short), consists of $n$ items.
+Each item $i\in[n]$ has a value $v_{i}$ and $m$ choices.
+Each choice $j\in[m]$ has a (deterministic) cost $p_{i}^{j}$, and a size distribution $F_{i}^{j}$.
+There is also a knapsack capacity of size $S$, but without loss of generality, $S=1$.
+When item $i$ is selected for the knapsack, a single choice $j$ must be specified.
+We refer to these as the item-choice pair (sometimes referring just to the item when the choice is clear from context).
+In this case, the cost $p_{i}^{j}$ is incurred up-front, and the item’s size $s_{i}$ is realized from distribution $F_{i}^{j}$.
+Item $i$’s realized utility is then $v_{i}-p_{i}^{j}$ (more precisely, this is the “potential” realized utility, i.e., the utility that will be gained if the item fits in the knapsack).
+For a given instance $I$, a policy $\varphi$ possibly inspects the knapsack, then either selects an item-choice pair, or terminates.
+Item-choice pairs are selected according to the policy sequentially, until the total size of selected items exceeds the capacity, or until the policy decides to terminate.
+In the former case, we refer to the last selected item as the overflow item.
+The realized utility of the policy is the total realized utility of items selected before the cumulative size exceeded the capacity, minus the cost of the overflow item, if one exists.
+Let $\mathrm{utility}(\varphi)$ be a random variable representing the utility obtained by the policy $\varphi$ on instance $I$.
+We also design algorithms that return policies.
+When referring to an algorithm $\mathcal{A}$ that is given an instance $I$ and returns a policy $\varphi_{I}$, we denote by $\mathcal{A}(I)=\mathrm{utility}(\varphi_{I})$ the random variable representing the utility obtained by the policy on instance $I$.
+We are also interested in the expected utility of a pair $(i,j)$ of item and its choice.
+This depends on the remaining capacity $t\in[0,1]$, and is equal to $v_{i}\cdot\Pr_{s_{i}\sim F_{i}^{j}}[s_{i}\leq t]\;-\;p_{i}^{j}.$ Let
+$w_{i,j}:=v_{i}\cdot\Pr_{s_{i}\sim F_{i}^{j}}[s_{i}\leq 1]-p_{i}^{j}$
+be the expected utility of inserting $(i,j)$ when the knapsack is empty ($t=1$).
+The lower bounds in this paper critically exploit instances with item-choice pairs $(i,j)$, which have high values and high costs, but only low expected utilities:
+$0<w_{i,j}\ll\min\{v_{i},p_{i}^{j}\}.$
+In such cases, an item-choice pair may yield a small positive expected utility when the knapsack is empty, yet incur a large negative expected utility, of up to $p_{i}^{j}$, when the capacity becomes limited.
+We show that this drastic instability property is an inherent obstacle in our model.
+To do so, we define the Inverse of the Return on Investment parameter:
+Definition 2.1 (ROI and $\alpha$ Parameters).
+We define
+$\mathrm{ROI}(i,j)\ :=\ w_{i,j}/p_{i}^{j},~~~\alpha\ :=\ \max_{i,j}\{1/\mathrm{ROI}(i,j)\}.$
+We also refer to $\alpha$ as the Inverse of ROI (IOR) parameter.
+Intuitively, a large value of $\alpha$ signifies the presence of at least one item-choice pair where the upfront cost of the item greatly exceeds the utility from successfully inserting the item.
+Our results show that $\alpha$ is the key parameter governing the adaptivity gaps between different levels of adaptivity, which we now specify.
+2.2 Adaptiveness Hierarchy
+For a fixed instance, recall that a policy is applied iteratively, possibly observing the state of the knapsack in each iteration, in order to sequentially fill the knapsack.
+After the formal definitions we give an informal interpretation.
 Definition 2.2.
-(Fully-Adaptive Policies).
-A fully-adaptive policy, $\varphi:2^{I}\times[0,1]\rightarrow[n]\times[m]\cup\{\perp\}$, takes as input the unassigned jobs $J\subseteq I$ and the remaining time budget $S^{\prime}\leq 1$, and chooses which agent to hire next and at which implementable effort level.
-This procedure is repeated until:
-  • an agent finishes their job after the deadline; or
-  • the policy decides not to execute another job (outputs $\perp$ as the next job).
-Let $\mathrm{profit}(\varphi)$ be a random variable representing the profit obtained by the principal when using policy $\varphi$ on instance $I$.
-We denote by $\mathrm{ADAPT}(I)=\max_{\varphi}\operatorname{\mathbb{E}}[\mathrm{profit}(\varphi)]$ the optimum expected profit obtained by an adaptive policy for instance $I$.
-When referring to an algorithm $\mathcal{A}$ that is given an instance $I$ and returns a policy $\varphi_{I}$, we denote by $\mathcal{A}(I)=\mathrm{profit}(\varphi_{I})$ the random variable representing the profit obtained by the principal when using the policy outputted by algorithm $\mathcal{A}$’s policy on instance $I$.
+(Non-Adaptive Policy) A non-adaptive policy $\varphi:[n]\rightarrow[n]\times[m]\cup\{\perp\}$ takes as input the iteration number, and returns an item-choice pair, or $\perp$ for termination.
+Let $\mathrm{NON\_ADAPT}(I)$ be the optimum expected utility obtained by a non-adaptive policy for instance $I$.
 Definition 2.3.
-(Non-Adaptive Policies) A non-adaptive policy is an ordering of agents $\pi:[n]\rightarrow[n]\cup\{\perp\}\times[m]$.
-For some $i\in[n]$, $\pi(i)$ outputs the $i$th selected agent and the effort level the policy incentivizes the agent to exert.
-Let $\Pi$ be the set of non-adaptive policies.
-We denote by $\mathrm{NON\_ADAPT}(I)=\max_{\pi\in\Pi}\operatorname{\mathbb{E}}[\mathrm{profit}(\pi)]$ the optimum expected profit obtained by a nonadaptive policy for instance $I$.
+(Stopping-Adaptive Policy) A stopping-adaptive policy $\pi:[n]\times[0,1]\rightarrow[n]\times[m]\cup\{\perp\}$ takes as input the iteration number and the remaining knapsack capacity, and returns an item-choice pair, or $\perp$ for termination.
+Let $\mathrm{STOP\_ADAPT}(I)$ be the optimum expected utility obtained by a stopping-adaptive policy for instance $I$.
 Definition 2.4.
-(Stopping-Time Adaptive Policies) Similarly to non-adaptive policies, a stopping-time adaptive policy chooses a non-adaptive ordering of the agents and their effort levels $\pi$.
-However, it may adaptively choose to stop at any point.
-Without loss of generality, we assume the policy stops if every remaining sequence of agents yields a negative expected profit.
-Let $\Pi^{\prime}$ be the set of stopping-time adaptive policies.
-$\mathrm{STOP\_ADAPT}(I)=\max_{\pi\in\Pi^{\prime}}\operatorname{\mathbb{E}}[\mathrm{profit}(\pi)]$ is the optimum expected profit obtained by a stopping-time policy for instance $I$.
-2.2 Multi-Choice Stochastic Knapsack Problem With Costs
-We define a generalization of Stochastic Knapsack, where items have multiple choices of assignment, and each is associated with a size distribution, and a cost.
-We show that our Knapsack Contracts problem reduces to this problem.
-This implies that our algorithmic results can be obtained in this setting, and the results will carry through to the contracts setting.
-Definition 2.5 (Multi-Choice Stochastic Knapsack Problem With Costs).
-An instance of the multi-choice stochastic knapsack problem with costs (or MSKC for short) consists of $n$ items, where each item has $m$ choices, each choice has a different cost and size distribution.
-That is $I=\{v_{i},\{p_{i}^{j},F_{i}^{j}\}_{j\in[m]}\}_{i\in[n]}.$ An item $i$ can be selected with a single choice $j$.
-In this case, the cost $p_{i}^{j}$ must be incurred up front, and the item’s size is realized from distribution $F_{i}^{j}$.
-There is a budget of size $S$ ($S=1$ without loss of generality).
-A policy selects item-choice pairs sequentially, and terminates either when it decides not to select any item, or when the total size of selected items exceeds the budget.
-The profit of a policy is the total value of items selected that do not exceed the budget minus the total cost of the selected items, including the cost of an item that exceeds the budget, if one exists.
-Adaptive, Stopping-time Adaptive and Non-adaptive policies for the MSKC problem are defined analogously to the way they are defined for the Knapsack Contracts problem.
-The following lemma (Proof in Appendix F) shows that in order to find an approximately optimal policy for the Knapsack Contracts problem, it suffices to find an approximately optimal policy for the MSKC problem.
-Lemma 2.1 (Reduction to Multi-Choice Stochastic Knapsack with Costs).
-Given an efficient algorithm that finds a policy that $c$-approximates the optimal policy MSKC problem, one can efficiently find a policy that $c$-approximates the optimal policy for the knapsack contracts problem.
-2.3 Inverse of Return on Investment Parameter
-Consider an instance of the Knapsack Contracts problem. For a job and implementable effort level pair, $(i,j)$, let
-$\tilde{t}_{i}^{j}=\operatorname{\mathbb{E}}_{s\sim F_{i}^{j}}[t_{i}(s)],$
-where $t_{i}$ is the contract that minimizes the expected payment while incentivizing effort level $j$, as given by the LP in Definition 2.1.
-That is, $\tilde{t}_{i}^{j}$ is the minimal expected transfer required in order to incentivize agent $i$ to take action $j$.
-The hardness results presented in the paper critically exploit problem instances characterized by pairs (job, action) with high values $v_{i}$ and expected transfers $\tilde{t}_{i}^{j}$, but a comparatively low net profit
-$v_{i}\cdot\Pr_{s_{i}^{j}\sim F_{i}^{j}}[s_{i}^{j}\leq 1]-\tilde{t}_{i}^{j}.$
-In such cases, a job may yield a small positive profit of $v_{i}-\tilde{t}_{i}^{j}$ when the knapsack is empty, yet incur a large negative penalty of up to $\tilde{t}_{i}^{j}$ when capacity is low.
-Moreover, this fluctuation can be abrupt: even an $\epsilon$ increase in load may decrease the expected profit to $\tilde{t}_{i}^{j}$.
-We show that this drastic fluctuation property is an inherent obstacle in our model.
-To do so, we define the Inverse of the Return on Investment parameter:
-Definition 2.6 (Inverse of ROI parameter).
-We define
-$\mathrm{ROI}(i,j)\ =\ \frac{v_{i}\cdot\Pr_{s_{i}^{j}\sim F_{i}^{j}}[\,s_{i}^{j}\leq 1\,]\;-\;\tilde{t}_{i}^{j}}{\tilde{t}_{i}^{j}},$
-and
-$\alpha\;=\;\max_{i,j\in\mathsf{I}_{i}}\frac{1}{\mathrm{ROI}(i,j)}\,$,
-where $\mathsf{I}_{i}$ is the set of implementable actions for job $i$. We sometimes refer to $\alpha$ as the Inverse of ROI (or IOR) parameter.
-Intuitively, a large value of $\alpha$ signifies the presence of at least one job where for some effort level, the transfer required to incentivize it is much larger than the net profit from successfully executing the job before the deadline.
-We show a $\Theta(\alpha)$ adaptivity gap and an $\Omega(\alpha)$ information gap (Section 5), suggesting that $\alpha$ captures precisely the “hardness” aspect for our problem.
-Analogously, we define the IOR parameter for the MSKC problem as in Definition 2.6, but using $p_{i}^{j}$ instead of $\tilde{t}_{i}^{j}$.
-Given a Knapsack Contracts instance with an $\alpha$ IOR parameter, applying the reduction in Lemma 2.1 yields an MSKC instance with the same $\alpha$ as the analogous IOR parameter.
-This implies that any approximation as a function of $\alpha$ a policy in the MSKC instance would obtain will imply the same approximation as a function of $\alpha$ in the Stochastic Knapsack instance.
-3 $O(\alpha)$-Approximate Non-Adaptive Contracts
-In this section, we show how to find a non-adaptive set of contracts $\pi:[n]\to[n]\times[m]\cup\{\perp\}$ that obtains an $O(\alpha)$ approximation to the optimal profit the principal can obtain by any (possibly highly adaptive) policy.
-By Lemma 2.1, we focus on finding an $O(\alpha)$-approximate non-adaptive policy for the MSKC problem, which yields an $O(\alpha)$-approximate non-adaptive set of contracts for the Knapsack Contracts problem.
-Our main theorem is the following:
+(Skipping-Adaptive Policy) A skipping-adaptive policy $\pi:[n]\times[n]\times[0,1]\rightarrow[n]\times[m]\cup\{\perp\}$ takes as input the iteration number, the last item in the knapsack and the remaining knapsack capacity, and returns an item-choice pair, or $\perp$ for termination.
+Let $\mathrm{SKIP\_ADAPT}(I)$ be the optimum expected utility obtained by a skipping-adaptive policy for instance $I$.
+Definition 2.5.
+(Fully-Adaptive Policy).
+A fully-adaptive policy $\varphi:[n]\times 2^{[n]}\times[0,1]\rightarrow[n]\times[m]\cup\{\perp\}$ takes as input the iteration number, the item set in the knapsack and the remaining knapsack capacity, and returns an item-choice pair, or $\perp$ for termination.
+Let $\mathrm{ADAPT}(I):=\max_{\varphi}\operatorname{\mathbb{E}}[\mathrm{utility}(\varphi(I))]$ be the optimum expected utility obtained by a fully-adaptive policy for instance $I$.
+Intuitively, a non-adaptive policy is an ordering of items.
+For each $i\in[n]$, it specifies the $i$-th selected item and its associated choice, or $\perp$ for termination.
+A stopping-adaptive policy, like a non-adaptive one, is an ordering of items, but it can adaptively stop at any point.
+A skipping-adaptive policy is like a stopping-adaptive policy, with the additional ability to skip an item before stopping.
+It, too, is an ordering of items, and given the last inserted item it scans the successive ones in the ordering and chooses the next one to insert.
+[^1] Related to skipping-adaptive policies, we can consider an ordering of items and target capacities — the maximum remaining capacity with which each item will be entered into the knapsack.
+When given the last item inserted, the policy finds the subsequent item in the ordering, whose target capacity is below the remaining knapsack capacity.
+[10] study fully-adaptive (and non-adaptive) policies.
+We remark that each policy can be represented by a decision tree; a fully-adaptive policy needs an exponentially-sized tree, but the other policies require only poly-sized trees.
+In fact stopping-adaptive and non-adaptive policies correspond to linear-sized decision trees.
+3 Tight $\Theta(\alpha)$ Adaptivity Gap
+In this section we establish a tight $\Theta(\alpha)$ gap between simple non-adaptive policies and the optimal fully-adaptive policy.
+Theorem 3.1 constructs an instance with an $\Omega(\alpha)$ gap between the best non-adaptive policy and the best stopping-adaptive policy.
+Theorem 3.2 gives a polynomial-time algorithm that outputs a non-adaptive policy whose expected utility is within an $O(\alpha)$ factor of the optimal fully-adaptive utility.
+In Appendix F, we study an even more demanding setup, where the policy chooses an ordered set of jobs in advance, and incurs the cost of all of them, not only jobs that are inserted before the capacity is met.
+In the setting without costs (e.g., [10]), this setup is equivalent to a non-adaptive policy, as all costs are 0.
+In the MSKC Problem, we show a separation between the two.
+In Appendix F we define non-stopping policies, and show that there is an instance with an $\Omega(\alpha)$ gap between the optimal non-stopping policy and the optimal non-adaptive policy, and an $O(\alpha)$ approximation non-stopping algorithm to the optimal fully-adaptive policy.
+3.1 Lower Bound
+We show that there exists an $\Omega(\alpha)$ gap between the optimal Non-Adaptive policy and the optimal Stopping-Adaptive policy.
+Recall that both models require the algorithm to choose an item ordering at the beginning; the stopping-adaptive policy can then adaptively choose when to halt, while the non-adaptive policy must attempt all items in the chosen order, until all items are inserted or an overflow occurs.
 Theorem 3.1.
-There exists a polynomial time algorithm for MSKC that computes for every input $I$ a non-adaptive policy $\pi_{I}:[n]\to[n]\times[m]\cup\{\perp\}$ such that
-$\frac{\mathrm{ADAPT}(I)}{\operatorname{\mathbb{E}}[\mathrm{profit}(\pi_{I})]}\leq O(\alpha).$
+For every $\alpha$ there exists an instance $I$ of the MSKC problem such that $\frac{\mathrm{STOP\_ADAPT}(I)}{\mathrm{NON\_ADAPT}(I)}=\Omega(\alpha).$
+We give a proof sketch here and defer the full proof of theorem 3.1 to section D.1.
+Proof sketch.
+Fix parameters $\varepsilon,\gamma\in(0,\tfrac{1}{2})$ and let $n=\lceil 1/\varepsilon\rceil$.
+Consider capacity $1$ and $n$ items with value $v=2$ and cost $p=2-\varepsilon$, so each successful insertion yields net gain $w=v-p=\varepsilon$ and hence $\alpha=\max_{i}p/w=(2-\varepsilon)/\varepsilon=\Theta(1/\varepsilon)$.
+For item $i\in[n]$, define a two-point size distribution
+$s_{i}=\begin{cases}a_{i}&\text{with prob. }1-\varepsilon,\\ b_{i}&\text{with prob. }\varepsilon,\end{cases}\qquad a_{i}=\gamma^{\,n-i+1},\qquad b_{i}=1-\sum_{j<i}a_{j}.$
+Since $\gamma<\tfrac{1}{2}$, the sequence $(a_{i})$ grows geometrically and satisfies $a_{i^{\prime}}>\sum_{j<i}a_{j}$ for all $i<i^{\prime}$, implying $b_{i}+a_{i^{\prime}}>1$.
+Thus, once any earlier item realizes its “big” size $b_{i}$, every subsequent attempt overflows even if it realizes “small”.
+A stopping-adaptive policy that attempts items in order and halts after the first big realization inserts a geometric number $X$ of items (truncated at $n$) with
+$\mathbb{E}[X]=\sum_{t=1}^{n}(1-\varepsilon)^{t-1}=\frac{1-(1-\varepsilon)^{n}}{\varepsilon}\geq\frac{1-e^{-1}}{\varepsilon},$
+and therefore achieves expected utility at least $\mathbb{E}[X]\cdot\varepsilon\geq 1-e^{-1}=\Omega(1)$.
+In contrast, we show (in the full proof) that the expected utility of any non-adaptive policy that attempts inserting at least $k>1$ items has negative utility.
+Hence the optimal non-adaptive policy attempts only one item and attains $\mathrm{NON\_ADAPT}=\varepsilon$.
+Consequently,
+$\frac{\mathrm{STOP\_ADAPT}}{\mathrm{NON\_ADAPT}}\geq\frac{1-e^{-1}}{\varepsilon}=\Omega(\alpha),$
+as claimed. ∎
+3.2 Upper Bound
+In this section, we present an $O(\alpha)$-approximate non-adaptive policy for the MSKC problem. Our main theorem is the following:
+Theorem 3.2.
+There exists a polynomial time algorithm for MSKC that computes for every input $I$ a non-adaptive policy $\pi_{I}:[n]\to[n]\times[m]\cup\{\perp\}$ such that $\frac{\mathrm{ADAPT}(I)}{\operatorname{\mathbb{E}}[\mathrm{utility}(\pi_{I})]}\leq O(\alpha).$
 We first define an LP that both guides our solution and gives an upper bound on the performance of every adaptive policy.
 For every $i\in[n],j\in[m]$ and for a knapsack of size $t$, denote by $w_{ij}=v_{i}\cdot\Pr_{{s_{i}^{j}\sim F_{i}^{j}}}[s_{i}^{j}\leq t]-p_{i}^{j}$ and $\mu_{ij}=\operatorname{\mathbb{E}}_{s_{i}^{j}\sim F_{i}^{j}}[\min\{s_{i}^{j},t\}]$.
-The first parameter can be viewed as a proxy to the profit from job $i$ when executed at effort level $j$ and the second parameter is the expected truncated size.
+The first parameter can be viewed as a proxy to the utility from item $i$ inserted with choice $j$ and the second parameter is the expected truncated size.
 Consider the following LP, parameterized by the size of the knapsack $t$.
@@
 Note that every solution returned by the Simplex algorithm satisfies the above properties; the basic feasible solution of the LP has at most $n+1$ non-negative variables, and the constraints ensure that for each item $i$, there is at least one choice $j$ with $x_{ij}>0$, hence at most one item has two non-zero choices.
-However, the Simplex algorithm may run at exponential time in the worst case.
+However, the Simplex algorithm may run in exponential time in the worst case.
 Thus, we describe a polynomial time algorithm to compute such a solution.
@@
 For an item $i$, choice $j^{\prime}$ single dominates choice $j$ if:
-$(1)$ $\quad w_{ij^{\prime}}\geq w_{ij}$
-$(2)$ $\quad\mu_{ij^{\prime}}\leq\mu_{ij}$
+$(1)\quad w_{ij^{\prime}}\geq w_{ij}\qquad(2)\quad\mu_{ij^{\prime}}\leq\mu_{ij}$
 Definition 3.2 (Double Dominance).
@@
 Suppose that for some feasible solution $x$, we have $x_{ij}\neq 0$.
-First, if $j^{\prime}$ single-dominates $j$, define a new solution by setting
-$x^{*}_{ij^{\prime}}=x_{ij^{\prime}}+x_{ij},\quad x^{*}_{ij}=0~.$
+First, if $j^{\prime}$ single-dominates $j$, define a new solution by setting $x^{*}_{ij^{\prime}}=x_{ij^{\prime}}+x_{ij},\ x^{*}_{ij}=0.$
 This yields a feasible solution whose objective value is at least as large as that of $x$. Hence, there exists an optimal solution in which $x_{ij}=0$.
@@
 This again yields a feasible solution whose objective value is at least as large as that of $x$ (see fig.
-1 for illustration, and Appendix E for the full proof).
+2 for illustration, and section D.2 for the full proof).
 Therefore, in either case, there exists an optimal solution in which $x_{ij}=0$.
@@
 Lemma 3.5.
-There exists an optimal solution $x$ to $\Phi(t)$ in which, for every item $i$, at most two choices satisfy $x_{ij}>0$.
+There exists
+an optimal solution $x$ to $\Phi(t)$ in which, for every item $i$, at most two choices satisfy $x_{ij}>0$.
 Proof.
-Let $x$ be an optimal solution in which no choice is dominated; such a solution exists by Lemma 3.3. Without loss of generality we may further assume that no additional choices remain and that
-$w_{i,j}\;<\;w_{i,j+1}\qquad\text{for every item }i\text{ and every }j\in\{1,\dots,m-1\}.$
-Lemma 3.3 then also yields
-$\mu_{i,j}\;<\;\mu_{i,j+1}\qquad\text{for all }i\text{ and }j\in\{1,\dots,m-1\}.$
+Let $x$ be an optimal solution in which no choice is dominated; such a solution exists by Lemma 3.4.
+Without loss of generality, we may further assume that no additional choices remain and that for every $i\in[n]$ and $j\in[m-1]$, $w_{i,j}<w_{i,j+1}$.
+Lemma 3.4 then also yields $\mu_{i,j}<\mu_{i,j+1}$.
 Consider the pairs $\{\bigl(w_{ij},\mu_{ij}\bigr)\}_{j\in[m]}$ for a fixed index $i$.
@@
 Proof of Lemma 3.3.
-By Lemmas 3.3, 3.4, it follows that there exists an optimal solution $x$ where at most two choices have positive values.
+By Lemmas 3.4, 3.5, it follows that there exists an optimal solution $x$ where at most two choices have positive values.
 Consider a solution where there exist indices $i_{1},i_{2}$ and $j^{\prime}_{1},j^{\prime\prime}_{1},j^{\prime}_{2},j^{\prime\prime}_{2}$ such that $x_{i_{k},j^{\prime}_{k}}>0$ and $x_{i_{k},j^{\prime\prime}_{k}}>0$ for each $k\in\{1,2\}$.
@@
 The solution for each item is a convex combination of the two choices, corresponding to the red dots in the graphs $\Gamma_{i_{1}}$ and $\Gamma_{i_{2}}$, which lie on the segments connecting the two choices.
-As the slope of the first item is larger, the modified solution indicated by the dashed arrows illustrates one possible update that maintain feasibility and increases the objective value.
+As the slope of the first item is larger, the modified solution indicated by the dashed arrows illustrates one possible update that maintains feasibility and increases the objective value.
 For a set of item choices $\mathcal{S}\subset[n]\times[m]$, denote $size(\mathcal{S})=\sum_{(i,j)\in\mathcal{S}}s_{i}^{j}$ and $\mu(S)=\sum_{(i,j)\in\mathcal{S}}\mu_{ij}$.
 The following Lemmas are direct extensions of key lemmas from Dean et al.
-[11], originally established for the classic Stochastic Knapsack problem with a budget of 1.
-We adapt these lemmas to our knapsack variant with multi-choice and costs, and for a general budget $t$.
-In this section, we only need the case where $t=1$; however, in Section 6 we use the generalized versions.
-The proofs are presented in Appendix E for completeness.
+[10], originally established for the classic Stochastic Knapsack problem with a capacity of 1.
+We adapt these lemmas to our knapsack variant with multi-choice and costs, and for a general capacity $t$.
+In this section, we only need the case where $t=1$; however, in appendix G we use the generalized versions.
+The proofs are presented in Appendix D for completeness.
 Lemma 3.6.
-For any set of item choices $\mathcal{S}$:
-$\Pr[size(\mathcal{S})<t]\geq 1-\frac{1}{t}\mu(\mathcal{S}).$
+For any set of item choices $\mathcal{S}$, $\Pr[size(\mathcal{S})<t]\geq 1-\frac{1}{t}\mu(\mathcal{S}).$
 Lemma 3.7.
-For any adaptive policy for the MSKC problem with budget $t$, let $\mathcal{S}$ be the random set of item choices that the policy inserts, then $\operatorname{\mathbb{E}}[\mu(\mathcal{S})]\leq 2t$.
+For any adaptive policy for the MSKC problem with capacity $t$, let $\mathcal{S}$ be the random set of item choices that the policy inserts, then $\operatorname{\mathbb{E}}[\mu(\mathcal{S})]\leq 2t$.
 Lemma 3.8.
-For any constant $c\geq 2$ and budget $t$,
-$\mathrm{ADAPT}\leq\Phi(ct)\leq c\cdot\Phi(t)~,$
-where $\mathrm{ADAPT}$ is the optimal adaptive policy for budget $t$.
-We are now ready to present our algorithm, SKC (Stochastic Knapsack with Costs):
+For any constant $c\geq 2$ and capacity $t$, $\mathrm{ADAPT}\leq\Phi(ct)\leq c\cdot\Phi(t),$ where $\mathrm{ADAPT}$ is the optimal adaptive policy for capacity $t$.
+We now present our algorithm, NA-SKC (Non-Adaptive Stochastic Knapsack with Costs):
 Input: Instance II with items $i\in[n]$, choices $j\in[m]$, values $v_{i}$, costs $p_{i}^{j}$, distributions $F_{i}^{j}$
@@
 // Step 1: LP Solution
-Solve linear program $\Phi(1)$ to obtain optimal solution $x$ with the structure from Lemma 3.2;
-Let $\mathcal{C}=\{(i,j)\mid x_{ij}>0\}$ be the candidate item-choice pairs;
+Solve linear program $\Phi(1)$ to obtain optimal solution $x$ with the structure from Lemma 3.3 Let $\mathcal{C}=\{(i,j)\mid x_{ij}>0\}$ be the candidate item-choice pairs
 // Step 2: Virtual Item Construction
 if $\exists$ item $i^{*}$ with two choices $j,j^{\prime}$ such that $x_{i^{*},j},x_{i^{*},j^{\prime}}>0$ then
-   Remove $(i^{*},j)$ and $(i^{*},j^{\prime})$ from $\mathcal{C}$;
-   Create virtual item $(i^{*},j^{*})$ with:
+   Remove $(i^{*},j)$ and $(i^{*},j^{\prime})$ from $\mathcal{C}$ Create virtual item $(i^{*},j^{*})$ with:
  $w_{i^{*},j^{*}}\leftarrow w_{i^{*},j}x_{i^{*},j}+w_{i^{*},j^{\prime}}x_{i^{*},j^{\prime}}$
  $\mu_{i^{*},j^{*}}\leftarrow\mu_{i^{*},j}x_{i^{*},j}+\mu_{i^{*},j^{\prime}}x_{i^{*},j^{\prime}}$
-   Add $(i^{*},j^{*})$ to $\mathcal{C}$;
- // Randomly select between original choices
-   When attempting to insert $(i^{*},j^{*})$, choose $j$ with probability $x_{i^{*},j}$ and $j^{\prime}$ with probability $x_{i^{*},j^{\prime}}$;
-end if
+   Add $(i^{*},j^{*})$ to $\mathcal{C}$ When attempting to insert $(i^{*},j^{*})$, choose $j$ with probability $x_{i^{*},j}$ and $j^{\prime}$ with probability $x_{i^{*},j^{\prime}}$
 // Step 3: Item Insertion with Threshold Check
@@
 if $w_{\max}>\mathcal{T}$ then
-   Insert item $(i_{\max},j_{\max})=\arg\max_{(i,j)\in\mathcal{C}}w_{ij}$;
- return policy that executes only this item;
-end if
+   Insert item $(i_{\max},j_{\max})=\arg\max_{(i,j)\in\mathcal{C}}w_{ij}$ return policy that executes only this item
 else
-   Sort items in $\mathcal{C}$ by decreasing density $\frac{w_{ij}}{\mu_{ij}}$;
- return policy that executes items in decreasing density order;
-end if
-Algorithm 1 SKC (Stochastic Knapsack with Costs)
+ return policy that executes items in $\mathcal{C}$ in decreasing order of density $\frac{w_{ij}}{\mu_{ij}}$
+Algorithm 1 NA-SKC (Non-Adaptive Stochastic Knapsack with Costs)
 An overview of the algorithm.
@@
 Note that from the properties of our solution $x$, it holds that $\sum_{k=1}^{r}{w_{k}}=\Phi(1)$ and $\sum_{k=1}^{r}{\mu_{k}}\leq 1$.
-We then define a threshold $\mathcal{T}(\alpha)$ according to which we decide whether to insert the item with the maximum profit $w_{\text{max}}$.
+We then define a threshold $\mathcal{T}(\alpha)$ according to which we decide whether to insert the item with the maximum utility $w_{\text{max}}$.
 If the threshold is not met, we insert the items in decreasing order of their density $\frac{w_{i}}{\mu_{i}}$, until the knapsack is full or no more items remain in the candidate set $\mathcal{C}$.
@@
 Claim 3.9.
-$\mathbb{E}[\textbf{SKC}]\geq\mathcal{T}=\frac{\Phi(1)}{4(1+\alpha)}$
+$\mathbb{E}[\textbf{NA-SKC}]\geq\mathcal{T}=\frac{\Phi(1)}{4(1+\alpha)}$
 Proof.
-First, if $w_{\max}>\mathcal{T}$, the algorithm attempts only item $(i_{\max},j_{\max})$, gaining an expected net reward of $w_{\max}$.
-Since $w_{\max}>\mathcal{T}$, we have $\mathbb{E}[\textbf{SKC}]=w_{\max}>\mathcal{T}$, as required.
-Now, consider the case $w_{\max}\leq\mathcal{T}$. Define $\textbf{SKC}_{k}$ as the contribution of item $k$ to the algorithm’s profit. It holds that
-$\operatorname{\mathbb{E}}[\textbf{SKC}]=\sum_{k=1}^{r}{\operatorname{\mathbb{E}}[\textbf{SKC}_{k}]}~.$
-We want to lower bound $\operatorname{\mathbb{E}}[\textbf{SKC}_{k}]$.
-If item $k$ is inserted and does not overflow, we obtain a reward of $v_{k}-p_{k}\geq w_{k}$.
+First, if $w_{\max}>\mathcal{T}$, the algorithm attempts only item $(i_{\max},j_{\max})$, gaining an expected net utility of $w_{\max}$.
+Since $w_{\max}>\mathcal{T}$, we have $\mathbb{E}[\textbf{NA-SKC}]=w_{\max}>\mathcal{T}$, as required.
+Now, consider the case $w_{\max}\leq\mathcal{T}$.
+Define $\textbf{NA-SKC}_{k}$ as the contribution of item $k$ to the algorithm’s utility.
+It holds that $\operatorname{\mathbb{E}}[\textbf{NA-SKC}]=\sum_{k=1}^{r}{\operatorname{\mathbb{E}}[\textbf{NA-SKC}_{k}]}.$ We want to lower bound $\operatorname{\mathbb{E}}[\textbf{NA-SKC}_{k}]$.
+If item $k$ is inserted and does not overflow, we obtain a utility of $v_{k}-p_{k}\geq w_{k}$.
 On the other hand, if the item overflows, we pay a cost of $p_{k}$.
-By the definition of the IOR parameter (Definition 2.6), $\alpha\geq\frac{p_{k}}{v_{k}\Pr[s_{k}\leq 1]-p_{k}}=p_{k}/w_{k}$, which implies that an overflowing job incurs a cost of $p_{k}\leq\alpha w_{k}\leq\alpha w_{\max}$.
+By the definition of the IOR parameter (Definition 2.1), $\alpha\geq\frac{p_{k}}{v_{k}\Pr[s_{k}\leq 1]-p_{k}}=p_{k}/w_{k}$, which implies that an overflowing item incurs a cost of $p_{k}\leq\alpha w_{k}\leq\alpha w_{\max}$.
 We get that
-$\operatorname{\mathbb{E}}[\textbf{SKC}_{k}]$ $=(v_{k}-p_{k})\cdot\Pr[k\mbox{ successfully inserted}]-p_{k}\cdot\Pr[k\mbox{ overflows}]$
+$\operatorname{\mathbb{E}}[\textbf{NA-SKC}_{k}]$ $=(v_{k}-p_{k})\cdot\Pr[k\mbox{ successfully inserted}]-p_{k}\cdot\Pr[k\mbox{ overflows}]$
 $\geq(v_{k}-p_{k})\left(1-\sum_{j=1}^{k}{\mu_{j}}\right)-p_{k}\cdot\Pr[k\mbox{ overflows}]$
@@
 where the first inequality follows from lemma 3.6 (with $t=1$), the second inequality follows from the definition of $w_{k}\leq v_{k}-p_{k}$ and the argument above that $p_{k}\leq\alpha\,w_{\max}$.
-In the above, this bound is also true for the virtual item $i^{*}$: the item’s expected truncated size is $\mu_{i^{*}}$, and its expected profit conditioned on the event that it was successfully inserted is at least $w_{i^{*}}$.
+In the above, this bound is also true for the virtual item $i^{*}$: the item’s expected truncated size is $\mu_{i^{*}}$, and its expected utility conditioned on the event that it was successfully inserted is at least $w_{i^{*}}$.
 Summing over all $k\in[r]$, we have that
-$\operatorname{\mathbb{E}}[\textbf{SKC}]$ $\geq\sum_{k=1}^{r}{w_{k}\left(1-\sum_{j=1}^{k}{\mu_{j}}\right)}-\alpha w_{\max}\sum_{k=1}^{r}{\Pr[k\mbox{ overflows}]}$
-$\geq\Phi(1)-\sum_{k=1}^{r}{w_{k}}\sum_{j=1}^{k}{\mu_{j}}-\alpha w_{\max}$ $(*)$
-$=\Phi(1)-\sum_{1\leq j\leq k\leq r}{w_{k}\mu_{j}}-\alpha w_{\max}$
+$\operatorname{\mathbb{E}}[\textbf{NA-SKC}]$ $\geq\sum_{k=1}^{r}{w_{k}\left(1-\sum_{j=1}^{k}{\mu_{j}}\right)}-\alpha w_{\max}\sum_{k=1}^{r}{\Pr[k\mbox{ overflows}]}$
+$\geq\Phi(1)-\sum_{1\leq j\leq k\leq r}{w_{k}\mu_{j}}-\alpha w_{\max}$ $(*)$
 $=\Phi(1)-\sum_{1\leq j<k\leq r}{w_{k}\mu_{j}}-\sum_{k=1}^{r}{w_{k}\mu_{k}}-\alpha w_{\max}$
@@
 $\geq\frac{\Phi(1)}{2}-\sum_{k=1}^{r}{w_{k}\mu_{k}}-\alpha w_{\max}$ $\sum_{k,j=1}^{r}w_{k}\mu_{j}\leq\Phi(1)$
-$\geq\frac{\Phi(1)}{2}-w_{\max}\cdot\left(\sum_{k=1}^{r}{\mu_{k}}+\alpha\right)$
-$\geq\frac{\Phi(1)}{2}-w_{\max}(1+\alpha)$ $\sum_{k=1}^{r}{\mu_{k}}\leq 1$
-$\geq\frac{\Phi(1)}{4}\geq\frac{\Phi(1)}{4(1+\alpha)}=\mathcal{T}~.$ $w_{\max}\leq\mathcal{T}=\frac{\Phi(1)}{4(1+\alpha)}.$ (4)
+$\geq\frac{\Phi(1)}{2}-w_{\max}(1+\alpha)\geq\frac{\Phi(1)}{4}\geq\mathcal{T}~,$ (3)
 In ()(*) we use the fact that $\sum_{k\in[r]}\Pr[k\mbox{ overflows}]\leq 1$ as at most one item overflows, and $\sum_{k=1}^{r}{w_{k}}=\Phi(1)$.
 In ()(**) we use the decreasing order of density: for every $j<k$ it holds that $w_{k}\mu_{j}\leq w_{j}\mu_{k}$, and thus $w_{k}\mu_{j}\leq\frac{1}{2}\left(w_{k}\mu_{j}+w_{j}\mu_{k}\right)$.
+In the final two inequalities we use the fact that $\sum_{k=1}^{r}\mu_{k}\leq 1$ and $w_{\max}\leq\mathcal{T}=\frac{\Phi(1)}{4(1+\alpha)}$.
 We are ready to prove the main theorem of this section:
 Proof of theorem 3.2.
-From 3.8 and Lemma 3.7, we conclude:
-$\frac{\mathrm{ADAPT}}{\operatorname{\mathbb{E}}[\textbf{SKC}]}\leq\frac{2\Phi(1)}{\frac{\Phi(1)}{4(1+\alpha)}}=8(1+\alpha)=O(\alpha)~.$
+From 3.9 and lemma 3.8, we conclude:
+$\frac{\mathrm{ADAPT}}{\operatorname{\mathbb{E}}[\textbf{NA-SKC}]}\leq\frac{2\Phi(1)}{\frac{\Phi(1)}{4(1+\alpha)}}=8(1+\alpha)=O(\alpha)~.$
 In our proof, we show that our algorithm finds a policy that approximates the LP value up to a factor $O(\alpha)$.
 In [10], a similar LP helped guide the solution as its value was a proxy (up to constant factors) of the optimal policy for the stochastic knapsack problem.
-In our case, this LP provides an upper bound to the optimal policy (see Lemma 3.7), but its value can be larger by a factor $\Omega(\alpha)$ than the profit of the optimal policy (see Appendix A).
+In our case, this LP provides an upper bound to the optimal policy (see Lemma 3.8), but its value can be larger by a factor $\Omega(\alpha)$ than the utility of the optimal policy (see Section D.3).
 This shows that we cannot hope for a tighter approximation to the LP’s value.
-In Section 4, we show that every non-adaptive policy cannot have a better approximation than $O(\alpha)$ to the optimal adaptive policy (Theorem 4.1).
-4 Adaptivity Gaps
-In this section we study the power of different policies.
-In Section 4.1, we start by comparing Non-Adaptive policies (Definition 2.3) to Stopping-time Adaptive policies (Definition 2.4).
-Theorem 4.1 shows that the profit of a principal that uses the latter can be higher by a factor of $\Omega(\alpha)$ (which is tight according to Theorem 3.1).
-In Section 4.2 we study the power of Fully-Adaptive policies (Definition 2.2), and show they are more powerful than policies that fix the order of jobs and only decide when to stop processing new jobs.
-Namely, the gap between Fully-Adaptive policies and Stopping-time Adaptive policies can be as high as $\Omega(\alpha^{1/3})$ (Theorem 4.2).
-We note that in the instances that give the lower bound, all jobs have a single choice, and therefore, in order to incentivize jobs, we need to pay their cost exactly.
-Specifically, when determining the IOR, we can substitute the expected payment in Definition 2.6 with the jobs’ costs.
-4.1 Stopping-time Adaptive vs. Non-Adaptive $\Omega(\alpha)$ Gap
-We show that there exists a $\Omega(\alpha)$ gap between the optimal Non-Adaptive policy and the optimal Stopping-time Adaptive policy.
-Recall that both models require the algorithm to choose an item ordering at the beginning; the stopping-time adaptive policy can then adaptively choose when to halt, while the non-adaptive policy must attempt all items in the chosen order.
+4 A Constant Approximation with Limited Adaptivity
+In this section, we study the power of limited adaptivity, namely skipping-adaptive and stopping-adaptive policies.
+In Section 4.1, we prove an $\Omega(\alpha^{1/3})$ gap between stopping-adaptive and skipping-adaptive policies.
+In Section 4.2, we present an $O(1)$-approximation skipping-adaptive algorithm, highlighting the power of adaptively skipping items in a given ordering.
+4.1 Lower Bound for Stopping-Adaptive Policies
+Recall that the stopping-adaptive model requires the algorithm to fix a non-adaptive ordering of the items at the outset, and it halts if all future sequences of items yield a negative expected utility.
+This is arguably the simplest mildly-adaptive class of policies.
+The following theorem shows that allowing the algorithm to adaptively halt is not sufficient to achieve a constant approximation to the optimal fully-adaptive policy in our model.
+In fact, the gap already exists between stopping-adaptive and skipping-adaptive policies.
 Theorem 4.1.
-For every $\alpha$ there exists an instance $I$ of the knapsack-contracts problem such that
-$\frac{\mathrm{STOP\_ADAPT}(I)}{\mathrm{NON\_ADAPT}(I)}=\Omega(\alpha).$
-Proof.
-For a constant $\alpha$, the claim is trivially true.
-We prove the claim for every $\alpha>3$.
-We build an instance $I$ as follows.
-Consider some $\epsilon<\frac{1}{2}$ and any $\gamma\in(0,\tfrac{1}{2})$, where $\epsilon$ will be used to determine $\alpha$.
-Our instance has $n$ jobs for $n=\lceil\frac{1}{\epsilon}\rceil$.
-Each job has value $v=2$ and cost $c=2-\epsilon$.
-The size distribution of job $i\in[n]$ is
-$s_{i}\;=\;\begin{cases}a_{i},&\text{with probability }1-\epsilon,\\[4.0pt] b_{i},&\text{with probability }\epsilon,\end{cases}$
-where
-$a_{i}=\gamma^{\,n-i+1},\qquad b_{i}=1-\displaystyle\sum_{j=1}^{i-1}a_{j}$
-By Definition 2.6, we have
-$\alpha=\frac{2-\epsilon}{\epsilon}\;=\;\Theta\!\left(\frac{1}{\epsilon}\right).$
-Note that for every $\alpha>3$, we can choose an $\epsilon<1/2$ in order to set the IOR of our instance to be $\alpha$.
-Since for every $i$, $\sum_{j=1}^{i-1}a_{j}+b_{i}=1,$ a stopping-time adaptive policy can safely schedule the jobs sequentially until it encounters the first job $i$ whose realized size equals $b_{i}$, and then halt.
-We show that since $b_{i}+a_{i^{\prime}}>1$ for every pair $i<i^{\prime}$, attempting to schedule job $i^{\prime}$ without considering the realization of earlier jobs’ size yields negative expected profit.
-Consequently, the optimal $\mathrm{NON\_ADAPT}$ policy executes at most a single job.
-Our proof combines these two facts to yield the desired bound.
-Formally, consider a stopping‑time adaptive policy that executes jobs in increasing index order ($i=1,2,\ldots$) until it encounters the first job whose size realizes as $s_{i}=b_{i}$, and then stops.
-Let $X$ denote the number of jobs the policy executes.
-We have
-$\operatorname{\mathbb{E}}[X]\;=\;\sum_{i=1}^{n}\Pr[X\geq i]\;=\;\sum_{i=1}^{n}(1-\epsilon)^{i-1}\;=\;\frac{1-(1-\epsilon)^{n}}{\epsilon}.$
-Since $n\geq 1/\epsilon$, we get that $\operatorname{\mathbb{E}}[X]\geq\frac{1-e^{-1}}{\epsilon}$. Therefore the expected profit is at least
-$\mathrm{STOP\_ADAPT}(I)\;=\;\operatorname{\mathbb{E}}[X]\,(v-c)\;\geq\;\frac{1-e^{-1}}{\epsilon}\,\cdot\,\epsilon\;=\;1-e^{-1}.$
-Consider now a non‑adaptive policy.
-This policy fixes the number of jobs to execute, $k$, and an ordering $\pi=i_{1},i_{2},\ldots,i_{k}$.
-For each $\ell\in[k]$, let $\mathrm{profit}(\ell)$ denote the profit from the $\ell$‑th job in the sequence.
-Clearly, $\operatorname{\mathbb{E}}[\mathrm{profit}(1)]=\epsilon$ for every non‑empty sequence of jobs.
-We will show that for every sequence with $k>1$ we have $\operatorname{\mathbb{E}}[\mathrm{profit}(k)]<0$, which implies that it is always better to execute at most a single job, since we can iteratively remove jobs and only improve the expected profit of the policy.
-Given a sequence $\pi=i_{1},i_{2},\ldots,i_{k}$ of jobs, consider first the case in which $i_{\ell}>i_{k}$ for some earlier job in the sequence ($\ell<k$).
-If $i_{k}$ is reached and its size realises as $s_{i_{k}}=b_{i_{k}}$, the job terminates after the deadline, because $s_{i_{\ell}}+s_{i_{k}}\geq a_{i_{\ell}}+b_{i_{k}}>1$.
-Hence
-[…]
-$\geq\Pr[i_{k}\text{ is reached}]\Bigl{(}\Pr[s_{i_{k}}=b_{i_{k}}](-c)+\Pr[s_{i_{k}}=a_{i_{k}}](v-c)\Bigr{)}$
-$=\Pr[i_{k}\text{ is reached}]\bigl{(}\epsilon(-c)+(1-\epsilon)(v-c)\bigr{)}$
-$=\Pr[i_{k}\text{ is reached}]\bigl{(}-2\epsilon+\epsilon^{2}+\epsilon-\epsilon^{2}\bigr{)}$
-$=-\,\epsilon\,\Pr[i_{k}\text{ is reached}]\;<\;0.$
-Now consider the case where $i_{\ell}<i_{k}$ for all $\ell<k$.
-Let $j=\arg\max_{\ell<k}i_{\ell}$, the largest‑index job among the first k1k-1 scheduled.
-Note that, $i_{j}<i_{k}$ by our assumption.
-We have that
-$\operatorname{\mathbb{E}}[\mathrm{profit}(k)]$ $=$ $\operatorname{\mathbb{E}}[\mathrm{profit}(k)|i_{k}\mbox{ is reached}]\cdot\Pr[i_{k}\mbox{ is reached}]$
-$=$ $\operatorname{\mathbb{E}}[\mathrm{profit}(k)|s_{i_{\ell}}=a_{i_{\ell}}\quad\forall\ell<k]\cdot\Pr[s_{i_{\ell}}=a_{i_{\ell}}\quad\forall\ell<k]$
-$+\operatorname{\mathbb{E}}[\mathrm{profit}(k)|i_{k}\mbox{ is reached}\ \wedge\ \exists\ell<k:\ s_{i_{\ell}}=b_{i_{\ell}}]\cdot\Pr[i_{k}\mbox{ is reached}\ \wedge\ \exists\ell<k:\ s_{i_{\ell}}=b_{i_{\ell}}]$
-$=$ $(v-c)\cdot(1-\epsilon)^{k-1}+(-c)\cdot\Pr[i_{k}\mbox{ is reached}\ \wedge\ \exists\ell<k:\ s_{i_{\ell}}=b_{i_{\ell}}]$
-$=$ $\epsilon\cdot(1-\epsilon)^{k-1}+(-c)\cdot\Pr[i_{k}\mbox{ is reached}\ \wedge\ \exists\ell<k:\ s_{i_{\ell}}=b_{i_{\ell}}]$
-$\leq$ $\epsilon\cdot(1-\epsilon)^{k-1}+(-c)\cdot\Pr[s_{i_{j}}=b_{i_{j}}\ \wedge\ (s_{i_{\ell}}=a_{i_{\ell}}\quad\forall\ell<k,\ell\neq j)]$
-$=$ $\epsilon\cdot(1-\epsilon)^{k-1}+(-2+\epsilon)\cdot(1-\epsilon)^{k-2}\cdot\epsilon$
-$=$ $\epsilon\cdot(1-\epsilon)^{k-2}(1-\epsilon-2+\epsilon)$
-$=$ $-\epsilon\cdot(1-\epsilon)^{k-2}<0.$
-In the above, the third equality follows since $i_{k}$ always overflows in case some $s_{i_{\ell}}=b_{i_{\ell}}$ for some $i_{\ell}<i_{k}$, as $b_{i_{\ell}}+a_{i_{k}}>1$.
-The first inequality follows from the fact that c-c is negative, and $\Pr[i_{k}\mbox{ is reached}\ \wedge\ \exists\ell<k:\ s_{i_{\ell}}=b_{i_{\ell}}]\geq\Pr[s_{i_{j}}=b_{i_{j}}\ \wedge\ (s_{i_{\ell}}=a_{i_{\ell}}\quad\forall\ell<k,\ell\neq j)]$ since for the event $s_{i_{j}}=b_{i_{j}}$ and for all other $\ell<k$, $s_{i_{\ell}}=a_{i_{\ell}},$ we have that $i_{k}$ is reached.
-Since
-$\sum_{\ell<k,\ell\neq j}a_{i_{\ell}}\leq\sum_{\ell=1}^{i_{j}-1}a_{\ell}=1-b_{i_{j}},$
-and there will be no overflowing job before $i_{k}$ is executed.
-We conclude that the optimal non-adaptive policy only executes a single job, and obtains a profit of $\epsilon$, while the optimal stopping-time adaptive policy obtains a profit of 1, as argued above.
-Therefore,
-$\frac{\mathrm{STOP\_ADAPT}(I)}{\mathrm{NON\_ADAPT}(I)}\geq\frac{1-e^{-1}}{\epsilon}=\Omega(\alpha).$
-4.2 Fully-adaptive vs. Stopping-time Adaptive $\Omega(\alpha^{1/3})$ Gap
-We further show that there is an $\Omega(\alpha^{1/3})$ gap between the optimal stopping–time adaptive policy and the optimal fully adaptive policy.
-Recall that the stopping-time adaptive model requires the algorithm to fix an ordering of the items at the outset, and it halts if all future sequences of jobs yield a negative expected profit.
-Theorem 4.2.
-For every $\alpha$ there exists an instance $I$ of the knapsack-contracts problem such that
-$\frac{\mathrm{ADAPT}(I)}{\mathrm{STOP\_ADAPT}(I)}=\Omega\bigl{(}{\alpha}^{1/3}\bigr{)}.$
-In order to prove the above theorem, we define an instance $I$ consisting of an infinite collection of jobs, each with with value $v=15/\epsilon$ and cost $c=15/\epsilon-\epsilon^{2}$.
+For every $\alpha$ there exists an instance $I$ of the MSKC problem such that
+$\frac{\mathrm{SKIP\_ADAPT}(I)}{\mathrm{STOP\_ADAPT}(I)}=\Omega\bigl({\alpha}^{1/3}\bigr).$
+In order to prove the above theorem, we define an instance $I$ consisting of an infinite collection of items, each with value $v=15/\epsilon$ and cost $p=15/\epsilon-\epsilon^{2}$.
 As in Section 3.1, $\epsilon$ is used to determine $\alpha$.
 Again, as the claim is trivial for bounded $\alpha$, we prove the claim for large values of $\alpha$.
-The job set is partitioned into $1/\epsilon$ types of jobs $L_{1},\ldots,L_{1/\epsilon}$, each containing an infinite number of jobs. A job of type $L_{i}$ has the size distribution
+The item set is partitioned into $1/\epsilon$ types of items $L_{1},\ldots,L_{1/\epsilon}$, each containing an infinite number of items. An item of type $L_{i}$ has the size distribution
 Probability Size
@@
 $\epsilon^{2}$ $b_{i}$
-where
-$a_{i}=\gamma^{\,n-i+1},\qquad b_{i}=1-\displaystyle\sum_{j=1}^{i-1}a_{j}$
-for $\gamma<\tfrac{1}{2}$.
-By the definition of the IOR parameter (Definition 2.6),
-$\alpha=\frac{15/\epsilon-\epsilon^{2}}{\epsilon^{2}}=15/\epsilon^{3}-1=O\left(\frac{1}{\epsilon^{3}}\right)$
+where $a_{i}=\gamma^{\,n-i+1},\qquad b_{i}=1-\displaystyle\sum_{j=1}^{i-1}a_{j}\qquad$ for $\gamma<\tfrac{1}{2}$.
+By the definition of the IOR parameter (definition 2.1), $\alpha=\frac{15/\epsilon-\epsilon^{2}}{\epsilon^{2}}=15/\epsilon^{3}-1=O\left(\frac{1}{\epsilon^{3}}\right)$.
 For every large enough $\alpha$, we can choose an appropriate $\epsilon$ that sets the IOR to be $\alpha$.
-We first analyze the expected reward from a fully adaptive policy.
+We first analyze the expected utility from a fully-adaptive policy.
 Lemma 4.2.
-For the instance given above,
-$\mathrm{ADAPT}(I)\;\geq\;\frac{1}{2e}.$
+For the instance given above, $\mathrm{SKIP\_ADAPT}(I)\;\geq\;\frac{1}{2e}.$
 Proof.
-Consider an adaptive policy that selects jobs by type $L_{1}$, then type $L_{2}$, and so on until type $L_{1/\epsilon}$, according to the following rule:
-  • When selecting jobs of type $L_{i}$, the policy continues to select such jobs as long as $s_{i}=0$.
-  • Once $s_{i}=a_{i}$, the policy moves on to select jobs of type $L_{i+1}$ if $i<1/\epsilon$.
+Consider a skipping-policy that selects items by type $L_{1}$, then type $L_{2}$, and so on until type $L_{1/\epsilon}$, according to the following rule:
+  • When selecting items of type $L_{i}$, the policy continues to select such items as long as $s_{i}=0$.
+  • Once $s_{i}=a_{i}$, the policy moves on to select items of type $L_{i+1}$ if $i<1/\epsilon$.
   • If $s_{i}=b_{i}$, the policy terminates.
-Notice that the execution order ensures each executed job finishes within the time budget, so the policy obtains a reward of exactly $v-c=\epsilon^{2}$ for every executed job.
-Let $X$ be a random variable for the number of jobs the adaptive policy completes.
-Since the profit from each job is $\epsilon^{2}$, the policy gets a reward of $\operatorname{\mathbb{E}}[X]\cdot\epsilon^{2}$.
-Let $X_{i}$ be the number of jobs of type $L_{i}$ completed by the policy.
+Notice that the execution order ensures each inserted item finishes and does not overflow, so the policy obtains a utility of exactly $v-p=\epsilon^{2}$ for every inserted item.
+Let $X$ be a random variable for the number of items the adaptive policy successfully completes.
+Since the utility from each item is $\epsilon^{2}$, the policy gets a utility of $\operatorname{\mathbb{E}}[X]\cdot\epsilon^{2}$.
+Let $X_{i}$ be the number of items of type $L_{i}$ completed by the policy.
 Notice that
 […]
-By the description of the policy,
-$\operatorname{\mathbb{E}}\bigl{[}X_{i}\mid\varphi\text{ executes jobs of type }L_{i}\bigr{]}=\frac{1}{\epsilon+\epsilon^{2}}.$
+By the description of the policy, $\operatorname{\mathbb{E}}\bigl[X_{i}\mid\varphi\text{ inserts items of type }L_{i}\bigr]=\frac{1}{\epsilon+\epsilon^{2}}.$
 We also have
 […]
-The second inequality follows because in order for the policy to reach jobs of type $L_{1/\epsilon}$, each time a job of type $L_{i}$ (for $i<1/\epsilon$) has its size realized at a value larger than 0, the policy moves on to jobs of type $L_{i+1}$ only if the job’s size is realized at $a_{i}$, which happens with probability $\frac{\epsilon}{\epsilon+\epsilon^{2}}$.
-Since this must occur at most $1/\epsilon$ times, the probability is at least
-$\left(\frac{\epsilon}{\epsilon+\epsilon^{2}}\right)^{1/\epsilon}.$
-Therefore,
-$\operatorname{\mathbb{E}}[X_{i}]\;\geq\;\frac{1}{e(\epsilon+\epsilon^{2})}\;>\;\frac{1}{2e\epsilon},$
-which implies
-$\operatorname{\mathbb{E}}[X]=\sum_{i=1}^{1/\epsilon}\operatorname{\mathbb{E}}[X_{i}]\;>\;\frac{1}{2e\epsilon}\cdot\frac{1}{\epsilon}=\frac{1}{2e\epsilon^{2}}.$
-This yields
-$\mathrm{ADAPT}(I)\;>\;\epsilon^{2}\cdot\frac{1}{2e\epsilon^{2}}\;=\;\frac{1}{2e},$
-as desired. ∎
-We next bound the expected reward from a stopping-time adaptive policy.
+The second inequality follows because in order for the policy to reach items of type $L_{1/\epsilon}$, each time an item of type $L_{i}$ (for $i<1/\epsilon$) has its size realized at a value larger than 0, the policy moves on to items of type $L_{i+1}$ only if the item’s size is realized at $a_{i}$, which happens with probability $\frac{\epsilon}{\epsilon+\epsilon^{2}}$.
+Since this must occur at most $1/\epsilon$ times, the probability is at least $\left(\frac{\epsilon}{\epsilon+\epsilon^{2}}\right)^{1/\epsilon}.$ Therefore, $\operatorname{\mathbb{E}}[X_{i}]\;\geq\;\frac{1}{e(\epsilon+\epsilon^{2})}\;>\;\frac{1}{2e\epsilon},$ which implies $\operatorname{\mathbb{E}}[X]=\sum_{i=1}^{1/\epsilon}\operatorname{\mathbb{E}}[X_{i}]\;>\;\frac{1}{2e\epsilon}\cdot\frac{1}{\epsilon}=\frac{1}{2e\epsilon^{2}}.$ This yields $\mathrm{ADAPT}(I)\;>\;\epsilon^{2}\cdot\frac{1}{2e\epsilon^{2}}\;=\;\frac{1}{2e},$ as desired.
+∎
+We next bound the expected utility from a stopping-adaptive policy.
 Lemma 4.3.
-For the instance given above,
-$\mathrm{STOP\_ADAPT}(I)\;\leq\;4\epsilon.$
+For the instance given above, $\mathrm{STOP\_ADAPT}(I)\;\leq\;4\epsilon.$
 Proof.
-Consider some sequence $\pi$ of jobs.
-Let $Y$ denote the following event: the policy is to process a job of type $L_{i}$, when some prior job of type $L_{i}$ was processed with a size realized to be larger than 0.
-We claim that once event $Y$ occurs, then for every sequence of future jobs, the expected profit is negative.
-Consider first the expected profit of only processing the next job, of type $L_{i}$.
+Consider some sequence $\pi$ of items.
+Let $Y$ denote the following event: the policy is to insert an item of type $L_{i}$, when some prior item of type $L_{i}$ was processed with a size realized to be larger than 0.
+We claim that once event $Y$ occurs, then for every sequence of future items, the expected utility is negative.
+Consider first the expected utility of only inserting the next item, of type $L_{i}$.
 This is at most
-$(1-\epsilon^{2})(v-c)+\epsilon^{2}(-c)=(1-\epsilon^{2})\cdot\epsilon^{2}-\epsilon^{2}(15/\epsilon-\epsilon^{2})=-15\epsilon+\epsilon^{2}<0,$
-since if the next job’s size is realized to be $b_{i}$, then the job finishes after the deadline for sure.
-As the profit of any job is at most $v-c=\epsilon^{2},$ this implies that for any sequence to have a positive profit, it has to be of length larger than $(15\epsilon-\epsilon^{2})/\epsilon^{2}$.
-Consider a partition of the sequence into blocks of size
-$k=\frac{1}{\epsilon+\epsilon^{2}}+\frac{1}{\epsilon},$
-If the sequence ends at a block of size smaller than $k$, we merge the last two blocks together to form a single block, meaning that the size of each block is between $k$ and $2k$.
+$(1-\epsilon^{2})(v-p)+\epsilon^{2}(-p)=(1-\epsilon^{2})\cdot\epsilon^{2}-\epsilon^{2}(15/\epsilon-\epsilon^{2})=-15\epsilon+\epsilon^{2}<0,$
+since if the next item’s size is realized to be $b_{i}$, then the item overflows for sure.
+As the utility of any item is at most $v-p=\epsilon^{2},$ this implies that for any sequence to have a positive utility, it has to be of size larger than $(15\epsilon-\epsilon^{2})/\epsilon^{2}$.
+Consider a partition of the sequence into blocks of size $k=\frac{1}{\epsilon+\epsilon^{2}}+\frac{1}{\epsilon}.$ If the sequence ends at a block of size smaller than $k$, we merge the last two blocks together to form a single block, meaning that the size of each block is between $k$ and $2k$.
 Denote by $r$ the number of blocks in the sequence.
-Consider some block $j\in[r]$, and let $n_{j}(i)$ denote the number of jobs of type $L_{i}$ it contains.
-From our observation, if $n_{j}(i)>1$, then the expected profit of processing the second job of type $L_{i}$ in the block is at most $15\epsilon+\epsilon^{2} \ZoharcommentThis$ is only true given $Y$, right?
-I think the proof might be more clear if it is partitioned (e.g.
-having a couple of clearly stated claims in the proof), in which case the expected profit of the entire block is at most
+Consider some block $j\in[r]$, and let $n_{j}(i)$ denote the number of items of type $L_{i}$ it contains.
+From our observation, if $n_{j}(i)>1$, then the expected utility of processing the second item of type $L_{i}$ in the block is at most $15\epsilon+\epsilon^{2}$, in which case the expected utility of the entire block is at most
 $15\epsilon+\epsilon^{2}+\epsilon^{2}(2k-1)<-15\epsilon+2\epsilon^{2}+4\epsilon=-11\epsilon+2\epsilon^{2}<0.$
-Therefore, to have a positive expected profit, for every $i$, if $b_{j}(i)>1$, then the first $b_{j}(i)-1$ jobs of type $L_{i}$ must have size 0, and in total, at least
-$\sum_{i=1}^{1/\epsilon}{\max\{n_{j}(i)-1,0\}}\geq k-\frac{1}{\epsilon}=\frac{1}{\epsilon+\epsilon^{2}}$
-jobs must have size 0. This occurs with probability at most
-$(1-(\epsilon+\epsilon^{2}))^{\frac{1}{\epsilon+\epsilon^{2}}}<\frac{1}{2}$~
-in which case the expected profit of the block is at most $2k\epsilon^{2}\leq 4\epsilon$.
-Else, with probability at least $1/2$, the expected profit of the block is at most $11\epsilon+2\epsilon^{2}$. Thus, the expected profit of a block bounded given by
-$\frac{1}{2}(-11\epsilon+2\epsilon^{2})+4\epsilon=-\frac{3}{2}\epsilon+\epsilon^{2}<0.$
-Thus, we get that once event $Y$ occurs, since every sequence of jobs has a negative profit for the policy, it will halt.
-Let $X$ be a random variable of the number of jobs processed before event $Y$ occurs.
-Since every job has profit at most $\epsilon^{2}$, we have that $\mathrm{STOP\_ADAPT}(I)\leq\operatorname{\mathbb{E}}[X]\epsilon^{2}$.
+Therefore, to have a positive expected utility, for every $i$, if $n_{j}(i)>1$, then the first $n_{j}(i)-1$ items of type $L_{i}$ must have size 0, and in total, at least $\sum_{i=1}^{1/\epsilon}{\max\{n_{j}(i)-1,0\}}\geq k-\frac{1}{\epsilon}=\frac{1}{\epsilon+\epsilon^{2}}$ items must have size 0.
+This occurs with probability at most $(1-(\epsilon+\epsilon^{2}))^{\frac{1}{\epsilon+\epsilon^{2}}}<\frac{1}{2},$ in which case the expected utility of the block is at most $2k\epsilon^{2}\leq 4\epsilon$.
+Else, with probability at least $1/2$, the expected utility of the block is at most $11\epsilon+2\epsilon^{2}$.
+Thus, the expected utility of a block is bounded by $\frac{1}{2}(-11\epsilon+2\epsilon^{2})+4\epsilon=-\frac{3}{2}\epsilon+\epsilon^{2}<0.$
+Thus, we get that once event $Y$ occurs, since every sequence of items has a negative utility for the policy, it will halt.
+Let $X$ be a random variable of the number of items inserted before event $Y$ occurs.
+Since every item has utility at most $\epsilon^{2}$, we have that $\mathrm{STOP\_ADAPT}(I)\leq\operatorname{\mathbb{E}}[X]\epsilon^{2}$.
 We again break $\pi$ into blocks of size $k$.
-As analyzed above, for every block, we have that event $Y$ occurs with probability at least $1/2$; therefore, the expected number of blocks processed is bounded by 2, and the expected number of jobs processed is bounded by $2k$.
+As analyzed above, for every block, we have that event $Y$ occurs with probability at least $1/2$; therefore, the expected number of blocks processed is bounded by 2, and the expected number of items inserted is bounded by $2k$.
 Therefore,
@@
 as desired. ∎
-Combining Lemmas 4.3 and 4.4 yields Theorem 4.2.
-5 $\Omega(\alpha)$ Informational Gap
+Combining Lemmas 4.2 and 4.3 yields Theorem 4.1.
+4.2 An $O(1)$ Skipping-Adaptive Policy
+theorem 4.1 implies that without skipping items we may not achieve an $\alpha$-independent approximation ratio.
+In this section we provide an algorithm which finds a skipping-adaptive policy with a constant approximation guarantee.
+The LP used in Section 3.2 has a $\Omega(\alpha)$ gap to the optimal solution, as shown in Section D.3.
+In fact, any LP that only uses limited distributional information (e.g.
+the expected truncated sizes) is insufficient to achieve a constant approximation, as shown in Appendix C.
+Thus, we turn to a policy that uses more information than just the moments.
+The idea is to use a global time-indexed LP as in [20, 24, 3], and design an algorithm that aims to get a constant fraction of its value.
+In this section, we assume all item sizes (or times) are in $\left\{0,1,\ldots,T\right\}$ and that the knapsack size is $T$ (equivalently, that all items have rational sizes) [^2] In the original problem we assumed that the knapsack size is $1$.
+Here we rescale all sizes by $T$ since it is more convenient.
+In section E.3 we discuss the handling of continuous distributions..
+In the global time-indexed LP, there is a variable not only for each item-choice $i,j$, but also for each time step $t\in\left\{0,\ldots T\right\}$ (each size is interpreted as a job duration).
+Thus the decision to insert any item depends also on the time of the decision.
+For every $i\in[n],\ j\in[m],\ t\in\left\{0,\ldots T\right\}$ let $w_{i,j,t}:=v_{i}\cdot\Pr_{{s_{i}^{j}\sim F_{i}^{j}}}[s_{i}^{j}\leq T-t]-p_{i}^{j}$ be the effective value of choice $i,j$ given remaining capacity of TtT-t and let $\mu_{i,j,t}=\operatorname{\mathbb{E}}_{s_{i}^{j}\sim F_{i}^{j}}[\min\{s_{i}^{j},t\}]$ be the $t$-truncated expected size of choice $i,j$.
+For any $b\in\mathbb{N}$, let $\Psi(b)$ be the following linear program:
+$\Psi(b)$ $:=\max_{x}\quad\sum_{i\in[n]}\sum_{j\in[m]}\sum_{t\in\left\{0,\ldots b\right\}}w_{i,j,t}\,x_{i,j,t}$
+s.t. $\sum_{i\in[n]}\sum_{j\in[m]}\sum_{t^{\prime}\leq t}x_{i,j,t^{\prime}}\,\mu_{i,j,t}\leq 2t,$ $\forall\,t\in\left\{0,\ldots b\right\},$ (4)
+$\sum_{j\in[m]}\sum_{t\in\left\{0,\ldots b\right\}}x_{i,j,t}\leq 1,$ $\forall\,i\in[n],$ (5)
+$0\leq x_{i,j,t}\leq 1,$ $\forall\,i\in[n],\,j\in[m],\,t\in\left\{0,\ldots b\right\}.$ (6)
+While the runtime of computing the optimal solution is pseudo-polynomial in $b$, it is possible to write a compact (approximate) LP and then use its solution to get a $2$-approximate solution for $\Psi(b)$ in polynomial time.
+All details of the compact (approximate) LP appear in section E.2.
+This $LP$ bounds the expected gain of the optimal adaptive policy:
+Lemma 4.4.
+$\operatorname{\mathbb{E}}[\mathrm{ADAPT}]\leq\Psi(T)$ where $\mathrm{ADAPT}$ is the utility of the optimal adaptive policy.
+We defer the full proof of lemma 4.4 to section E.1.
+The idea is to define $x_{i,j,t}$ as the probability that $\sigma_{A}$ (an optimal adaptive policy) attempts to insert item-choice $(i,j)$ at time $t$, and show that $x$ is a feasible solution to the $LP$.
+This is used to show that the sum of contributions of all item-choice pairs to the expected utility of the optimal adaptive policy is upper bounded by the $LP$.
+We now give the algorithm that finds an adaptive policy that achieves a constant approximation of the LP $\Psi(T)$.
+Input: Instance II with items $i\in[n]$, choices $j\in[m]$, values $v_{i}$, costs $p_{i}^{j}$, distributions $F_{i}^{j}$
+Output: Adaptive policy $\pi$
+// Step 1: Solve LP
+Compute an optimal solution $x^{*}$ to $\Psi(T)$
+// Step 2: Per-item randomized start time (at most one $(j,t)$ per item)
+[…]
+for $i\in[n]$ do
+ $\alpha_{i}\leftarrow\sum_{j\in[m],\,t\in\left\{0,\ldots,T\right\}}x^{*}_{i,j,t}$
+  Sample $A_{i}\sim\textnormal{{Bernoulli}}\!\left(\frac{\alpha_{i}}{4}\right)$
+ if $A_{i}=0$ then continue;
+   Draw $(j_{i},t_{i})\sim\textnormal{{Categorical}}\!\left(\frac{x^{*}_{i,j,t}}{\alpha_{i}}\right)$,   $b_{i,j_{i}}\leftarrow t$ // Implicitly $b_{i,j^{\prime}}\leftarrow\infty$ for $j^{\prime}\neq j_{i}$
+[…]
+// Step 3: Adaptive execution in the random order
+[…]
+foreach $(i,j,b_{i,j})\in\operatorname{\mathcal{I}}$ in order do
+ if $S\leq b_{i,j}$ then
+      Insert $(i,j)$; Observe $s_{i}^{j}$ $S\leftarrow S+s_{i}^{j}$
+Algorithm 2 SA-SKC (Skipping-Adaptive Stochastic Knapsack with Costs)
+The next lemma tells us that given that the randomly assigned start time of item-choice $i,j$ is $b_{i,j}=t$, the probability that all previous item-choices took less than $t$ time is at least half.
+Lemma 4.5.
+[…]
+Proof.
+Let $S=\min\left\{\sum_{(i^{\prime},j^{\prime})\in\operatorname{\mathcal{I}}_{i,j}}s_{i^{\prime}}^{j^{\prime}},t\right\}$ denote how much of $[0,t]$ interval is occupied by items inserted before time $t$.
+Let $x^{*}$ be the optimal solution to $\Psi(T)$.
+Note that if $(i^{\prime},j^{\prime})\in\operatorname{\mathcal{I}}_{i,j}$ it must be that $b_{i^{\prime},j^{\prime}}\leq t$.
+Thus:
+$\operatorname{\mathbb{E}}[S\mid b_{i,j}=t]$ $\leq\sum_{i^{\prime},j^{\prime}\in\operatorname{\mathcal{I}}_{i,j}}\operatorname{\mathbb{E}}\left[\min\left\{s_{i^{\prime}}^{j^{\prime}},t\right\}\ \middle|\ b_{i,j}=t\right]$
+$=\sum_{i^{\prime},j^{\prime}\in\operatorname{\mathcal{I}}_{i,j}}\operatorname{\mathbb{E}}\left[\min\left\{s_{i^{\prime}}^{j^{\prime}},t\right\}\ \middle|\ b_{i,j}=t,b_{i^{\prime},j^{\prime}}\leq t\right]\Pr\left(b_{i^{\prime},j^{\prime}}\leq t\right)$
+$\leq\sum_{i^{\prime},j^{\prime}\in[n]\times[m]}\mu_{i^{\prime},j^{\prime},t}\cdot\left(\frac{\alpha_{i^{\prime}}}{4}\cdot\sum_{t^{\prime}\leq t}\frac{x^{*}_{i^{\prime},j^{\prime},t^{\prime}}}{\alpha_{i^{\prime}}}\right)$
+$=\frac{1}{4}\sum_{i^{\prime},j^{\prime}\in[n]\times[m]}\sum_{t^{\prime}\leq t}x^{*}_{i^{\prime},j^{\prime},t^{\prime}}\cdot\mu_{i^{\prime},j^{\prime},t}\leq\frac{t}{2},$
+where the first inequality follows from the definition of $S$, the second inequality holds since $\operatorname{\mathcal{I}}_{i,j}\subseteq[n]\times[m]$ and the independence between the size of the item $s_{i^{\prime}}^{j^{\prime}}$ and the assigned starting times of the items $b_{i,j},b_{i^{\prime},j^{\prime}}$ and the starting time assignment probability.
+The last inequality holds due to constraint eq.
+4 of the LP (as $x^{*}$ is a solution to $\Psi(T)$).
+By Markov inequality: $\Pr\left(S\geq t\mid b_{i,j}=t\right)\leq\frac{\operatorname{\mathbb{E}}[S\mid b_{i,j}=t]}{t}\leq\frac{1}{2}$, implying the desired.
+We can now show the constant approximation via an adaptive policy result:
+Theorem 4.6.
+Algorithm 2 returns an adaptive policy that achieves a $O(1)$ approximation of the optimal adaptive policy.
+Proof.
+[…]
+$\Pr\left(\mathbb{I}_{\left\{i,j,t\right\}}^{ins}\right)\geq\Pr\left(b_{i,j}=t,s(\operatorname{\mathcal{I}}_{i,j})\leq t\right)=\Pr\left(s(\operatorname{\mathcal{I}}_{i,j})\leq t\mid b_{i,j}=t\right)\cdot\Pr\left(b_{i,j}=t\right)\geq\frac{1}{2}\cdot\frac{x^{*}_{i,j,t}}{4}=\frac{x^{*}_{i,j,t}}{8},$ (7)
+where the last inequality is due to lemma 4.5 and the fact that
+$\Pr\left(b_{i,j}=t\right)=\Pr\left(A_{i}=1\right)\cdot\Pr\left((j_{i},t_{i})=(j,t)\mid A_{i}=1\right)=\frac{\alpha_{i}}{4}\cdot\frac{x^{*}_{i,j,t}}{\alpha_{i}}=\frac{x^{*}_{i,j,t}}{4}.$
+For any $i\in[n],j\in[m]$:
+$\operatorname{\mathbb{E}}[W_{i,j}]$ $=\sum_{t\in\left\{0,\ldots,T\right\}}\operatorname{\mathbb{E}}[W_{i,j,t}]=\sum_{t\in\left\{0,\ldots,T\right\}}\operatorname{\mathbb{E}}[W_{i,j,t}\mid\mathbb{I}_{\left\{i,j,t\right\}}^{ins}]\cdot\Pr\left(\mathbb{I}_{\left\{i,j,t\right\}}^{ins}\right)\geq\sum_{t\in\left\{0,\ldots,T\right\}}w_{i,j,t}\cdot\Pr\left(\mathbb{I}_{\left\{i,j,t\right\}}^{ins}\right)$
+$\geq\frac{1}{8}\sum_{t\in\left\{0,\ldots,T\right\}}x^{*}_{i,j,t}\cdot w_{i,j,t},$ (8)
+where the last inequality holds due to eq. 7.
+From linearity of expectation we get:
+$\operatorname{\mathbb{E}}[\textbf{SA-SKC}]$ $=\sum_{i,j}\operatorname{\mathbb{E}}[W_{i,j}]\geq\sum_{i,j}\frac{1}{8}\sum_{t\in\left\{0,\ldots,T\right\}}x^{*}_{i,j,t}\cdot w_{i,j,t}=\frac{1}{8}\Psi(T)\geq\frac{1}{8}\mathrm{ADAPT},$
+where the first inequality follows from eq. 7 and the second inequality is due to lemma $4.$ ∎
+5 Conclusion and Future Directions
+In this work, we study an economically motivated extension of the canonical Stochastic Knapsack problem through the lens of simple-vs-optimal policies.
+By introducing costs, we show that the adaptivity gap can be unbounded, and that this gap is governed by a natural return-on-investment (ROI) parameter.
+We also introduce a hierarchy of increasingly adaptive policies, and show that limited adaptivity can still achieve near-optimal utility.
+An open question concerns the power of stopping-adaptive policies: while we show an $\Omega(\alpha^{1/3})$ gap relative to skipping-adaptive policies, the best known upper bound remains the $O(\alpha)$ guarantee inherited from non-adaptive policies.
+Closing this gap would complete the picture of our proposed hierarchy.
+A broader direction is to revisit other classic decision-making problems where introducing costs is well motivated, and to study how this changes the optimization landscape.
+We expect simple-vs-optimal approaches to remain useful in such settings, along with the ROI parameter.
+Finally, we suggest re-examining other optimization problems with mixed-sign objectives, and exploring whether our approach can yield new insights there as well.
+Appendix A Further Related Work
+Limited adaptivity.
+Different limited adaptivity notions have been studied in the literature.
+Instead of studying the size of the decision tree, other works study the number of adaptive rounds/choices.
+Several recent works in stochastic optimization explore the power of $k$ adaptive queries compared to full adaptivity: [17, 1] study Stochastic Submodular Cover, while [26] study Informative Path Planning from an information-theoretic perspective (putting aside computational considerations).
+Esfandiari et al.
+[14] consider both Stochastic Submodular Maximization and Stochastic Minimum Cost Coverage.
+The studied problems and techniques are distinct from ours, and focus on the number of adaptive decision rounds/choices; to our knowledge our notions of limited adaptivity have not been previously studied.
+At a high level, our work is related to free-order prophet inequalities and Pandora’s box problems [4, 21], which study adaptive ordering under uncertainty, though the objectives and cost structures differ substantially from ours.
+The power of limited adaptivity has also been studied in the context of non-stochastic optimization, in particular, in maximization of submodular and other complement-free set functions: see e.g.
+the works of [8, 6, 7, 9, 23].
+Algorithmic contract theory
+The study of sequential contracts in [15] investigates the problem of incentivizing a single agent to perform multiple jobs in a sequential manner.
+In contrast, our work considers a sequence of contracts in which the jobs are executed sequentially by different agents, with no interdependence between the jobs.
+The works of [25, 16, 2, 18] consider contract design subject to a limited monetary budget constraining the transfers from principal to agent.
+In our work, we study a multi-agent setting under a time budget constraint on the agents’ actions, where time is limited and the principal must incentivize the agents to complete their jobs on time.
+In recent years, a series of studies has emerged that study contracts through the lens of the theory of computation, initiated by [5, 13]; a recent survey paper [12] provides an overview of developments in algorithmic contract theory.
+The authors of [13] raise concerns regarding the complexity and unintuitiveness of optimal contracts, while exploring the approximation ratios of simpler contract forms, such as linear contracts.
+The multi-agent setting was initially explored by Babaioff et al.
+[5] and further developed by [11], but in their setting the agents only choose between effort and no-effort, and the outcome is only success or failure of the project; we allow multiple actions and outcomes.
+Several works consider agent decisions under costs and incentives, including sequential exploration settings studied in Hoefer et al.
+[22], where optimal contracts are designed for a principal–agent exploration task.
+The work of Hoefer et al.
+[22] is similar in flavor to ours as it combines stochastic optimization with contract design and delegation.
+LP-based approach.
+Our approach builds on LP-based upper bounds for adaptive policies introduced in Dean et al.
+[10], extending them to settings with costs and mixed-sign objectives.
+We use convexity and dominance arguments to obtain well-structured LP solutions.
+In the context of limited-adaptivity, we adapt global time-indexed LP techniques from [3, 20, 24] to design constant-approximation skipping-adaptive policies.
+Unlike [20, 24], our setting also captures multiple choices for each item and negative item utility.
+Appendix B Knapsack Contracts: An Application to Contract Design
+In this section we show the formal connection between MSKC and between a natural multi-agent contract setting, in which a principal sequentially delegates jobs to agents up to a deadline.
+Intuitively, this setting is very similar to MSKC, since the principal has a value for each completed job, there are multiple choices of effort levels at which each job can be executed, and incentivizing different choices incurs different costs on the principal (the cost of paying the agent enough to incentivize the chosen effort level; this cost is well-understood based on existing contract theory).
+There is however one main difference: in the contract setting (unlike MSKC), the principal’s costs are random variables, since the principal pays each agent according to the realized size of each job.
+In section B.1 we set up the contract problem, and in section B.2 we show how to reduce it to MSKC despite the aforementioned difference.
+B.1 Setup
+An instance $I$ of the Knapsack Contracts problem consists of a principal, $n$ agents, and a knapsack of size $S$.
+Without loss of generality, let $S=1$, where $S$ represents a budget of some limited resource.
+As our running example, we assume $S$ is the time remaining until a set deadline.
+Each agent $i$ has a job which they can carry out to the benefit of the principal.
+Agent $i$ is represented by the following:
+  • $v_{i}$: The value which the principal obtains if agent $i$ completes their job before the deadline.
+  • $\{(F_{i}^{j},c_{i}^{j})\}_{j\in[m]}$: The possible effort levels of agent ii.
+[^3] All agents have the same number of effort levels without loss of generality, since we can add “dummy” effort levels that would never be incentivized by the principal.
+Each effort level $j\in[m]$ determines a distribution $F_{i}^{j}$ over the time it takes the agent to complete the job, as well as an agent-cost $c_{i}^{j}\geq 0$.
+If the agent chooses to exert effort level $j$, they incur agent-cost $c_{i}^{j}$, and complete their job within time $s_{i}$ drawn (independently) from distribution $F_{i}^{j}$.
+We assume non-decreasing agent-costs $c_{i}^{1}\leq\dots\leq c_{i}^{m}$.
+Let $\mathbf{F}_{i}=\left(F_{i}^{1},\ldots,F_{i}^{m}\right)$.
+Let $\mathrm{supp}(F_{i}^{j})$ denote the support of $F_{i}^{j}$, and $\mathrm{supp}(\mathbf{F}_{i})=\cup_{j\in[m]}\mathrm{supp}(F_{i}^{j}).$ We assume the support of each distribution is finite and is explicitly given.
+The principal knows $I=(v_{i},\{F_{i}^{j},c_{i}^{j}\}_{j\in[m]})_{i\in[n]}$ in advance.
+After an agent completes a job, the principal observes the time it took the agent to complete it, but it does not observe the effort level chosen by the agent.
+The principal offers each agent $i$ a contract $t_{i}:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$, mapping $i$’s observed processing time to a transfer paid by the principal to the agent.
+Given contract $t_{i}$, agent $i$ chooses an effort level
+$j\in\arg\max_{\ell}\operatorname{\mathbb{E}}_{s_{i}\sim F_{i}^{\ell}}[t_{i}(s_{i})]-c_{i}^{\ell}.$
+Definition B.1 (Implementable effort level).
+For agent $i$, an effort level $j$ is implementable if there exists a contract $t_{i}$ for which this action maximizes the agent’s expected utility among all actions given the contract.
+In this case, we say contract $t_{i}$ incentivizes action $j$.
+Any implementable effort level $j$ for agent $i$ has an optimal contract that incentivizes it, which can be obtained by solving the following LP:
+$\min\quad$ $\sum_{s\in\mathrm{supp}(F_{i}^{j})}t_{i}(s)\cdot F_{i}^{j}(s)$
+s.t.
+$\sum_{s\in\mathrm{supp}(F_{i}^{j})}t_{i}(s)\cdot F_{i}^{j}(s)-c_{i}^{j}\geq\sum_{s\in\mathrm{supp}(F_{i}^{j^{\prime}})}t_{i}(s)\cdot F_{i}^{j^{\prime}}(s)-c_{i}^{j^{\prime}}$ $\forall j^{\prime}\in[m]$
+$t_{i}(s)\geq t_{i}(s^{\prime})$ $\forall s<s^{\prime}\in\mathrm{supp}(\mathbf{F}_{i})$ (9)
+$t_{i}(s)\geq 0$ $\forall s\in\mathrm{supp}(\mathbf{F}_{i})$
+Constraint (9) ensures monotonicity of payments. This is required as without it, an agent might have an incentive to hover around after completing the job and get a higher payment.
+The principal’s problem can then be described as: when hiring agent $i$, the principal chooses an effort level $j$ to incentivize, and offers the optimal contract that incentivizes that effort level.
+The principal’s collected value is the sum of the values of the jobs that the agents get to complete with a total processing time less than the budget $S$.
+The principal aims to maximize their utility: the expected total value minus the expected sum of transfers made to the agents.
+B.2 Reduction
+The following proposition shows that in order to find an approximately optimal policy for the Knapsack Contracts problem, it suffices to find an approximately optimal policy for the MSKC problem.
+See 1.1
+Proof.
+Given a Knapsack Contracts instance, for each agent $i\in[n]$, we create a corresponding item $i$ in the MSKC instance.
+For each implementable effort level $j$ of agent $i$, we create a corresponding choice $j$ for item $i$.
+The parameters are mapped as follows:
+  • The value of item $i$ is set to the principal’s value $v_{i}$.
+  • The size of choice $(i,j)$ is a random variable with the distribution $F_{i}^{j}$.
+  • The cost of choice $(i,j)$ is set to $p_{i}^{j}=\operatorname{\mathbb{E}}_{s\sim F_{i}^{j}}[t_{i}(s)]$, where $t_{i}$ is the contract that minimizes the expected payment, as given by the LP in Definition B.1.
+Now, consider any policy $\varphi$ for the contracts problem, and denote by $\varphi^{\prime}$ the corresponding policy for the knapsack problem.
+$\varphi$ specifies the choice of (agent, effort level) pairs for the contracts problem, and $\varphi^{\prime}$ specifies the corresponding item-choice pairs for the knapsack problem.
+It is left to show that the expected profits are identical.
+Denote by $S$ the set of successfully completed jobs in the contracts and knapsack problems, and by $A$ the set of agents/items attempted in the contracts and knapsack problems.
+Notice that $S$ and $A$ have exactly the same distribution under $\varphi$ and $\varphi^{\prime}$, as both policies make the same choices under the same randomization.
+The utility for policy $\varphi$ is a random variable that depends on the (possibly random) choices of the policy and the random sizes of the attempted jobs. Its expectation is given by
+$\operatorname{\mathbb{E}}[\mathrm{utility}(\varphi)]=\operatorname{\mathbb{E}}\left[\sum_{(i,j)\in S}{v_{i}}-\sum_{(i,j)\in A}{t_{i}(s_{i})}\right]=\operatorname{\mathbb{E}}\left[\sum_{(i,j)\in S}{v_{i}}\right]-\operatorname{\mathbb{E}}\left[\sum_{(i,j)\in A}{t_{i}(s_{i})}\right].$ (10)
+Expanding the transfer term, we get:
+$\operatorname{\mathbb{E}}\left[\sum_{(i,j)\in A}{t_{i}(s_{i})}\right]$ $=\operatorname{\mathbb{E}}\left[\sum_{(i,j)\in A}{t_{i}(s_{i})}\right]=\operatorname{\mathbb{E}}\left[\sum_{(i,j)}{\mathbbm{1}_{(i,j)\in A}\cdot t_{i}(s_{i})}\right]=\sum_{(i,j)}\operatorname{\mathbb{E}}\left[{\mathbbm{1}_{(i,j)\in A}\cdot t_{i}(s_{i})}\right]$
+$\stackrel{{\scriptstyle(\star)}}{{=}}\sum_{(i,j)}\operatorname{\mathbb{E}}\left[\mathbbm{1}_{(i,j)\in A}\right]\cdot\operatorname{\mathbb{E}}_{s_{i}\sim F_{i}^{j}}\left[t_{i}(s_{i})\right]\stackrel{{\scriptstyle(\star\star)}}{{=}}\sum_{(i,j)}\operatorname{\mathbb{E}}\left[\mathbbm{1}_{(i,j)\in A}\right]\cdot p_{i}^{j}$
+$=\operatorname{\mathbb{E}}\left[\sum_{(i,j)}\mathbbm{1}_{(i,j)\in A}\cdot p_{i}^{j}\right]=\operatorname{\mathbb{E}}\left[\sum_{(i,j)\in A}p_{i}^{j}\right],$
+where equality $(\star)$ is due to the independence between $\mathbbm{1}_{(i,j)\in A}$ and $s_{i}^{j}$, and equality $(\star\star)$ is due to the definition of $p_{i}^{j}$.
+Plugging this back in eq.
+10 yields:
+$\operatorname{\mathbb{E}}[\mathrm{utility}(\varphi)]=\operatorname{\mathbb{E}}\left[\sum_{(i,j)\in S}{v_{i}}-\sum_{(i,j)\in A}p_{i}^{j}\right]=\operatorname{\mathbb{E}}[\mathrm{utility}(\varphi^{\prime})]~,$
+which implies the desired.
+Thus, given a $\varphi^{\prime}$ that $c$-approximate the optimal policy in the constructed instance of the MSKC problem, the corresponding $\varphi$ $c$-approximates the optimal policy in the original Knapsack Contracts instance.
+∎
+Implications of the reduction to the Knapsack Contracts Problem.
+The above implies: $(i)$ there exists a skipping-adaptive policy that gives an $O(1)$ approximation to the optimal principal’s utility (by Theorem 4.6); and $(ii)$ one can naturally adapt the ROI parameter to the Knapsack Contracts setting, treating expected transfer that the principal makes when incentivizing agent $i$ to exert effort level $j$ as $p_{i}^{j}$ in Definition 2.1.
+This implies that an instance with $\alpha$ IOR for the Knapsack Contracts problem is reduced to an MSKC instance with $\alpha$ IOR.
+Thus, Theorem 3.2 implies an $O(\alpha)$-approximate non-adaptive policy for the Knapsack Contracts problem.
+As all of our lower bound instances use items with a single choice, the lower bounds also trivially extend to the Knapsack Contracts problem, where each agent has a single effort level, and the principal need only cover the agent-cost exactly.
+Therefore, all results obtained for the MSKC problem, summarized by Figure 1, also apply to the contract problem.
+Appendix C $\Omega(\alpha)$ Informational Gap
 In this section we establish a lower bound that exposes a fundamental information gap in the problem of Stochastic Knapsack with Costs.
 Previous constant-factor approximation algorithms for Stochastic Knapsack (without costs) have used limited distributional information, specifically the expected (or truncated) sizes and the probability of fitting within the current capacity.
-In Proposition 5.1 we demonstrate that this information is insufficient.
-We show that any algorithm, even one that finds an adaptive policy, which relies solely on such limited data, cannot achieve a constant-factor approximation in the presence of costs.
-More generally, we prove in Theorem 5.2 that even providing the first $k$ moments of each distribution for any constant $k$ is insufficient.
-This showcases the inherent difficulty of the added costs, implying that future constant-approximation algorithms will need to utilize almost the full distributional structure in novel ways.
-5.1 Warm-Up: First Moment
-Proposition 5.1.
+In Proposition C.1 we demonstrate that this information is insufficient.
+We show that any algorithm, even one that finds a fully-adaptive policy, which relies solely on such limited data, cannot achieve a constant-factor approximation in the presence of costs.
+More generally, we prove in Theorem C.2 that even providing the first $k$ moments of each distribution for any constant $k$ is insufficient.
+C.1 Warm-Up: First Moment
+Proposition C.1.
 Let $\mathcal{A}$ be any algorithm whose knowledge of the item size distributions is restricted to the expected values $\operatorname{\mathbb{E}}[s_{i}]$, the truncated expected values $\mu_{i}=\operatorname{\mathbb{E}}[\min\{s_{i},1\}]$, and oracle access to the function $q_{i}(B)=\Pr[s_{i}\leq B]$ for each $i\in[n]$, where $B$ is the current remaining knapsack capacity.
@@
 By creating an instance where the number of “bad” items far exceeds the number of “good” items, algorithm $\mathcal{A}$ will almost surely pick a “bad” item first, resulting in our lower bound.
-Proof of Proposition 5.1.
+Proof of Proposition C.1.
 Fix $\epsilon\in(0,1)$. The two item types are defined as follows:
 Property Type 1 Item (“good”) Type 2 Item (“bad”)
-Reward $v_{i}$ 22 22
+Value $v_{i}$ 22 22
 Cost $p_{i}$ $2-\epsilon$ $2-\epsilon$
@@
 Note that in this instance the expected values and the truncated expected values are identical: $\operatorname{\mathbb{E}}[s_{1}]=\operatorname{\mathbb{E}}[s_{2}]=\mu_{1}=\mu_{2}=\epsilon$ and $q_{1}(1)=q_{2}(1)=1$, hence, algorithm $\mathcal{A}$ cannot distinguish between the two items.
-Inserting one item of type 2 results in a net profit of $\epsilon$, and reduces the remaining capacity $B$ by at least $\epsilon/2$.
-Inserting any further items (of either types) results in a negative expected profit:
+Inserting one item of type 2 results in a net utility of $\epsilon$, and reduces the remaining capacity $B$ by at least $\epsilon/2$.
+Inserting any further items (of either types) results in a negative expected utility:
 $\operatorname{\mathbb{E}}[\text{inserting item of type $i$}\;|\;B\leq 1-\epsilon/2]$ $\leq\Pr[s_{i}\neq 1]\cdot v_{i}-p_{i}$
@@
 The optimal policy will attempt to insert items of type 1, and stop if an item’s size instantiates $s_{i}=1$, or if no more items of type 1 remain.
-The expected number of items successfully inserted by OPT is $\frac{1}{\epsilon}$, each item contributing $\epsilon$ net profit, resulting in
+The expected number of items successfully inserted by OPT is $\frac{1}{\epsilon}$, each item contributing $\epsilon$ net utility, resulting in
 $\mathrm{ADAPT}(I)\geq\frac{1}{\epsilon}\cdot\epsilon=1.$
 Since $\mathcal{A}$’s only strategy is to randomly select an item to insert, with probability $1-\frac{1}{n}$ it begins by inserting a “bad” item.
-By taking $n\to\infty$, the expected profit of $\mathcal{A}$ approaches $\epsilon$:
+By taking $n\to\infty$, the expected utility of $\mathcal{A}$ approaches $\epsilon$:
 $\lim_{n\to\infty}\operatorname{\mathbb{E}}[\mathcal{A}(I)]=\epsilon.$
@@
 $\frac{\mathrm{ADAPT}(I)}{\operatorname{\mathbb{E}}[\mathcal{A}(I)]}\geq\frac{1}{2\epsilon}\geq\frac{1}{4}\alpha.$
-5.2 $k$ Moments
+C.2 $k$ Moments
 We generalize the previously shown distributional informational gap to show that, even an algorithm with knowledge of the first $k$ (or truncated) moments of the distribution, has an $\Omega(\alpha)$ gap compared to an algorithm with knowledge of the exact distribution.
-Theorem 5.2.
+Theorem C.2.
 Fix $k\geq 1$.
@@
 Similarly to the previous proof, we construct two types of items, one “good” and one “bad”, which are indistinguishable to $\mathcal{A}$ yet are critically different for the optimal algorithm.
-Proof of Theorem 5.2.
+Proof of Theorem C.2.
 Fix $\epsilon\in(0,1)$. Fix distinct sizes
@@
 Property Type 1 Item (“good”) Type 2 Item (“bad”)
-Reward $v_{i}$ 22 22
+Value $v_{i}$ 22 22
 Cost $c_{i}$ $2-\epsilon$ $2-\epsilon$
@@
 Hence, the two distributions are indistinguishable to $\mathcal{A}$.
-Inserting a single item of type 2 results in a net profit of $\epsilon$, and reduces the remaining capacity $B$ by at least $\delta$.
-Inserting any further items (of either types) results in a negative expected profit:
-$\operatorname{\mathbb{E}}[\text{inserting item of type $i$}\;|\;B\leq 1-\delta]$ $\leq\Pr[s_{i}\neq 1]\cdot v_{i}-c_{i}$
+Inserting a single item of type 2 results in a net utility of $\epsilon$, and reduces the remaining capacity $B$ by at least $\delta$.
+Inserting any further items (of either types) results in a negative expected utility:
+$\operatorname{\mathbb{E}}[\text{inserting item of type $i$}\;|\;B\leq 1-\delta]$ $\leq\Pr[s_{i}\neq 1]\cdot v_{i}-p_{i}$
 $=(1-\epsilon)\cdot 2-(2-\epsilon)$
@@
 The optimal policy does as well as the following policy: it will attempt to insert items of type 1, and stop if an item’s size instantiates $s_{i}\neq 0$, or if no more items of type 1 remain.
-The expected number of items successfully inserted by OPT is $\frac{1}{2\epsilon}$, each item contributing $\epsilon$ net profit, resulting in
+The expected number of items successfully inserted by OPT is $\frac{1}{2\epsilon}$, each item contributing $\epsilon$ net utility, resulting in
 $\mathrm{ADAPT}(I)\geq\frac{1}{2\epsilon}\cdot\epsilon=\frac{1}{2}.$
 Since $\mathcal{A}$’s only strategy is to randomly select an item to insert, with probability $1-\frac{1}{n}$ it begins by inserting a “bad” item.
-By taking $n\to\infty$, the expected profit of $\mathcal{A}$ approaches $\epsilon$:
+By taking $n\to\infty$, the expected utility of $\mathcal{A}$ approaches $\epsilon$:
 $\lim_{n\to\infty}\operatorname{\mathbb{E}}[\mathcal{A}(I)]=\epsilon.$
@@
 $\frac{\mathrm{ADAPT}(I)}{\operatorname{\mathbb{E}}[\mathcal{A}(I)]}\geq\frac{\frac{1}{2}}{2\epsilon}\geq\frac{1}{8}\alpha.$
-6 $O(1/\delta)$-approximation for Bounded-Size Jobs
-In this section, we assume that all jobs are size-bounded with high probability; specifically, we assume that for some $\delta\in(0,1)$, for all $i,j$,
-$\Pr_{s_{i}^{j}\sim F_{i}^{j}}[s_{i}^{j}>1-\delta]\leq\frac{1}{2n\alpha}.$
-By Lemma 2.1, we focus on providing an approximately optimal policy for the MSKC problem.
-As a first step, we analyze a modified problem, where there is no bound on the job sizes, and if a job exceeds the deadline, the principal collects the job’s reward, and then the process terminates.
-For this case, we give a policy that approximates the profit of the optimal policy that does not receive the reward of the job that overflows.
-We then show how to transform this policy into a policy that handles bounded-size jobs and does not collect the reward of overflowing jobs.
-Our algorithm for the case where jobs can overflow, SKC-OF, runs SKC with budget $\delta$ and a different threshold $\mathcal{T}=\Phi(\delta)/4$.
-Note that, unlike the SKC algorithm, this threshold is independent of $\alpha$.
-We denote by $\textbf{SKC-OF}(t)$, the algorithm that runs SKC with a budget $t$, and a threshold $\mathcal{T}=\Phi(t)/4$.
-Specifically, $\textbf{SKC-OF}=\textbf{SKC-OF}(\delta)$.
-We have the following.
-Lemma 6.1.
-For any constant $c\geq 1$,
-$\operatorname{\mathbb{E}}[\textbf{SKC-OF}(t)]\geq\frac{1}{c}\Phi(ct)/4.$
+Appendix D Omitted Proofs from Section 3
+D.1 Proof of theorem 3.1
 Proof.
-In case $w_{\max}\geq\Phi(t)/4$, we only take $w_{\max}$ and obtain a profit of
-$v_{i_{\max}}-p_{i_{\max}}^{j_{\max}}\geq w_{\max}\geq\Phi(t)/4\geq\frac{1}{c}\Phi(ct)/4,$
-where the last inequality follows Lemma 3.7. Therefore, assume $w_{\max}<\Phi(t)/4$.
-Recall that $\textbf{SKC-OF}_{k}$ is the profit of item $k$ in the SKC-OF algorithm, as defined in Section 3.
-In our case, the profit from item $k$ will be as large as in the MSKC instance, and will not incur a cost in case of an overflow.
-Therefore, by adapting Eq.
-(3), our bound on $\textbf{SKC-OF}_{k}$ will be
-$\operatorname{\mathbb{E}}[\textbf{SKC-OF}_{k}]\geq w_{k}\left(1-\frac{1}{t}\sum_{j=1}^{k}{\mu_{j}}\right).$
-We adapt the bound on SKC-OF from Eq. (4) accordingly to be
-$\operatorname{\mathbb{E}}[\textbf{SKC-OF}]\geq\Phi(t)/2-w_{\max}\geq\Phi(t)/4\geq\frac{1}{c}\Phi(ct)/8,$ (7)
-where the second inequality follows our assumption that $w_{\max}\leq\Phi(t)/4$ and the last inequality follows Lemma 3.7. ∎
-We now turn to the bounded job size case.
-Our algorithm for this case, SKC-BOUND, will run SKC-OF, and pretend that the budget is $\delta$, i.e.
-it will terminate if the total size of jobs exceeds $\delta$.
-Theorem 6.2.
-$\frac{\mathrm{ADAPT}}{\operatorname{\mathbb{E}}[\textbf{SKC-BOUND}]}\leq O\left(\frac{1}{\delta}\right)$
-Proof.
-If all items are instantiated with sizes $s_{i}^{j}<1-\delta$, then since the policy terminates when we exceed the budget of $\delta$, the policy never overflows, achieving the expected value of SKC-OF.
-However, if $s_{i}^{j}>1-\delta$ for some $i,j$, the policy terminates and an overflow may occur, resulting in a loss of at most $\alpha w_{\max}$.
+For a constant $\alpha$, the claim is trivially true.
+We prove the claim for every $\alpha>3$.
+We build an instance $I$ as follows.
+Consider some $\epsilon<\frac{1}{2}$ and any $\gamma\in(0,\tfrac{1}{2})$, where $\epsilon$ will be used to determine $\alpha$.
+Our instance has $n$ items for $n=\lceil\frac{1}{\epsilon}\rceil$.
+Each item has value $v=2$ and cost $p=2-\epsilon$.
+The size distribution of item $i\in[n]$ is
+$s_{i}\;=\;\begin{cases}a_{i},&\text{with probability }1-\epsilon,\\[4.0pt] b_{i},&\text{with probability }\epsilon,\end{cases}$ where $\qquad a_{i}=\gamma^{\,n-i+1},\qquad b_{i}=1-\displaystyle\sum_{j=1}^{i-1}a_{j}$
+By Definition 2.1, we have $\alpha=\frac{2-\epsilon}{\epsilon}\;=\;\Theta\!\left(\frac{1}{\epsilon}\right).$ Note that for every $\alpha>3$, we can choose an $\epsilon<1/2$ in order to set the IOR of our instance to be $\alpha$.
+Since for every $i$, $\sum_{j=1}^{i-1}a_{j}+b_{i}=1,$ a stopping-adaptive policy can safely insert the items sequentially until it encounters the first item $i$ whose realized size equals $b_{i}$, and then halt.
+We show that since $b_{i}+a_{i^{\prime}}>1$ for every pair $i<i^{\prime}$, attempting to insert item $i^{\prime}$ without considering the realization of earlier items’ size yields negative expected utility.
+Consequently, the optimal $\mathrm{NON\_ADAPT}$ policy executes at most a single item.
+Our proof combines these two facts to yield the desired bound.
+Formally, consider a stopping-adaptive policy that inserts items in increasing index order ($i=1,2,\ldots$) until it encounters the first item whose size realizes as $s_{i}=b_{i}$, and then stops.
+Let $X$ denote the number of items the policy inserts.
+We have
+$\operatorname{\mathbb{E}}[X]\;=\;\sum_{i=1}^{n}\Pr[X\geq i]\;=\;\sum_{i=1}^{n}(1-\epsilon)^{i-1}\;=\;\frac{1-(1-\epsilon)^{n}}{\epsilon}.$
+Since $n\geq 1/\epsilon$, we get that $\operatorname{\mathbb{E}}[X]\geq\frac{1-e^{-1}}{\epsilon}$. Therefore the expected utility is at least
+$\mathrm{STOP\_ADAPT}(I)\;=\;\operatorname{\mathbb{E}}[X]\,(v-p)\;\geq\;\frac{1-e^{-1}}{\epsilon}\,\cdot\,\epsilon\;=\;1-e^{-1}.$
+Consider now a non‑adaptive policy.
+This policy fixes the number of items to insert, $k$, and an ordering $\pi=i_{1},i_{2},\ldots,i_{k}$.
+For each $\ell\in[k]$, let $\mathrm{utility}(\ell)$ denote the utility from the $\ell$‑th item in the sequence.
+Clearly, $\operatorname{\mathbb{E}}[\mathrm{utility}(1)]=\epsilon$ for every non‑empty sequence of items.
+We will show that for every sequence with $k>1$ we have $\operatorname{\mathbb{E}}[\mathrm{utility}(k)]<0$, which implies that it is always better to execute at most a single item, since we can iteratively remove items and only improve the expected utility of the policy.
+Given a sequence $\pi=i_{1},i_{2},\ldots,i_{k}$ of items, consider first the case in which $i_{\ell}>i_{k}$ for some earlier item in the sequence ($\ell<k$).
+If $i_{k}$ is reached and its size realises as $s_{i_{k}}=b_{i_{k}}$, the item overflows, because $s_{i_{\ell}}+s_{i_{k}}\geq a_{i_{\ell}}+b_{i_{k}}>1$.
+Hence
+[…]
+$\leq\Pr[i_{k}\text{ is reached}]\Bigl(\Pr[s_{i_{k}}=b_{i_{k}}](-p)+\Pr[s_{i_{k}}=a_{i_{k}}](v-p)\Bigr)$
+$=\Pr[i_{k}\text{ is reached}]\bigl(\epsilon(-p)+(1-\epsilon)(v-p)\bigr)$
+$=\Pr[i_{k}\text{ is reached}]\bigl(-2\epsilon+\epsilon^{2}+\epsilon-\epsilon^{2}\bigr)$
+$=-\,\epsilon\,\Pr[i_{k}\text{ is reached}]\;<\;0.$
+Now consider the case where $i_{\ell}<i_{k}$ for all $\ell<k$.
+Let $j=\arg\max_{\ell<k}i_{\ell}$, the largest‑index item among the first k1k-1 inserted.
+Note that, $i_{j}<i_{k}$ by our assumption.
 We have that
-$\operatorname{\mathbb{E}}[\textbf{SKC-BOUND}]$ $\geq\operatorname{\mathbb{E}}[\textbf{SKC-OF}]-\alpha w_{\max}\cdot\Pr[\exists\text{ item $i$ with }s_{i}^{j}>1-\delta]$
-$\geq\Phi(\delta)/4-\alpha w_{\max}\cdot\sum_{i=1}^{n}\Pr[s_{i}^{j}>1-\delta]$
-$\geq\Phi(\delta)/4-\alpha w_{\max}\cdot n\cdot\frac{1}{2n\alpha}$
-$\geq\Phi(\delta)/4-\Phi(\delta)/8$
-$=\Phi(\delta)/8.$
-The theorem follows from Lemma 3.7,
-$\frac{\mathrm{ADAPT}}{\operatorname{\mathbb{E}}[\textbf{SKC-BOUND}]}\leq\frac{\Phi(2)}{\Phi(\delta/8)}\leq\frac{\frac{16}{\delta}\Phi(\delta/8)}{\Phi(\delta/8)}=O\left(\frac{1}{\delta}\right)~.$
-We note that SKC-BOUND determines a non-adaptive order of jobs-effort levels (as it applies SKC), but it applies an adaptive stopping rule, as it stops after the total realized sizes first exceed $\delta$.
-In Appendix B, we show that even for bounded-sized jobs, any non-adaptive policy is bounded away by a factor $\Omega(\sqrt{\alpha})$ from the optimal policy (Theorem B.1).
-In Appendix C, we show that if indeed the principal gets the reward of any job that is initiated before the deadline (even if it finishes after the deadline), then we have a constant adaptivity gap Lemma C.1.
-7 Conclusion and Future Directions
-In this work, we initiate the study of contract design where rewards are fixed, but the principal incentivizes the agents to invest effort in finishing their jobs on time.
-We identify an economically meaningful Inverse of the Return-on-Investment parameter, $\alpha$, which captures many aspects of the problem’s complexity.
-Dean et al.
-[11] show a constant adaptivity gap, while we show that in our case, the adaptivity gap is $\Theta(\alpha)$.
-As the optimal adaptive contract is intractable, a remaining open problem is whether one can efficiently compute an adaptive contract whose profit is a good approximation to the optimal adaptive contract.
-In Section 4.2, we consider a policy with limited adaptivity: the order is non-adaptive, but the policy stops processing jobs once it expects a negative profit.
-We show that for such policies, their profit can be a factor $\Omega(\alpha^{1/3})$ away from the optimal profit.
-As our algorithm in Section 3 guarantees an $O(\alpha)$-approximation, this leaves a gap between the lower and upper bound for such policies.
-It is also interesting to consider other limited-adaptivity policies, e.g., a policy that predetermines the agents’ order, but not their effort levels.
-Another direction is quantifying the profit loss from using a family of restricted contracts, such as contracts that do not pay agents who finish after the deadline.
-Finally, many additional stochastic optimization problems, such as stochastic routing and others, have natural strategic variants were the uncertainty is controlled by an agent’s actions.
-E.g., in routing, service times are likely to result from a strategic choice of an economic entity.
-Marrying the classic algorithms developed for such problems with contractual incentives is an intriguing avenue for future research.
-Appendix A LP Overconfidence: An $\Omega(\alpha)$ Overestimation of the True Profit
-In this section, we provide an instance of the MSKC problem where the value of the LP introduced in Eq.
-(2) is larger by a factor $\Omega(\alpha)$ from the obtainable value of any policy.
-This is in stark contrast to the case without costs, where Dean et al.
-[11] show a non-adaptive policy who’s reward is a constant factor away from the value of the analogous LP.
-As our non-adaptive policy in Section 3 gives $O(\alpha)$-approximation to the value of the LP in Eq.
-(2), this approximation is the best possible with respect to the value of the LP.
-In our instance, we have an infinite number of identical jobs with value $1$ and cost $1-\epsilon$. The size-distribution of the jobs is
-$s=\begin{cases}\epsilon^{n},&\text{w.p. }1-\epsilon,\\ 1,&\text{w.p. }\epsilon;\end{cases}.$
-By Definition 2.6, the value of the IOR parameter is $\alpha=\frac{1-\epsilon}{\epsilon}=O(1/\epsilon)$.
-For every job $i$, we have $\mu_{i}=(1-\epsilon)\cdot\epsilon^{n}+\epsilon\cdot 1=\epsilon^{n}+\epsilon-\epsilon^{n+1}\leq 2\epsilon$ and $w_{i}=\epsilon$.
-Setting $x_{i}=$ for $1/2\epsilon$ is a feasible assignment to the LP, which implies that $\Phi(1)\geq\frac{1}{2\epsilon}\cdot\epsilon=1/2.$ We next bound the reward of any policy for this instance.
-For any policy, after executing the first job, we have that the any job fits with probability at most $1-\epsilon$; thus, we have that the benefit from trying to execute any job after the first one is at most $(1-\epsilon)\epsilon-\epsilon\cdot(1-\epsilon)=0.$ Therefore, the policy that only executes one job is optimal, and it obtains a reward of $\epsilon\leq\frac{\epsilon}{2}\Phi(1)=\Phi(1)/\Omega(\alpha).$
-Appendix B The Need for Adaptivity for Bounded-Size Jobs
-Theorem B.1.
-There exists an instance $I$ of the Knapsack Contracts problem such that
-$\frac{\mathrm{STOP\_ADAPT}(I)}{\mathrm{NON\_ADAPT}(I)}=\Omega\bigl{(}\sqrt{\alpha}\bigr{)}.$
-Proof.
-Let item:
-Property Item
-Reward $v$ 22
-Cost $c$ $2-\epsilon^{2}$
-Size $s$ dist. $s_{1}=\begin{cases}\epsilon^{n},&\text{w.p. }1-\epsilon,\\ \frac{1}{2},&\text{w.p. }\epsilon;\end{cases}$
-We consider the case where $S^{\prime}<\frac{1}{2}$, and compute the expected profit from each job.
-$\operatorname{\mathbb{E}}[w|S^{\prime}<\frac{1}{2}]=v\cdot\Pr[s\leq S^{\prime}]-c=2\cdot(1-\epsilon)-(2-\epsilon^{2})=2-2\epsilon-2+\epsilon^{2}=-2\epsilon+\epsilon^{2}<0$
-In expectation, we can include approximately $1/\epsilon$ jobs before reaching a state where $S^{\prime}<\frac{1}{2}$ so, the expected number of jobs executed by the optimal stopping-time adaptive policy is $1/\epsilon$, and therefore, it obtains an expected profit of $\frac{1}{\epsilon}\cdot\epsilon^{2}=\epsilon.$
-Consider now a non-adaptive policy.
-This policy decides on the number of jobs to execute, and ends whether it overflows or completes the number of jobs it sets to execute.
-Consider first a policy that only executes one job.
-Its reward is exactly $v-c=\epsilon^{2}$.
-Let $\mathrm{profit}(i)$ be the expected profit from the $i$th job in the sequence.
-Since the first job always finishes before the deadline, it is clear that $\mathrm{profit}(1,2)=c-v=\epsilon^{2}$.
-Consider now $\mathrm{profit}(i)$ for $i>2$.
-We have that
-$\mathrm{profit}(i)$ $=$ $\operatorname{\mathbb{E}}[\mathrm{profit}(i)|s_{j}=\epsilon^{n}\ \forall j<i]\cdot\Pr[s_{j}=\epsilon^{n}\quad\forall j<i]$
-$+\operatorname{\mathbb{E}}[\mathrm{profit}(i)|\exists!j<i:\ s_{j}>\epsilon^{n}]\cdot\Pr[\exists!j<i:\ s_{j}>\epsilon^{n}]$
-$=$ $(v-c)\cdot(1-\epsilon)^{i-1}+((v-c)\cdot(1-\epsilon)-c\cdot\epsilon)\cdot\Pr[\exists!j<i:\ s_{j}>0]$
-$=$ $\epsilon^{2}\cdot(1-\epsilon)^{i-1}+((1-\epsilon)v-c)\cdot\Pr[\exists!j<i:\ s_{j}>0]$
-$=$ $\epsilon^{2}\cdot(1-\epsilon)^{i-1}+(\epsilon^{2}-2\epsilon)\cdot\Pr[\exists!j<i:\ s_{j}>0]$
-$\leq$ $\epsilon^{2}\cdot(1-\epsilon)^{i-1}+(\epsilon^{2}-2\epsilon)\cdot\Pr[s_{i-1}>0\ \wedge\ s_{j}=0\quad\forall j<i-1]$
-$=$ $\epsilon^{2}\cdot(1-\epsilon)^{i-1}+(\epsilon^{2}-2\epsilon)\cdot(1-\epsilon)^{i-2}\cdot\epsilon$
-$=$ $\epsilon\cdot(1-\epsilon)^{i-2}(\epsilon\cdot(1-\epsilon)+\epsilon^{2}-2\epsilon)$
-$\leq$ $\epsilon\cdot(1-\epsilon)^{i-2}(\epsilon^{2}-\epsilon)$
-$<$ $0.$
-In the above, the first inequality follows from the fact that $\epsilon^{2}-2\epsilon$ is negative, and that the probability that exactly one of the first i1i-1 jobs has size greater than $\epsilon^{n}$ is larger than the probability that only job i1i-1 has size greater than $\epsilon^{n}$.
-By Definition 2.6, the value of the IOR parameter is $\alpha=\frac{2-\epsilon^{2}}{\epsilon^{2}}=O(1/\epsilon^{2})$.
-Therefore, we have that a non-adaptive policy only loses from having a sequence larger than 2, and the optimal non-adaptive policy has a profit of $2\epsilon^{2}$.
-We conclude that for this instance, we have $\frac{\mathrm{STOP\_ADAPT}(I)}{\mathrm{NON\_ADAPT}(I)}=\frac{1}{2\epsilon}=O(\sqrt{\alpha}).$
-Appendix C Constant Approximation for MSKC with Overflows
-We first provide an upper-bound for expected profit of the optimal adaptive policy that can overflow as well.
-Recall the definition of $\Phi$ given in Eq (2).
-Let $\mathrm{ADAPT}_{OF}$ be the optimal adaptive policy that can overflow.
-Lemma C.1.
-For every instance of the knapsack contract problem $I$, $\mathrm{ADAPT}_{OF}(1)\leq 2\Phi(1)+w_{max}$.
-Proof.
-let $\mathcal{S}$ be the random set of item choices that the policy inserts, and $w_{ij}=v_{i}\cdot\Pr_{{s_{i}^{j}\sim F_{i}^{j}}}[s_{i}^{j}\leq 1]-p_{i}^{j}$ (like in Section 3), $w_{\max}=\max_{(i,j)\in\mathcal{C}}w_{ij}$ (Algorithm 1).
-$\Phi(t)$ is defined in Eq.
-(2), .
-$\mathrm{ADAPT}_{OF}(1)$ $=$ $\sum_{ij}w_{ij}\cdot\Pr[(i,j)\in\mathcal{S}]$ (8)
-$=$ $\sum_{ij}w_{ij}\cdot\Pr[(i,j)\in\mathcal{S}\wedge s_{i}\leq 1]+\sum_{ij}w_{ij}\cdot\Pr[(i,j)\in\mathcal{S}\wedge s_{i}>1]$
-$\leq$ $\sum_{ij}w_{ij}\cdot\Pr[s_{i}\leq 1]\cdot\Pr[(i,j)\in\mathcal{S}]+w_{max}\sum_{ij}\Pr[(i,j)\in\mathcal{S}\wedge s_{i}>1]$
-$\leq$ $\sum_{ij\in\mathcal{S}}w_{ij}\cdot x_{ij}+w_{max}$
-$\leq$ $\Phi(2)+w_{max}$
-$\leq$ $2\Phi(1)+w_{max}.$
-Here the first inequity follows since the probability $s_{i}>1$ is independent from the fact $(i,j)\in\mathcal{S}$.
-The last inequality follows from the fact that, at each run of the adaptive policy, at most one element has a size larger than 1, and therefore the sum of probabilities of the right summand is at most 1.
-By Lemma 3.6, we know that $\operatorname{\mathbb{E}}[\mu(\mathcal{S})]\leq 2$, therefore, we get $\sum_{ij\in A}w_{ij}\cdot x_{ij}\leq\Phi(2)$.
-∎
-Theorem C.2.
-$\textbf{SKC}_{OF}$ is a 9-approximation algorithm to $\mathrm{ADAPT}_{OF}$.
-Proof.
-By Equations (7),(8):
-$\frac{ADAPT_{OF}}{\textbf{SKC-OF}}$ $\leq$ $\frac{2\Phi(1)+w_{max}}{\Phi(1)/4}.$
-If $w_{max}>\Phi(1)/4$: so $\textbf{SKC-OF}=w_{max}$ and:
-$\frac{ADAPT_{OF}}{\textbf{SKC-OF}}$ $\leq$ $\frac{8w_{max}+w_{max}}{w_{max}}=9.$
-Else:
-$\frac{ADAPT_{OF}}{\textbf{SKC-OF}}$ $\leq$ $\frac{2\Phi(1)+\Phi(1)/4}{\Phi(1)/4}=9.$
-Appendix D Ordered Adaptive Contracts
-In many real-world scenarios, the principal lacks the flexibility to reorder agents at will.
-For example, this can be due to procedural requirements, like a fixed sequence of suppliers in a supply chain, or pre-determined project plan with sequential dependencies.
-In these settings, the principal’s adaptive power is limited to deciding, after observing the outcome of each task, whether or not to continue with the next agent, and if so – which effort level to incentivize.
-We define the Ordered Adaptive Contracts model:
-Definition D.1 (Ordered Adaptive Contracts Model).
-In the Ordered Adaptive Contracts model, the principal is given a fixed ordering $\sigma=(1,2,\ldots,n)$ of the $n$ agents.
-For each agent $i$ in order, the principal offers a contract $t_{i}$, and the agent chooses an effort level $j$ to maximize their utility.
-The agents execute their jobs sequentially in the given order.
-The principal observes the completion time of each job and may adaptively decide whether to proceed to the next agent or stop, but cannot change the order.
-The objective is to maximize the expected net profit (total value from completed jobs before the budget is exhausted minus total payments to agents), given the fixed order $\sigma$.
-Since the order is fixed, we are interested in comparing our policy’s performance to the performance of the optimal policy operating on the same order of agents.
-By using the same reduction to the multi-agent contract problem as shown in Lemma 2.1, we focus on solving the matching stochastic knapsack with costs problem.
-In a similar manner to the algorithm that finds an ordered adaptive policy presented in [11] and to the PTAS presented in [5] for the regular Stochastic Knapsack variant (single-choice, no-cost), we use a resource augmentation approach: the policy is allowed a knapsack of size $(1+\epsilon)$, while the optimal adaptive policy has a knapsack of size $1$.
-Our algorithm runs in time $O(\frac{n^{3}m}{\epsilon^{2}})$, offering a tradeoff between the augmented size of the knapsack and the running time of the algorithm.
-We use a dynamic programming algorithm with discretized sizes.
-We denote by $\mathcal{D}(i,k)$ the optimal expected value that can be obtained from inserting items i,,ni,...,n, with $\delta k$ capacity left in the knapsack, for some small $\delta>0$ to be determined later.
-We are interested in computing $\mathcal{D}(1,\lfloor\frac{1}{\delta}\rfloor)$.
-If all of the real sizes are indeed multiples of $\delta k$, then one can obtain an optimal policy by computing:
-$\mathcal{D}(i,k)=\max\left\{\mathcal{D}(i+1,k),\max_{j\in[m]}\left\{v_{i}\cdot\Pr[s_{i}^{j}\leq\delta k]-p_{i}^{j}+\sum_{t=0}^{k}{\mathcal{D}(i+1,k-t)\cdot\Pr[s_{i}^{j}=t\delta]}\right\}\right\}.$
-With the base cases of $\mathcal{D}(n+1,k)=0$ for all $k$, and $\mathcal{D}(i,k)=0$ for $k<0$.
-However, when the distributions are arbitrary, a slight deviation from the discretization can result in an arbitrarily bad solution.
-To overcome this, we must utilize the fact our policy is allowed a knapsack of size $1+\epsilon$.
-We do so by rounding down the true distribution at each step to the nearest multiple of $\delta$.
-By picking a small enough $\delta$, we ensure that the overall rounding error is not too large.
-Theorem D.1.
-There exists an $O(\frac{n^{3}m}{\epsilon^{2}})$ time algorithm that finds a policy for the ordered adaptive knapsack with costs problem, where the policy has a knapsack of size $(1+\epsilon)$, and the optimal adaptive policy has a knapsack of size $1$.
-The policy achieves an expected profit of at least the profit of the optimal policy with a knapsack of size $1$.
-Proof.
-Let OAC be our algorithm and $\mathrm{ADAPT}$ be the optimal adaptive policy. We set the discretization parameter $\delta=\epsilon/n$.
-First, consider a modified problem where all item sizes $s_{i}^{j}$ are rounded down to the nearest multiple of $\delta$: $\hat{s}_{i}^{j}=\delta\lfloor s_{i}^{j}/\delta\rfloor$.
-Let $\widehat{\mathrm{ADAPT}}$ be the optimal expected profit for this problem with rounded-down sizes.
-Since for any policy, the probability of successfully fitting a set of items can only increase when item sizes are reduced, the optimal value for the rounded-down problem is an upper bound on the optimal value for the original problem.
-Thus, $\widehat{\mathrm{ADAPT}}\geq\mathrm{ADAPT}.$
-Our algorithm solves the dynamic program $\mathcal{D}$ described above for the rounded-down sizes.
-This DP computes the exact optimal policy for the rounded-down problem.
-Denote by $\varphi_{\textbf{OAC}}$ the policy derived from this DP.
-When given a knapsack of capacity $1$, this policy achieves the optimal value for the rounded-down problem.
-Let $\widehat{\mathrm{profit}}(\varphi_{\textbf{OAC}},\text{capacity}=1)$ be a random variable of the profit of the policy with the original capacity and rounded-down job sizes.
-We have that $\operatorname{\mathbb{E}}[\widehat{\mathrm{profit}}(\varphi_{\textbf{OAC}},\text{capacity}=1)]=\widehat{\mathrm{ADAPT}}$.
-Combining these, we have $\operatorname{\mathbb{E}}[\widehat{\mathrm{profit}}(\varphi_{\textbf{OAC}},\text{capacity}=1)]\geq\mathrm{ADAPT}.$
-Now, we analyze the performance of our algorithm, which executes the policy $\varphi_{\textbf{OAC}}$ on the original item sizes $s_{i}^{j}$ with an augmented capacity of $1+\epsilon$.
-Consider any execution path of $\varphi_{\textbf{OAC}}$ where a sequence of items and choices $\mathcal{S}\subset[n]\times[m]$ is attempted and successfully completed.
-According to the policy $\varphi_{\textbf{OAC}}$ (which was computed for capacity 1 and sizes $\hat{s}$), the sum of the rounded-down sizes must be no more than 1:
-$\sum_{(i,j)\in\mathcal{S}}\hat{s}_{i}^{j}\leq 1.$
-The true total size of these items is $\sum_{(i,j)\in\mathcal{S}}s_{i}^{j}$.
-We know that for each item, $s_{i}^{j}=\hat{s}_{i}^{j}+\eta_{i}^{j}$, where $0\leq\eta_{i}^{j}<\delta$ is the rounding error.
-The total size in the original problem is:
-$\sum_{(i,j)\in\mathcal{S}}s_{i}^{j}=\sum_{(i,j)\in\mathcal{S}}\hat{s}_{i}^{j}+\sum_{(i,j)\in\mathcal{S}}\eta_{i}^{j}.$
-Since at most $n$ items can be placed in the knapsack, the total accumulated rounding error is bounded:
-$\sum_{(i,j)\in\mathcal{S}}\eta_{i}^{j}\leq n\cdot\delta=n\cdot(\epsilon/n)=\epsilon.$
-Therefore, the true total size of the items is:
-$\sum_{(i,j)\in\mathcal{S}}s_{i}^{j}\leq\sum_{(i,j)\in\mathcal{S}}\hat{s}_{i}^{j}+\epsilon\leq 1+\epsilon.$
-This shows that any sequence of items that fits within capacity 1 in the rounded-down instance is guaranteed to fit within the augmented capacity of $1+\epsilon$ in the original instance.
-Since the policy $\varphi_{\textbf{OAC}}$ makes decisions based on the sequence of outcomes (which are independent of the rounding), every successful path under capacity 1 with sizes $\hat{s}$ is also a successful path under capacity $1+\epsilon$ with sizes $s$.
-Moreover, it crucially follows that an item can cause an overflow under capacity $1+\epsilon$ with sizes $s$ only if it causes an overflow under capacity 1 with sizes $\hat{s}$.
-Therefore, the expected profit can only be higher:
-$\operatorname{\mathbb{E}}[\mathrm{profit}(\varphi_{\textbf{OAC}},\text{capacity}=1+\epsilon)]\geq\operatorname{\mathbb{E}}[\widehat{\mathrm{profit}}(\varphi_{\textbf{OAC}},\text{capacity}=1)].$
-Chaining the inequalities, we get the final result:
-$\textbf{OAC}=\operatorname{\mathbb{E}}[\mathrm{profit}(\varphi_{\textbf{OAC}},\text{capacity}=1+\epsilon)]\geq\operatorname{\mathbb{E}}[\widehat{\mathrm{profit}}(\varphi_{\textbf{OAC}},\text{capacity}=1)]\geq\mathrm{ADAPT}.$
-The dynamic program $\mathcal{D}$ has $O(n)\cdot O(n/\epsilon)$ states, and computing each state takes up to $O(m\cdot n/\epsilon)$ time.
-This results in a running time of $O(\frac{n^{3}m}{\epsilon^{2}})$.
-Appendix E Omitted Proofs from Section 3
+$\operatorname{\mathbb{E}}[\mathrm{utility}(k)]$ $=$ $\operatorname{\mathbb{E}}[\mathrm{utility}(k)|i_{k}\mbox{ is reached}]\cdot\Pr[i_{k}\mbox{ is reached}]$
+$=$ $\operatorname{\mathbb{E}}[\mathrm{utility}(k)|s_{i_{\ell}}=a_{i_{\ell}}\quad\forall\ell<k]\cdot\Pr[s_{i_{\ell}}=a_{i_{\ell}}\quad\forall\ell<k]$
+$+\operatorname{\mathbb{E}}[\mathrm{utility}(k)|i_{k}\mbox{ is reached}\ \wedge\ \exists\ell<k:\ s_{i_{\ell}}=b_{i_{\ell}}]$
+$\ \ \cdot\Pr[i_{k}\mbox{ is reached}\ \wedge\ \exists\ell<k:\ s_{i_{\ell}}=b_{i_{\ell}}]$
+$=$ $(v-p)\cdot(1-\epsilon)^{k-1}+(-p)\cdot\Pr[i_{k}\mbox{ is reached}\ \wedge\ \exists\ell<k:\ s_{i_{\ell}}=b_{i_{\ell}}]$
+$=$ $\epsilon\cdot(1-\epsilon)^{k-1}+(-p)\cdot\Pr[i_{k}\mbox{ is reached}\ \wedge\ \exists\ell<k:\ s_{i_{\ell}}=b_{i_{\ell}}]$
+$\leq$ $\epsilon\cdot(1-\epsilon)^{k-1}+(-p)\cdot\Pr[s_{i_{j}}=b_{i_{j}}\ \wedge\ (s_{i_{\ell}}=a_{i_{\ell}}\quad\forall\ell<k,\ell\neq j)]$
+$=$ $\epsilon\cdot(1-\epsilon)^{k-1}+(-2+\epsilon)\cdot(1-\epsilon)^{k-2}\cdot\epsilon$
+$=$ $\epsilon\cdot(1-\epsilon)^{k-2}(1-\epsilon-2+\epsilon)$
+$=$ $-\epsilon\cdot(1-\epsilon)^{k-2}<0.$
+In the above, the third equality follows since $i_{k}$ always overflows in case some $s_{i_{\ell}}=b_{i_{\ell}}$ for some $i_{\ell}<i_{k}$, as $b_{i_{\ell}}+a_{i_{k}}>1$.
+The first inequality follows from the fact that p-p is negative, and $\Pr[i_{k}\mbox{ is reached}\ \wedge\ \exists\ell<k:\ s_{i_{\ell}}=b_{i_{\ell}}]\geq\Pr[s_{i_{j}}=b_{i_{j}}\ \wedge\ (s_{i_{\ell}}=a_{i_{\ell}}\quad\forall\ell<k,\ell\neq j)]$ since for the event $s_{i_{j}}=b_{i_{j}}$ and for all other $\ell<k$, $s_{i_{\ell}}=a_{i_{\ell}},$ we have that $i_{k}$ is reached.
+Since $\sum_{\ell<k,\ell\neq j}a_{i_{\ell}}\leq\sum_{\ell=1}^{i_{j}-1}a_{\ell}=1-b_{i_{j}},$ there will be no overflowing item before $i_{k}$ is executed.
+We conclude that the optimal non-adaptive policy only inserts a single item, and obtains a utility of $\epsilon$, while the optimal stopping-adaptive policy obtains a utility of 1, as argued above.
+Therefore, $\frac{\mathrm{STOP\_ADAPT}(I)}{\mathrm{NON\_ADAPT}(I)}\geq\frac{1-e^{-1}}{\epsilon}=\Omega(\alpha).$
+D.2 Proof of lemma 3.4
 Full proof of lemma 3.4.
@@
 $\mu_{ij^{\prime}}(w_{ij^{\prime\prime}}-w_{ij})+\mu_{ij^{\prime\prime}}(w_{ij}-w_{ij^{\prime}})-\mu_{ij}(w_{ij^{\prime\prime}}-w_{ij^{\prime}})\leq 0~.$
-This confirms the numerator is non-positive. Since $w_{ij^{\prime\prime}}>w_{ij^{\prime}}$ we conclude that $\Delta_{\mu}\leq 0$.
+This confirms the numerator is non-positive. Since $w_{ij^{\prime\prime}}>w_{ij^{\prime}}$ we conclude that $\Delta_{\mu}\leq 0$. ∎
 Proof of lemma 3.6.
@@
 $\operatorname{\mathbb{E}}[\mu(\mathcal{S})]$ $=\operatorname{\mathbb{E}}[\sum_{(i,j)\in\mathcal{S}}{\mu_{ij}}]=\operatorname{\mathbb{E}}[\sum_{(i,j)\in\mathcal{S}^{-}}{\mu_{ij}}]+\operatorname{\mathbb{E}}[\mu_{i^{*}j^{*}}]$
-$=\operatorname{\mathbb{E}}[\sum_{(i,j)\in\mathcal{S}^{-}}{s_{i}^{j}}]+\operatorname{\mathbb{E}}[s_{i^{*}}^{j^{*}}]=\operatorname{\mathbb{E}}[size(\mathcal{S}^{-})]+\operatorname{\mathbb{E}}[s_{i^{*}}^{j^{*}}]$
+$=\operatorname{\mathbb{E}}[\sum_{(i,j)\in\mathcal{S}^{-}}{s_{i}^{j}}]+\operatorname{\mathbb{E}}[\min\{t,s_{i^{*}}^{j^{*}}\}]=\operatorname{\mathbb{E}}[size(\mathcal{S}^{-})]+\operatorname{\mathbb{E}}[\min\{t,s_{i^{*}}^{j^{*}}\}]$
 $\leq t+t=2t~.$
 Proof of lemma 3.8.
-For any adaptive policy, denote by $\mathcal{S}$ the random set of pairs $(i,j)$ that the policy inserts.
+For any adaptive policy and capacity $t$, denote by $\mathcal{S}$ the random set of pairs $(i,j)$ that the policy inserts.
 Set $x_{ij}=\Pr[(i,j)\in\mathcal{S}]$.
 From the previous lemma, the expected mass that $\mathrm{ADAPT}$ attempts to insert is $\operatorname{\mathbb{E}}(\mu(\mathcal{S}))=\sum_{ij}\mu_{ij}x_{ij}\leq 2t$, therefore, it is a feasible solution to $\Phi(2t)$ with $\sum_{ij}w_{ij}x_{ij}\leq\Phi(2t)$.
-Since $w_{ij}x_{ij}$ is an upper bound on the expected profit of inserting the item choice $(i,j)$, we have that
+Since $w_{ij}x_{ij}$ is an upper bound on the expected utility of inserting the item choice $(i,j)$, we have that
 $\mathrm{ADAPT}\leq\sum_{ij}{w_{ij}x_{ij}}\leq\Phi(2t)\leq\Phi(ct)~,$
@@
 $\mathrm{ADAPT}\leq\Phi(ct)\leq c\cdot\Phi(t)$
-Appendix F Proof of Lemma 2.1
+D.3 LP Overconfidence: An $\Omega(\alpha)$ Overestimation of the True Utility
+In this section, we provide an instance of the MSKC problem where the value of the LP introduced in Eq.
+(1) is larger by a factor $\Omega(\alpha)$ from the obtainable value of any policy.
+This is in stark contrast to the case without costs, where Dean et al.
+[10] show a non-adaptive policy whose utility is a constant factor away from the value of the analogous LP.
+As our non-adaptive policy in Section 3.2 gives $O(\alpha)$-approximation to the value of the LP in Eq.
+(1), this approximation is the best possible with respect to the value of the LP.
+In our instance, we have an infinite number of identical items with value $1$ and cost $1-\epsilon$. The size-distribution of the items is
+$s=\begin{cases}\epsilon^{n},&\text{w.p. }1-\epsilon,\\ 1,&\text{w.p. }\epsilon;\end{cases}.$
+By Definition 2.1, the value of the IOR parameter is $\alpha=\frac{1-\epsilon}{\epsilon}=O(1/\epsilon)$.
+For every item $i$, we have $\mu_{i}=(1-\epsilon)\cdot\epsilon^{n}+\epsilon\cdot 1=\epsilon^{n}+\epsilon-\epsilon^{n+1}\leq 2\epsilon$ and $w_{i}=\epsilon$.
+Setting $x_{i}=1$ for $\ell=\lfloor\frac{1}{2\epsilon}\rfloor$ items, and for an additional item $i^{\prime}$ set $x_{i^{\prime}}=\frac{1}{2\epsilon}$-$\ell$, implies that
+$\sum_{i}x_{i}\cdot\mu_{i}\leq 2\epsilon\cdot\sum_{i}x_{i}=2\epsilon\cdot\frac{1}{2\epsilon}=1,$
+which is a feasible assignment to the LP. Therefore, $\Phi(1)\geq\frac{1}{2\epsilon}\cdot\epsilon=1/2.$ We next bound the utility of any policy for this instance.
+For any policy, after inserting the first item, we have that any item fits with probability at most $1-\epsilon$; thus, we have that the benefit from trying to insert any item after the first one is at most $(1-\epsilon)v-p=0.$ Therefore, the policy that only inserts one item is optimal, and it obtains a utility of $\epsilon\leq 2\epsilon\cdot\Phi(1)=\Phi(1)/\Omega(\alpha).$
+Appendix E Appendix for Section 4
+E.1 Proof of lemma 4.4
 Proof.
-Given a knapsack contracts instance, for each agent $i\in[n]$, we create a corresponding item $i$ in the MSKC instance.
-For each implementable effort level $j$ of agent $i$, we create a corresponding choice $j$ for item $i$.
-The parameters are mapped as follows:
-  • The value of item $i$ is set to the principal’s reward $v_{i}$.
-  • The size of choice $(i,j)$ is a random variable with the distribution $F_{i}^{j}$.
-  • The cost of choice $(i,j)$ is set to $p_{i}^{j}=\operatorname{\mathbb{E}}_{s\sim F_{i}^{j}}[t_{i}(s)]$, where $t_{i}$ is the contract that minimizes the expected payment, as given by the LP in Definition 2.1.
-Now, consider any policy $\varphi$ for the contracts problem, and denote by $\varphi^{\prime}$ the corresponding policy for the knapsack problem.
-$\varphi$ specifies the choice of (agent, effort level) pairs for the contracts problem, and $\varphi^{\prime}$ specifies the corresponding (item, choice) pairs for the knapsack problem.
-It is left to show that the expected profits are identical.
-Denote by $S$ the set of successfully completed jobs in the contracts and knapsack problems, and by $A$ the set of agents/items attempted in the contracts and knapsack problems.
-Notice that $S$ and $A$ have exactly the same distribution under $\varphi$ and $\varphi^{\prime}$, as both policies make the same choices under the same randomization.
-The profit for policy $\varphi$ is a random variable that depends on the (possibly random) choices of the policy and the random sizes of the attempted jobs. Its expectation is given by
-$\operatorname{\mathbb{E}}[\mathrm{profit}(\varphi)]=\operatorname{\mathbb{E}}\left[\sum_{(i,j)\in S}{v_{i}}-\sum_{(i,j)\in A}{t_{i}(s_{i})}\right]=\operatorname{\mathbb{E}}\left[\sum_{(i,j)\in S}{v_{i}}\right]-\operatorname{\mathbb{E}}\left[\sum_{(i,j)\in A}{t_{i}(s_{i})}\right].$ (9)
-Expanding the transfer term, we get:
-$\operatorname{\mathbb{E}}\left[\sum_{(i,j)\in A}{t_{i}(s_{i})}\right]$ $=\operatorname{\mathbb{E}}\left[\sum_{(i,j)\in A}{t_{i}(s_{i})}\right]=\operatorname{\mathbb{E}}\left[\sum_{(i,j)}{\mathbb{1}_{(i,j)\in A}\cdot t_{i}(s_{i})}\right]=\sum_{(i,j)}\operatorname{\mathbb{E}}\left[{\mathbb{1}_{(i,j)\in A}\cdot t_{i}(s_{i})}\right]$
-$\stackrel{{\scriptstyle(\star)}}{{=}}\sum_{(i,j)}\operatorname{\mathbb{E}}\left[\mathbb{1}_{(i,j)\in A}\right]\cdot\operatorname{\mathbb{E}}_{s_{i}\sim F_{i}^{j}}\left[t_{i}(s_{i})\right]\stackrel{{\scriptstyle(\star\star)}}{{=}}\sum_{(i,j)}\operatorname{\mathbb{E}}\left[\mathbb{1}_{(i,j)\in A}\right]\cdot p_{i}^{j}$
-$=\operatorname{\mathbb{E}}\left[\sum_{(i,j)}\mathbb{1}_{(i,j)\in A}\cdot p_{i}^{j}\right]=\operatorname{\mathbb{E}}\left[\sum_{(i,j)\in A}p_{i}^{j}\right],$
-where equality $(\star)$ is due to the independence between $\mathbb{1}_{(i,j)\in A}$ and $s_{i}^{j}$, and equality $(\star\star)$ is due to the definition of $p_{i}^{j}$.
-Plugging this back in Eq.
-9 yields:
-$\operatorname{\mathbb{E}}[\mathrm{profit}(\varphi)]=\operatorname{\mathbb{E}}\left[\sum_{(i,j)\in S}{v_{i}}-\sum_{(i,j)\in A}p_{i}^{j}\right]=\operatorname{\mathbb{E}}[\mathrm{profit}(\varphi^{\prime})]~,$
-which implies the desired.
-Thus, given a $\varphi^{\prime}$ that $c$-approximate the optimal policy in the constructed instance of the MSKC problem, the corresponding $\varphi$ $c$-approximates the optimal policy in the original knapsack contracts instance.
+Let $\sigma_{A}$ be an optimal adaptive policy.
+Let $x_{i,j,t}$ be the probability that $\sigma_{A}$ attempts to insert $(i,j)$ at time $t$.
+We claim that $x$ is a feasible solution to the LP $\Psi(T)$.
+It is immediate that eq.
+6, eq.
+5 hold (by the definition of $x_{i,j,t}$, each item and choice are inserted at most once in the run of $\sigma_{A}$).
+It remains to show that eq.
+4 holds.
+For any $i\in[n]$, $j\in[m]$, $t^{\prime}\in\left\{0,\ldots,T\right\}$: let $\mathbb{I}_{\left\{i,j,t^{\prime}\right\}}^{ins}$ be the indicator for the event where item-choice $(i,j)$ is inserted at time $t^{\prime}$ and let $\mathbb{I}_{\left\{i,j,s\right\}}^{size}$ be the indicator for the event where item-choice $(i,j)$ realizes to size $s$; that is: $s_{i}^{j}=s$.
+For any time $t$, let $(i_{t},j_{t})$ be the last item-choice pair that $\sigma_{A}$ inserts up until or at time $t$, and let $P_{t}$ be the set of all item-choices successfully inserted by $\sigma_{A}$ before inserting $(i_{t},j_{t})$.
+So the sum of all inserted item sizes is at most $t$.
+$\sum_{i,j\in P_{t}}\sum_{t^{\prime}\leq t}\sum_{s\leq T}\mathbb{I}_{\left\{i,j,s\right\}}^{size}\cdot s\leq t.$
+Let $P^{\prime}_{t}=P_{t}\cup\left\{(i_{t},j_{t})\right\}$. We get $\sum_{i,j\in P^{\prime}_{t}}s_{i}^{j}=\sum_{i,j\in P_{t}}s_{i}^{j}+s_{i_{t}}^{j_{t}}$ and thus:
+$\sum_{i,j\in P^{\prime}_{t}}\sum_{t^{\prime}\leq t}\sum_{s\leq T}\mathbb{I}_{\left\{i,j,s\right\}}^{size}\cdot\min(s,t)\leq 2t.$
+For any $t^{\prime}\leq t$ and $i\in[n],j\in[m]$: $\mathbb{I}_{\left\{i,j,t^{\prime}\right\}}^{ins}=1\iff{i,j}\in P^{\prime}_{t}$ and thus we have:
+$\sum_{i,j\in[n]\times[m]}\ \sum_{t^{\prime}\leq t}\ \sum_{s\leq T}\mathbb{I}_{\left\{i,j,t^{\prime}\right\}}^{ins}\cdot\mathbb{I}_{\left\{i,j,s\right\}}^{size}\cdot\min(s,t)\leq 2t.$
+By taking expectation and using linearity of expectation we have:
+$\sum_{i,j\in[n]\times[m]}\ \sum_{t^{\prime}\leq t}\ \sum_{s\leq T}\operatorname{\mathbb{E}}\left[\mathbb{I}_{\left\{i,j,t^{\prime}\right\}}^{ins}\cdot\mathbb{I}_{\left\{i,j,s\right\}}^{size}\right]\cdot\min(s,t)=$
+$\qquad=\sum_{i,j\in[n]\times[m]}\ \sum_{t^{\prime}\leq t}\Pr\left(\mathbb{I}_{\left\{i,j,t^{\prime}\right\}}^{ins}\right)\left(\sum_{s\leq T}\Pr\left(\mathbb{I}_{\left\{i,j,s\right\}}^{size}\right)\min(s,t)\right)$
+$\qquad=\sum_{i,j\in[n]\times[m]}\ \sum_{t^{\prime}\leq t}x_{i,j,t^{\prime}}\ \mu_{i,j,t},$
+where the first equality holds due to the independence between $\mathbb{I}_{\left\{i,j,t^{\prime}\right\}}^{ins}$, $\mathbb{I}_{\left\{i,j,s\right\}}^{size}$ (the events are independent as the policy inserts item-choice $(i,j)$ before seeing the realized value of $s_{i}^{j}$).
+Hence
+$\sum_{i,j\in[n]\times[m]}\ \sum_{t^{\prime}\leq t}x_{i,j,t^{\prime}}\ \mu_{i,j,t}\leq 2t.$
+We’ve shown that $x$ is indeed a feasible solution to the LP.
+We now bound $E[\mathrm{ADAPT}]$.
+Let $W_{i,j}$ be the utility $\sigma_{A}$ gets from item-choice $i,j$ (if $\sigma_{A}$ did not attempt inserting the item-choice then the corresponding utility is 0).
+From linearity of expectation we get:
+$\operatorname{\mathbb{E}}[\mathrm{ADAPT}]$ $=\sum_{i,j}\operatorname{\mathbb{E}}[W_{i,j}]=\sum_{i,j}\sum_{t\in\left\{0,\ldots,T\right\}}\operatorname{\mathbb{E}}[W_{i,j}\mid\sigma_{A}\text{ inserts }(i,j)\text{ at time }t^{\prime}]\cdot\Pr\left(\mathbb{I}_{\left\{i,j,t\right\}}^{ins}\right)$
+$=\sum_{i,j}\sum_{t\in\left\{0,\ldots,T\right\}}\operatorname{\mathbb{E}}[W_{i,j}\mid\sigma_{A}\text{ inserts }(i,j)\text{ at time }t]\cdot x_{i,j,t}$
+$\leq\sum_{i,j}\sum_{t\in\left\{0,\ldots,T\right\}}x_{i,j,t}\cdot w_{i,j,t}\leq\sum_{i,j}\sum_{t\in\left\{0,\ldots,T\right\}}x^{*}_{i,j,t}\cdot w_{i,j,t}=\Psi(T),$
+where the last inequality holds due to the fact that $x$ is a feasible solution to the LP. ∎
+E.2 From Pseudo-Polytime to Truly-Polytime Algorithm
+In this section, we consider the $O(1)$ skipping-adaptive algorithm presented in section 4.2 (see algorithm 2), showing how to turn the algorithm from pseudo-polynomial to fully-polynomial.
+We do so by giving a compact LP that may be solved in polynomial time yet gives a $2$ approximation to the original LP.
+We formulate an LP $\Psi_{P}(b)$ - a compact version of $\Psi(b)$ (defined in section 4.2) with $O(nm\log(b))$ constraints.
+We first show that $\Psi_{P}(b)\geq\frac{1}{2}\Psi(b)$ (lemma E.1), and then give a way to create an approximate solution to $\Psi(b)$ given a solution to $\Psi_{P}(b)$ (lemma E.2).
+Thus, it is enough to use $\Psi_{P}(b)$ in our algorithm.
+We finally discuss how to modify the algorithm to use the compact LP.
+For any $b\in\mathbb{N}$, let $\Psi_{P}(b)$ be the following linear program:
+$\Psi_{P}(b)$ $:=\max_{x}\quad\sum_{i\in[n]}\sum_{j\in[m]}\sum_{t=0}^{\log(b)}w_{i,j,2^{t+1}}\,x_{i,j,2^{t}}$
+s.t. $\sum_{i\in[n]}\sum_{j\in[m]}\sum_{t^{\prime}\leq t}x_{i,j,2^{t^{\prime}}}\,\mu_{i,j,2^{t+1}}\leq 2\cdot 2^{t},$ $\forall\,t\in[0,\log(b)],$ (13)
+$\sum_{j\in[m]}\sum_{t\in[0,\log(b)]}x_{i,j,2^{t}}\leq 1,$ $\forall\,i\in[n],$ (14)
+$0\leq x_{i,j,2^{t}}\leq 1,$ $\forall\,i\in[n],\,j\in[m],\,t\in[0,\log(b)].$ (15)
+Lemma E.1.
+$\Psi_{P}(b)\geq\frac{1}{2}\Psi(b)$.
+Proof.
+Let $x$ be the optimal solution to $\Psi(b)$.
+Let $\hat{x}_{i,j,1}=\frac{1}{2}\sum_{t\in\left\{1,2,3\right\}}x_{i,j,t}$, and $\hat{x}_{i,j,2^{t}}=\frac{1}{2}\sum_{t\in[2^{t+1},2^{t+2})}x_{i,j,t}$ for any $t\in(1,\log(b)]$.
+Let us show that all constraints of $\Psi_{P}(b)$ hold for $\hat{x}$.
+For any $i\in[n]$, by constraint 5 of $\Psi(b)$:
+$\sum_{j\in[m]}\sum_{t\in[0,\log(b)]}\hat{x}_{i,j,2^{t}}\leq\frac{1}{2}\sum_{j\in[m]}\sum_{t\geq 0}x_{i,j,t}\leq\frac{1}{2}.$
+Also,
+$\sum_{i\in[n]}\sum_{j\in[m]}\sum_{t^{\prime}\leq t}\hat{x}_{i,j,2^{t^{\prime}}}\,\mu_{i,j,2^{t+1}}$ $=\sum_{i\in[n]}\sum_{j\in[m]}\sum_{t^{\prime}=1}^{2^{t+2}-1}\frac{1}{2}x_{i,j,t}\cdot\mu_{i,j,2^{t+1}}$
+$\leq\frac{1}{2}\sum_{i\in[n]}\sum_{j\in[m]}\sum_{t^{\prime}=1}^{2^{t+2}-1}x_{i,j,t}\cdot\mu_{i,j,2^{t+2}-1}\leq\frac{1}{2}\left(2^{t+2}-1\right)\leq 2\cdot 2^{t},$
+where the first inequality is due to the fact that $\mu_{i,j,2^{t+1}}\leq\mu_{i,j,2^{t+2}-1}$ by the definition of $\mu$, and the second inequality is due to the fact that $x$ is a feasible solution to $\Psi(b)$ and thus constraint 4 holds.
+So all constraints of $\Psi_{P}(b)$ hold and thus $\hat{x}$ is a feasible solution. Thus, the value obtained from this solution is a lower bound for $\Psi_{P}(b)$ and therefore:
+$\Psi_{P}(b)\geq\sum_{i\in[n]}\sum_{j\in[m]}\sum_{t=0}^{\log(b)}w_{i,j,2^{t+1}}\,\hat{x}_{i,j,2^{t}}\geq\frac{1}{2}\sum_{i\in[n]}\sum_{j\in[m]}\sum_{t\in\left\{0,\ldots,b\right\}}w_{i,j,t}\,x_{i,j,t}=\frac{1}{2}\Psi(b).$
+Let $x^{*}$ be the optimal solution of $\Psi_{P}(b)$.
+Let $y$ be defined as follows: $y_{i,j,t}=\frac{x^{*}_{i,j,2^{l}}}{2^{l}}$ for any $i,j\in[n]\times[m]$ and $l=\lfloor{\log(t)}\rfloor$.
+In the following lemma we show that $y$ is a feasible solution of $\Psi(b)$ which yields the same value as $x^{*}$.
+Lemma E.2.
+Let $x^{*}$ be the optimal solution of $\Psi_{P}(b)$, and let $y\in[n]\times[m]\times\left\{0,\ldots,b\right\}$ be defined by $y_{i,j,t}=\frac{x^{*}_{i,j,2^{l}}}{2^{l}}$ for any $i,j\in[n]\times[m]$ and $l=\lfloor{\log(t)}\rfloor$.
+Then $y$ is a feasible solution of $\Psi(b)$ yielding a value of at least half the optimal LP value.
+Proof.
+Since $x^{*}$ is a feasible solution of $\Psi_{P}(b)$ then constraints 5, 6 hold (due to constraints 14, 15 of $\Psi_{P}(b)$).
+Let $t\in[2^{l},2^{l+1})$ for some $l\in[\log(b)]$.
+$\sum_{i\in[n]}\sum_{j\in[m]}\sum_{t^{\prime}\leq t}y_{i,j,t^{\prime}}\,\mu_{i,j,t}$ $=\sum_{i\in[n]}\sum_{j\in[m]}\sum_{t^{\prime}\leq t}\frac{x^{*}_{i,j,2^{\lfloor\log(t^{\prime})\rfloor}}}{2^{\lfloor\log(t^{\prime})\rfloor}}\,\mu_{i,j,t}\leq\sum_{i\in[n]}\sum_{j\in[m]}\sum_{t^{\prime}\leq t}\frac{x^{*}_{i,j,2^{\lfloor\log(t^{\prime})\rfloor}}}{2^{\lfloor\log(t^{\prime})\rfloor}}\,\mu_{i,j,2^{l+1}}$
+$\leq\sum_{i\in[n]}\sum_{j\in[m]}\sum_{l^{\prime}\leq l}\sum_{t\in[2^{l^{\prime}},2^{l^{\prime}+1}-1]}\frac{x^{*}_{i,j,2^{l^{\prime}}}}{2^{l^{\prime}}}\,\mu_{i,j,2^{l+1}}$
+$=\sum_{i\in[n]}\sum_{j\in[m]}\sum_{l^{\prime}\leq l}x^{*}_{i,j,2^{l^{\prime}}}\,\mu_{i,j,2^{l+1}}\leq 2\cdot 2^{l}\leq 2\cdot t,$
+where the first inequality is due to the monotonicity of $\mu_{i,j,t}$ in $t$, the second inequality is due to summing over more elements and the last inequality is due to constraint 13 of $\Psi_{P}(b)$.
+We deduce that constraint 4 holds, and thus $y$ is a feasible solution to $\Psi(b)$.
+Finally:
+$\sum_{i\in[n]}\sum_{j\in[m]}\sum_{l\in[0,\log(b)]}\sum_{t\in[2^{l},2^{l+1}-1]}w_{i,j,t}\,\frac{x^{*}_{i,j,2^{l}}}{2^{l}}$ $\geq\sum_{i\in[n]}\sum_{j\in[m]}\sum_{l\in[0,\log(b)]}\sum_{t\in[2^{l},2^{l+1}-1]}w_{i,j,2^{l+1}}\,\frac{x^{*}_{i,j,2^{l}}}{2^{l}}$
+$=\sum_{i\in[n]}\sum_{j\in[m]}\sum_{l\in[0,\log(b)]}w_{i,j,2^{l+1}}\,x^{*}_{i,j,2^{l}}=\Psi_{P}(b),$
+where the first inequality holds due to the monotonicity of $w_{i,j,t}$ in $t$. We thus get the desired by lemma E.1. ∎
+The complete fully polynomial algorithm can simply solve the $\Psi_{P}(T)$ LP, and use its solution $x^{*}$ such that for each $x^{*}_{i,j,2^{l}}$ we add an additional step of choosing $t$ in a uniform way: $t\sim Unif{[2^{l},2^{l+1}-1]}$.
+This way, the probability to choose item choice $(i,j)$ with $b_{i,j_{i}}=t$ is $\frac{x^{*}_{i,j,2^{l}}}{2^{l}}$ which is indeed $y$ (and so the proof works as before).
+Since this time $y$ is not the optimal solution, but a $2$-approximate solution, we pay an additional constant factor and still get a $O(1)$ approximation ratio.
+E.3 Handling Continuous Distributions
+In algorithm 2 we assume that (similar to [20]) the size distributions are discrete; namely, that all the items’ sizes are integers between $1$ and $T$ (or, equivalently, that the sizes are $\left\{0,\frac{1}{T},\frac{2}{T},\cdots,1\right\}$).
+In this section, we describe how to apply our algorithm to instances with continuous distributions, using discretization and $\epsilon$-resource augmentation.
+Fix $\epsilon>0$.
+We set the discretization parameter $\delta=\epsilon/n$.
+Given an instance $I$ with continuous size distributions, we construct a rounded instance $\hat{I}$ where all item sizes $s_{i}^{j}$ are rounded down to the nearest multiple of $\delta$: $\hat{s}_{i}^{j}=\delta\lfloor s_{i}^{j}/\delta\rfloor$.
+Let $OPT(I)$ and OPT(I^)OPT($\hat${I}) be the optimal expected utilities for the original and rounded instances, respectively, with capacity 1.
+Since reducing item sizes can only increase the probability of successfully fitting any set of items, the optimal value for the rounded problem is an upper bound on the optimal value for the original problem.
+Thus, $OPT(\hat{I})\geq OPT(I)$.
+algorithm 2 returns a policy $\varphi$ that achieves a $O(1)$ approximate policy for $\hat{I}$.
+We claim that given an augmented capacity of $1+\epsilon$, the expected utility of $\varphi$ on $I$ is the same as the expected utility of $\varphi$ running on $\hat{I}$.
+For any execution path (realization) of $\varphi$, let $\mathcal{S}$ be the set of item-choice pairs successfully inserted by $\varphi$.
+By the knapsack constraint: $\sum_{(i,j)\in\mathcal{S}}\hat{s}_{i}^{j}\leq 1$.
+In the original instance, the true size of each item is $s_{i}^{j}=\hat{s}_{i}^{j}+\eta_{i}^{j}$, where $0\leq\eta_{i}^{j}<\delta$.
+Since at most $n$ items can be placed in the knapsack, the total rounding error is bounded by $n\delta=n(\epsilon/n)=\epsilon$.
+Thus, the total size in the original instance is:
+$\sum_{(i,j)\in\mathcal{S}}s_{i}^{j}<\sum_{(i,j)\in\mathcal{S}}\hat{s}_{i}^{j}+\epsilon\leq 1+\epsilon.$
+We deduce that the same policy $\varphi$ applied to instance $I$ given an $\epsilon$ additional capacity may get at least the same utility as $\varphi$ applied to instance $\hat{I}$ with no additional capacity.
+Appendix F Non-Stopping Policies
+In this section, we consider the non-stopping model, where the policy must first commit to a bundle of items alongside their order of insertion, and then insert all the items according to the order, without stopping at the first overflow; therefore, such a policy incurs the costs of all items in the bundle regardless of how many of these actually fit in the knapsack.
+This is clearly a restriction of the non-adaptive model, where the policy stops after the first overflow.
+In the original work of stochastic knapsack without costs by Dean et al.
+[10], non-adaptive and non-stopping policies are equivalent, since inserting additional items after the first overflow does not increase the expected value and incurs no additional cost.
+However, in our setting with costs, an overflow may result in payments for many items without obtaining their corresponding values.
+This model describes real-world scenarios where payments for jobs are made in advance, before any jobs have been executed and their durations are realized.
+We formally define:
+Definition F.1 (Non-Stopping Policy).
+A non-stopping policy is an ordering of a subset of items and choices $\pi$.
+The policy inserts all items in the given order, paying all their associated costs, and collecting the values of all items that fit in the knapsack before overflow.
+Let $\Pi$ be the set of non-stopping policies.
+We denote by $\mathrm{NON\_STOP}(I)=\max_{\pi\in\Pi}\operatorname{\mathbb{E}}[\mathrm{utility}(\pi)]$ the optimum expected utility obtained by a non-stopping policy for instance $I$.
+In Section F.1 we show that in our model there is a $\Omega(\alpha)$-gap between the optimal non-stopping policy and the optimal non-adaptive policy.
+In Section F.2, we present an $O(\alpha)$-approximate non-stopping policy to the optimal fully-adaptive policy.
+F.1 Non-Stopping vs. Non-Adaptive Policies
+Unlike the no-costs case, in our setting, non-stopping policies can be much worse than non-adaptive policies.
+Theorem F.1.
+For every $\alpha$, there exists an instance $I$ of the MSKC problem such that
+$\frac{\mathrm{NON\_ADAPT}(I)}{\mathrm{NON\_STOP}(I)}=\Omega(\alpha).$
+Proof.
+Consider $n$ items with value $v_{i}=1$, cost $p_{i}=1-\epsilon$, and size distribution
+$s_{i}=\begin{cases}0,&\text{w.p. }1-\epsilon^{2},\\ 2,&\text{w.p. }\epsilon^{2}.\end{cases}$
+Note that $\alpha=\frac{1}{\epsilon}$.
+Denote by $\pi_{NA}$ a non-adaptive policy that inserts all items until the first overflow (i.e., until an item realizes to $s_{i}=2$).
+The expected number of items inserted before the first overflow is $\frac{1}{\epsilon^{2}}$, each contributing a utility of $\epsilon$, except the last (overflowing) item, incurring a cost of $1-\epsilon$.
+This results in a total expected utility of
+$\mathrm{utility}(\pi_{NA})\geq 1/\epsilon-1.$
+For a non-stopping policy, since all items are identical, the only degree of freedom for the algorithm is choosing how many items it would attempt to insert.
+Consider the non-stopping policy $\pi_{NS-k}$ that inserts $k$ items, for some $1\leq k\leq n$.
+We prove that the expected utility of this policy is at most 1.
+The fixed total cost of inserting $k$ items is $k(1-\epsilon)$.
+The expected value from the $i$th item is $(1-\epsilon^{2})^{i}$, as it is successfully inserted only if all the first $i$ items fit.
+Hence, the expected value overall is
+$\sum_{i=1}^{k}(1-\epsilon^{2})^{i}=(1-\epsilon^{2})\frac{1-(1-\epsilon^{2})^{k}}{1-(1-\epsilon^{2})}=\frac{1-\epsilon^{2}}{\epsilon^{2}}[1-(1-\epsilon^{2})^{k}].$
+Therefore, the expected utility of $\pi_{NS-k}$ is
+$\mathrm{utility}(\pi_{NS-k})=\frac{1-\epsilon^{2}}{\epsilon^{2}}[1-(1-\epsilon^{2})^{k}]-k(1-\epsilon).$
+To maximize this expression over $k$, consider the marginal utility of inserting an additional item:
+$\Delta p(k)=\mathrm{utility}(\pi_{NS-k+1})-\mathrm{utility}(\pi_{NS-k})=(1-\epsilon^{2})^{k+1}-(1-\epsilon).$
+$\Delta p(k)$ is decreasing in $k$, thus, the optimal $k$ is the largest $k$ such that $\Delta p(k)>0$. Solving for $k$, we get
+$\Delta p(k)>0$
+$\iff$ $(1-\epsilon^{2})^{k+1}>1-\epsilon$
+$\iff$ $(k+1)\ln(1-\epsilon^{2})>\ln(1-\epsilon)$
+$\iff$ $k<\frac{\ln(1-\epsilon)}{\ln(1-\epsilon^{2})}-1$
+We rewrite the RHS:
+$\frac{\ln(1-\epsilon)}{\ln(1-\epsilon^{2})}-1$ $=\frac{\ln(1-\epsilon)-\ln(1-\epsilon^{2})}{\ln(1-\epsilon^{2})}$
+$=\frac{\ln(1-\epsilon)-\left[\ln(1-\epsilon)+\ln(1+\epsilon)\right]}{\ln(1-\epsilon^{2})}$
+$=\frac{-\ln(1+\epsilon)}{\ln(1-\epsilon^{2})}$
+$\leq\frac{-\epsilon}{-\epsilon^{2}}=\frac{1}{\epsilon}$
+Where the final inequality uses the fact that $\ln(1+x)\leq x$ for x>1x>-1, for both the numerator and denominator.
+Therefore, the optimal $k$ is at most $\frac{1}{\epsilon}$.
+Even if all $k$ items fit, the algorithm’s utility is at most
+$\mathrm{utility}(\pi_{NS-k})\leq k\cdot\epsilon\leq\frac{1}{\epsilon}\cdot\epsilon=1.$
+We conclude that for this instance,
+$\frac{\mathrm{NON\_ADAPT}(I)}{\mathrm{NON\_STOP}(I)}\geq\frac{\mathrm{utility}(\pi_{NA})}{\max_{k}[\mathrm{utility}(\pi_{NS-k})]}\geq\frac{1/\epsilon-1}{1}=\Omega(\alpha).$
+F.2 An $O(\alpha)$ Non-Stopping Policy
+In this section, we present an $O(\alpha)$-approximate non-stopping algorithm for the MSKC problem.
+Theorem F.2.
+There exists a polynomial time algorithm for MSKC that computes for every input $I$ a non-stopping policy $\pi_{I}$ such that $\frac{\mathrm{ADAPT}(I)}{\operatorname{\mathbb{E}}[\mathrm{utility}(\pi_{I})]}\leq O(\alpha).$
+The algorithm is a variation of Algorithm 1: We start by solving the LP $\Phi(1)$ and consider items sorted by their density.
+However, in contrast to Algorithm 1, we then insert only a small ($\approx 1/\alpha$) mass of items, to ensure that with high probability no overflowing occurs.
+An overview of the algorithm.
+First, we solve the LP $\Phi(1)$ to obtain an optimal solution $x$ of item-choices.
+As shown in Section 3.2, we can calculate a solution such that every item has one choice with $x_{ij}=1$ except for at most one item with two choices $x_{ij}+x_{ij^{\prime}}=1$, which can be converted into a single choice.
+Also as in Section 3.2, since each item now has a single (possibly virtual) choice, we refer to each item-choice pair $(i,j)$ by the item $i$ for convenience.
+We then sort the items by decreasing order of density $d_{i}=\frac{w_{i}}{\mu_{i}}$, and define a threshold $U=\frac{1}{3(1+\alpha)}$.
+Denote by $S$ the maximal prefix of items $\{1,\dots,k\}$ such that their cumulative truncated size satisfies
+$\mu(S)=\sum_{i=1}^{k}\mu_{i}\leq U<\sum_{i=1}^{k+1}\mu_{i}.$
+The non-stopping algorithm NS-SKC either inserts the items in $S$, or the single most profitable item-choice $(i_{\max},j_{\max})$ with $w_{\max}=\max_{i}\{w_{i}\}$.
+Specifically, it inserts $S$ if $\frac{2}{3}\sum_{i=1}^{k}w_{i}\geq w_{\max}$, and inserts $(i_{\max},j_{\max})$ otherwise.
+Denote by $\pi_{NS}(S)$ and by $\pi_{NS}(w_{\max})$ the policies that insert $S$ and $(i_{\max},j_{\max})$, respectively.
+The following lemma lower bounds the expected utility obtained when inserting $S$.
+Lemma F.3.
+$\operatorname{\mathbb{E}}[\mathrm{utility}(\pi_{NS}(S))]\geq\frac{2}{3}\sum_{i=1}^{k}w_{i}$
+Proof.
+Denote by $E_{i}$ the event that all items $1,\dots,i$ fit in the knapsack. The expected utility of the policy is
+$\operatorname{\mathbb{E}}[\mathrm{utility}(\pi_{NS}(S))]$ $=\sum_{i=1}^{k}v_{i}\cdot\Pr[E_{i}]-p_{i}$
+$\geq\sum_{i=1}^{k}v_{i}\cdot(1-\mu(S))-p_{i}$ (by lemma 3.6)
+$\geq\sum_{i=1}^{k}(w_{i}+p_{i})\cdot(1-\mu(S))-p_{i}$
+$=(1-\mu(S))\cdot\sum_{i=1}^{k}w_{i}-\mu(S)\cdot\sum_{i=1}^{k}p_{i}$
+$\geq(1-\mu(S))\cdot\sum_{i=1}^{k}w_{i}-\mu(S)\cdot\alpha\sum_{i=1}^{k}w_{i}$ (since $\forall i:p_{i}\leq\alpha w_{i}$)
+$=\sum_{i=1}^{k}w_{i}\cdot(1-(1+\alpha)\mu(S))$
+$\geq\sum_{i=1}^{k}w_{i}\cdot\left(1-\frac{1+\alpha}{3(1+\alpha)}\right)$ (since $\mu(S)\leq U$)
+$=\frac{2}{3}\sum_{i=1}^{k}w_{i}.$
+We can now conclude with the main result of this section.
+Proof of Theorem F.2.
+Since $\sum_{i=1}^{k+1}\mu_{i}>U$ and the items are ordered by density, we have that
+$\sum_{i=1}^{k}w_{i}+w_{\max}\geq\sum_{i=1}^{k+1}w_{i}>U\cdot\Phi(1)=\frac{1}{3(1+\alpha)}\Phi(1).$
+It follows that
+$\max\left\{\sum_{i=1}^{k}w_{i},w_{\max}\right\}\geq\frac{1}{2}\cdot\left(\sum_{i=1}^{k}w_{i}+w_{\max}\right)\geq\frac{1}{6(1+\alpha)}\Phi(1).$
+Our non-stopping policy inserts $S$ if $\frac{2}{3}\sum_{i=1}^{k}w_{i}\geq w_{\max}$, and inserts $(i_{\max},j_{\max})$ otherwise. Therefore, by Lemma F.3 the expected utility of NS-SKC is bounded by
+$\operatorname{\mathbb{E}}[\textbf{NS-SKC}]$ $\geq\max\left\{\operatorname{\mathbb{E}}[\mathrm{utility}(\pi_{NS}(S))],\operatorname{\mathbb{E}}[\mathrm{utility}(\pi_{NS}(w_{\max}))]\right\}$
+$\geq\max\left\{\frac{2}{3}\sum_{i=1}^{k}w_{i},w_{\max}\right\}$
+[…]
+We remark that [10] studies a policy class termed fixed-set, in which the policy must choose an unordered subset of items to insert, and receives the cumulative values of the items if and only if the entire set is successfully inserted into the knapsack; otherwise, it obtains zero utility.
+In the classic stochastic knapsack problem without costs, it is shown that there exists a fixed-set policy that achieves a constant approximation of the fully-adaptive optimal policy.
+Note that a fixed-set policy can only perform worse than a non-stopping policy, since an overflow results in no received value at all (and possibly a large negative utility in our model due to the presence of costs).
+This is in contrast to a non-stopping policy that, at the very least, receives the value of the items inserted prior to the overflow.
+However, our non-stopping policy can be easily converted into a fixed-set policy: either insert $S$ in its entirety, or insert $(i_{\max},j_{\max})$.
+It can be similarly shown that such a policy achieves an $O(\alpha)$ approximation of the optimal fully-adaptive policy.
+Appendix G Relaxations
+In this section, we consider several relaxations of the original model.
+In particular, we focus on the setting where items are size-bounded with high probability; specifically, for some $\delta\in(0,1)$, for all $i,j$,
+$\Pr_{s_{i}^{j}\sim F_{i}^{j}}\!\left[s_{i}^{j}>1-\delta\right]\leq\frac{1}{2n\alpha}.$
+To facilitate the analysis, we first study a relaxed variant in which there is no explicit bound on item sizes.
+In this variant, the policy receives the value of each item inserted before the capacity is met, and the process terminates upon overflow.
+We show that this policy approximates the utility of the optimal policy in the original model, where no value is collected from an overflowing item.
+We then demonstrate how to transform this policy into one that operates under the bounded-size assumption and does not collect value from overflowing items.
+G.1 Start-by-Deadline Model
+We define SKC-OF to be the algorithm that accepts all items inserted before reaching the capacity.
+SKC-OF runs NA-SKC (Algorithm 1) with a different threshold $\mathcal{T}=\Phi(\delta)/4$.
+Note that, unlike the NA-SKC algorithm, this threshold is independent of $\alpha$.
+We denote by $\textbf{SKC-OF}(t)$, the algorithm that runs NA-SKC with a capacity $t$, and a threshold $\mathcal{T}=\Phi(t)/4$.
+In particular, $\textbf{SKC-OF}=\textbf{SKC-OF}(\delta)$.
+Lemma G.1.
+For any constant $c\geq 1$,
+$\operatorname{\mathbb{E}}[\textbf{SKC-OF}(t)]\geq\frac{1}{c}\Phi(ct)/4.$
+Proof.
+In case $w_{\max}\geq\Phi(t)/4$, we only take $w_{\max}$ and obtain a utility of
+$v_{i_{\max}}-p_{i_{\max}}^{j_{\max}}\geq w_{\max}\geq\Phi(t)/4\geq\frac{1}{c}\Phi(ct)/4,$
+where the last inequality follows Lemma 3.8. Hence, assume that $w_{\max}<\Phi(t)/4$.
+Recall that $\textbf{SKC-OF}_{k}$ is the utility of item $k$ in the SKC-OF algorithm, as defined in Section 3.2.
+In our case, the utility from item $k$ will be as large as in the MSKC instance, and will not incur a cost in case of an overflow.
+Therefore, by adapting Eq.
+(2), our bound on $\textbf{SKC-OF}_{k}$ will be
+$\operatorname{\mathbb{E}}[\textbf{SKC-OF}_{k}]\geq w_{k}\left(1-\frac{1}{t}\sum_{j=1}^{k}{\mu_{j}}\right).$
+We adapt the bound on SKC-OF from Eq. (3) accordingly to be
+$\operatorname{\mathbb{E}}[\textbf{SKC-OF}]\geq\Phi(t)/2-w_{\max}\geq\Phi(t)/4\geq\frac{1}{c}\Phi(ct)/8,$ (16)
+where the second inequality follows our assumption that $w_{\max}\leq\Phi(t)/4$ and the last inequality follows Lemma 3.8. ∎
+If the policy gets the value of any item inserted when there is still some capacity left (even if it overflows), then the adaptivity gap becomes constant, as formalized below.
+Lemma G.2.
+For every instance of the MSKC problem $I$, $\mathrm{ADAPT}_{OF}(1)\leq 2\Phi(1)+w_{max}$.
+Proof.
+let $\mathcal{S}$ be the random set of item choices selected by the policy, and recall that $w_{ij}=v_{i}\cdot\Pr_{{s_{i}^{j}\sim F_{i}^{j}}}[s_{i}^{j}\leq 1]-p_{i}^{j}$ (like in section 3.2, and w.l.o.g we assume that: $\Pr_{{s_{i}^{j}\sim F_{i}^{j}}}[s_{i}^{j}\leq 1]=1$), $w_{\max}=\max_{(i,j)\in\mathcal{C}}w_{ij}$ (Algorithm 1).
+Using $\Phi(t)$ (Eq.
+(1)), we obtain:
+$\mathrm{ADAPT}_{OF}(1)$ $=$ $\sum_{ij}w_{ij}\cdot\Pr[(i,j)\in\mathcal{S}]$ (17)
+$=$ $\sum_{ij}w_{ij}\cdot\Pr[(i,j)\in\mathcal{S}\wedge s_{i}\leq 1]+\sum_{ij}w_{ij}\cdot\Pr[(i,j)\in\mathcal{S}\wedge s_{i}>1]$
+$\leq$ $\sum_{ij}w_{ij}\cdot\Pr[s_{i}\leq 1]\cdot\Pr[(i,j)\in\mathcal{S}]+w_{max}\sum_{ij}\Pr[(i,j)\in\mathcal{S}\wedge s_{i}>1]$
+$\leq$ $\sum_{ij\in\mathcal{S}}w_{ij}\cdot x_{ij}+w_{max}$
+$\leq$ $\Phi(2)+w_{max}$
+$\leq$ $2\Phi(1)+w_{max}.$
+The first inequality holds because $\Pr[s_{i}>1]$ is independent of $(i,j)\in\mathcal{S}$.
+The final inequality follows since at most one element per run has size greater than 1, implying the summation of those probabilities is at most one.
+By Lemma 3.7, we know $\operatorname{\mathbb{E}}[\mu(\mathcal{S})]\leq 2$, giving $\sum_{ij\in A}w_{ij}x_{ij}\leq\Phi(2)$.
 ∎
+Theorem G.3.
+SKC-OF is a 9-approximation algorithm to $\mathrm{ADAPT}_{OF}$.
+Proof.
+By Equations (16),(17):
+$\frac{\mathrm{ADAPT}_{OF}}{\textbf{SKC-OF}}$ $\leq$ $\frac{2\Phi(1)+w_{max}}{\Phi(1)/4}.$
+If $w_{max}>\Phi(1)/4$, then $\textbf{SKC-OF}=w_{max}$ giving:
+$\frac{\mathrm{ADAPT}_{OF}}{\textbf{SKC-OF}}$ $\leq$ $\frac{8w_{max}+w_{max}}{w_{max}}=9.$
+Else:
+$\frac{\mathrm{ADAPT}_{OF}}{\textbf{SKC-OF}}$ $\leq$ $\frac{2\Phi(1)+\Phi(1)/4}{\Phi(1)/4}=9.$
+G.2 Bounded-Size Items
+We now turn to the bounded item size case.
+Our algorithm, SKC-BOUND, runs SKC-OF, using a capacity of $\delta$ (instead of $1$), i.e., it terminates once the total size of accepted items exceeds $\delta$.
+Theorem G.4.
+$\frac{\mathrm{ADAPT}}{\operatorname{\mathbb{E}}[\textbf{SKC-BOUND}]}\leq O\left(\frac{1}{\delta}\right)$
+Proof.
+If $w_{\max}\geq\Phi(\delta)/4$, SKC-OF selects only the item with $w_{\max}$ and we obtain that:
+$\operatorname{\mathbb{E}}[\textbf{SKC-BOUND}]\geq\Phi(\delta)/4$.
+Otherwise: If all items have sizes $s_{i}^{j}<1-\delta$, then the policy never overflows (since it stops at total size $\delta$), achieving the same expected value as SKC-OF.
+However, if $s_{i}^{j}>1-\delta$ for some $i,j$, the process may terminate early, losing at most $\alpha w_{\max}$.
+Hence:
+$\operatorname{\mathbb{E}}[\textbf{SKC-BOUND}]$ $\geq\operatorname{\mathbb{E}}[\textbf{SKC-OF}]-\alpha w_{\max}\cdot\Pr[\exists\text{ item $i$ with }s_{i}^{j}>1-\delta]$
+$\geq\Phi(\delta)/4-\alpha w_{\max}\cdot\sum_{i=1}^{n}\Pr[s_{i}^{j}>1-\delta]$
+$\geq\Phi(\delta)/4-\alpha w_{\max}\cdot n\cdot\frac{1}{2n\alpha}$
+$\geq\Phi(\delta)/4-\Phi(\delta)/8$
+$=\Phi(\delta)/8.$
+The theorem follows from Lemma 3.8,
+$\frac{\mathrm{ADAPT}}{\operatorname{\mathbb{E}}[\textbf{SKC-BOUND}]}\leq\frac{\Phi(2)}{\Phi(\delta/8)}\leq\frac{\frac{16}{\delta}\Phi(\delta/8)}{\Phi(\delta/8)}=O\left(\frac{1}{\delta}\right)~.$
+We note that SKC-BOUND determines a non-adaptive order of item choices (as in the other algorithms), but it uses an adaptive stopping rule, stopping once the total realized sizes first exceed $\delta$.
+In Section G.3, we show that even for bounded-sized items, any non-adaptive policy is bounded away by a factor $\Omega(\sqrt{\alpha})$ from the optimal policy (Theorem G.5).
+G.3 $\Omega(\sqrt{\alpha})$ Adaptivity Gap of Bounded-Size Items
+Theorem G.5.
+There exists an instance $I$ of the MSKC-BOUND problem such that
+$\frac{\mathrm{STOP\_ADAPT}(I)}{\mathrm{NON\_ADAPT}(I)}=\Omega\bigl(\sqrt{\alpha}\bigr).$
+Proof.
+Consider an item with the following parameters:
+Property Item
+Value $v$ 22
+Cost $p$ $2-\epsilon^{2}$
+[…]
+We consider the case where $S^{\prime}<\frac{1}{2}$, and compute the expected utility per item.
+$\operatorname{\mathbb{E}}[w|S^{\prime}<\frac{1}{2}]=v\cdot\Pr[s\leq S^{\prime}]-p=2\cdot(1-\epsilon)-(2-\epsilon^{2})=2-2\epsilon-2+\epsilon^{2}=-2\epsilon+\epsilon^{2}<0$
+In expectation, the adaptive policy can include approximately $1/\epsilon$ items before reaching $S^{\prime}<1/2$, yielding expected total utility
+$\frac{1}{\epsilon}\cdot\epsilon^{2}=\epsilon.$
+For a non-adaptive policy, the number of attempted insertions is fixed in advance.
+If the policy inserts only one item, its expected utility is $v-p=\epsilon^{2}$.
+Let $\mathrm{utility}(i)$ denote the expected utility from the $i$th item.
+For $i>2$:
+$\mathrm{utility}(i)$ $=$ $\operatorname{\mathbb{E}}[\mathrm{utility}(i)|s_{j}=\epsilon^{n}\ \forall j<i]\cdot\Pr[s_{j}=\epsilon^{n}\quad\forall j<i]$
+$+\operatorname{\mathbb{E}}[\mathrm{utility}(i)|\exists!j<i:\ s_{j}>\epsilon^{n}]\cdot\Pr[\exists!j<i:\ s_{j}>\epsilon^{n}]$
+$=$ $(v-p)\cdot(1-\epsilon)^{i-1}+((v-p)\cdot(1-\epsilon)-p\cdot\epsilon)\cdot\Pr[\exists!j<i:\ s_{j}>0]$
+$=$ $\epsilon^{2}\cdot(1-\epsilon)^{i-1}+(\epsilon^{2}-2\epsilon)\cdot\Pr[\exists!j<i:\ s_{j}>0]$
+$\leq$ $\epsilon^{2}\cdot(1-\epsilon)^{i-1}+(\epsilon^{2}-2\epsilon)\cdot\Pr[s_{i-1}>0\ \wedge\ s_{j}=0\quad\forall j<i-1]$
+$=$ $\epsilon\cdot(1-\epsilon)^{i-2}(\epsilon\cdot(1-\epsilon)+\epsilon^{2}-2\epsilon)$
+$\leq$ $\epsilon\cdot(1-\epsilon)^{i-2}(\epsilon^{2}-\epsilon)<0.$
+In the above, the first inequality follows from the fact that $\epsilon^{2}-2\epsilon$ is negative, and that the probability that exactly one of the first i1i-1 items has size greater than $\epsilon^{n}$ is larger than the probability that only item i1i-1 has size greater than $\epsilon^{n}$.
+By Definition 2.1, the value of the IOR parameter is $\alpha=\frac{2-\epsilon^{2}}{\epsilon^{2}}=O(1/\epsilon^{2})$.
+Therefore, we have that a non-adaptive policy only loses from having a sequence larger than 2, and the optimal non-adaptive policy has a utility of $2\epsilon^{2}$.
+We conclude that for this instance, we have
+$\frac{\mathrm{STOP\_ADAPT}(I)}{\mathrm{NON\_ADAPT}(I)}=\frac{1}{2\epsilon}=O(\sqrt{\alpha}).$